### v1

# Memory-Dominated Quantum Criticality as a Universal Route to High-Temperature Superconductivity

Abstract
Understanding the dynamical origin of high–temperature superconductivity remains a central challenge of strongly correlated quantum matter. Conventional approaches to quantum criticality assume overdamped Markovian dissipation governed by Ohmic Landau damping. Here we show that infrared collective dynamics is instead generically controlled by the time–scale density of states (TDOS) of relaxation modes. Within the Martin–Siggia–Rose–Janssen–De Dominicis formalism, we derive an exact spectral representation of the collective susceptibility in terms of the TDOS. A finite infrared TDOS defines a new universality class of memory–dominated critical dynamics characterized by long–time kernels $K(t)\sim 1/t$ and nonanalytic dynamical response. This spectral reorganization produces a strong infrared amplification of the particle–particle channel, converting the marginal instabilities of BCS and Eliashberg theories into algebraic superconducting transitions. As a result, the transition temperature scales linearly with the infrared spectral weight of slow collective modes, naturally yielding superconducting domes and Uemura scaling without invoking bosonic glue or fine tuning. The same slow–mode reservoir governs anomalous normal–state dynamics, including long–time correlations and strange–metal behavior, providing a unified description of thermodynamic and dynamical phenomena in correlated superconductors. Our results establish dynamical spectral organization as a fundamental principle of quantum critical matter and identify memory–dominated criticality as a generic route to high–temperature superconductivity.
I I. Introduction
High–temperature superconductivity remains one of the central unsolved problems of strongly correlated quantum matter [1, 2]. Despite decades of intensive theoretical and experimental effort, no consensus has emerged on the microscopic mechanism responsible for robust pairing far above conventional phonon scales. A central difficulty lies in the intimate proximity of superconductivity to quantum critical regimes, where strong collective fluctuations dominate low–energy dynamics.
Within the prevailing theoretical paradigm, quantum critical fluctuations are assumed to be overdamped by fast electronic continua [3, 4]. This framework, formalized in Hertz–Millis theory, predicts a universal Ohmic Landau damping kernel of the form $|\omega|$ governing collective dynamics. While remarkably successful in weakly correlated metals, this assumption has increasingly come into tension with experiments in strongly correlated systems, including cuprate superconductors and strange metals, where anomalous temporal correlations, non-Markovian response, and $1/f$-like noise are ubiquitously observed [5, 6, 7, 8, 9].
These observations raise a fundamental question: Is overdamped Markovian critical dynamics truly universal, or does it rely on implicit assumptions about the structure of dynamical spectra that break down in strongly correlated regimes?
A key limitation of conventional approaches is their formulation in terms of a small number of collective order-parameter fields [10, 11]. Such low-dimensional reductions necessarily generate a finite set of soft modes and therefore preclude the emergence of a dense continuum of relaxation time scales. A genuine slow-mode spectrum can only arise when the full high-dimensional collective dynamical state space is retained.
In this work we show that infrared collective dynamics is generically controlled not by a single overdamped mode but by the time–scale density of states (TDOS) of correlated internal relaxation modes. By formulating a dynamical coarse–graining theory within the Martin–Siggia–Rose–Janssen–De Dominicis (MSRJD) framework [12, 13, 14], we derive an exact spectral representation of the collective susceptibility in terms of the TDOS.
We demonstrate that a TDOS scaling $\rho(\lambda)\sim\lambda^{\alpha}$ directly classifies universality classes of quantum critical dynamics. While the conventional Hertz–Millis kernel emerges only for $\alpha=1$, a finite TDOS at the origin ($\alpha=0$) produces a marginal nonanalytic retarded kernel proportional to $\mathrm{sgn}(\omega)$, corresponding in the time domain to a universal long–memory response $K(t)\sim 1/t$. This memory–dominated infrared dynamics represents a breakdown of Markovian overdamped criticality and defines a new universality class governed by self–organized slow modes rather than fast external baths.
Crucially, we show that such flat relaxation mode spectra arise naturally in strongly correlated collective systems, including self–organized dynamical manifolds explored in our previous work [15, 16, 17]. The resulting long–memory kernel dynamically amplifies intrinsic electronic pairing tendencies, providing a natural route to strong superconductivity without invoking finely tuned bosonic mediators.
Figure 1 summarizes the resulting dynamical universality map of correlated quantum matter. The infrared structure of the relaxation mode spectra determines collective kernels in the time domain and controls physical consequences including pairing enhancement, superconducting domes, Uemura scaling, and anomalous dynamical response [1, 18, 19]. Charge, spin, and pairing channels reorganize simultaneously within the same slow–mode reservoir, with superconductivity corresponding to condensation in the particle–particle sector of a common infrared spectral continuum.
Our results establish the relaxation mode spectra as a fundamental organizing principle of quantum critical dynamics and reveal memory–dominated criticality as a generic route toward high–temperature superconductivity. This framework reconciles anomalous dynamical phenomena observed in strange metals with robust pairing and provides a unified perspective on collective behavior across strongly correlated quantum matter [5, 7, 10].
The remainder of this paper is organized as follows. In Section II we derive the exact TDOS representation of the collective susceptibility and classify infrared universality classes. In Sections III and IV we analyze the resulting memory–dominated pairing kernel and its implications for superconducting instabilities. In Section V we discuss experimental consequences including superconducting domes, Uemura scaling, and dynamical anomalies. We conclude in Section VI with a summary and outlook.
Figure 1: Dynamical universality map of correlated quantum matter. (a) Infrared time–scale density of states $\rho(\lambda)$ defining distinct dynamical universality classes. Vanishing TDOS yields conventional Hertz–Millis and phonon dynamics, marginal spectra produce Ohmic damping, while a flat TDOS defines memory–dominated criticality. (b) Corresponding long–time kernels in the time domain, ranging from Markovian dissipation to non-Markovian power–law memory. (c) Physical consequences of memory–dominated criticality, including algebraic pairing enhancement, superconducting domes, Uemura scaling, strange metal transport, and long–time noise correlations.
II II. Emergence of a Slow-Mode Reservoir from Correlated Dynamical Spectra
Strongly correlated electronic matter hosts many intertwined collective fluctuation channels, including charge, spin, current, pairing, and phase–coherence sectors. These degrees of freedom continuously exchange energy and information and cannot, in general, be reduced to a small set of independent order parameters.
To capture this collective complexity in a unified manner, we represent the correlated system as a high–dimensional interacting dynamical system whose instantaneous collective configuration is encoded in a state vector
$x(t)\in\mathbb{R}^{N},$ (1)
which evolves under nonlinear many–body feedback.
The coarse-grained dynamics may therefore be written as a stochastic field theory
$\dot{x}(t)=F(x(t))+\xi(t),$ (2)
where $F$ encodes correlated interactions and $\xi$ is Gaussian noise with
$\langle\xi_{i}(t)\xi_{j}(t^{\prime})\rangle=2D\,\delta_{ij}\delta(t-t^{\prime}).$ (3)
II.1 A. Dynamical stability spectrum and time-scale density of states
Near correlated critical regimes the system does not soften a single collective mode but undergoes a reorganization of its full dynamical stability structure. Linearizing around a fluctuating metastable configuration $x^{*}$,
$\delta\dot{x}(t)=-A\,\delta x(t)+\xi(t),$ (4)
with
$A=-\left.\frac{\partial F}{\partial x}\right|_{x^{*}},$ (5)
yields a spectrum of collective relaxation modes
$Av_{i}=\lambda_{i}v_{i},$ (6)
whose retarded propagators are
$G_{i}(\omega)=\frac{1}{-i\omega+\lambda_{i}}.$ (7)
The infrared structure of correlated dynamics is therefore controlled by the distribution of relaxation rates $\lambda_{i}$. We characterize this spectrum by the time-scale density of states,
$\rho(\lambda)=\Big\langle\sum_{i}\delta(\lambda-\lambda_{i})\Big\rangle.$ (8)
In weakly interacting systems the TDOS vanishes as $\lambda\to 0$, reflecting a finite number of quasiparticle-like modes. By contrast, strongly correlated critical regimes generically generate an extensive accumulation of near-marginal directions in phase space,
$\rho(\lambda\to 0)=\rho_{0}\neq 0,$ (9)
forming a slow-mode reservoir that governs infrared collective dynamics.
This accumulation is the dynamical signature of correlated criticality: the system approaches marginal stability along many collective directions rather than along a single order-parameter coordinate.
II.2 B. Exact spectral representation from MSRJD field theory
To compute collective response functions while retaining the full dynamical spectrum, we employ the Martin–Siggia–Rose–Janssen–De Dominicis formalism associated with Eq. (2),
$S[x,\tilde{x}]=\int dt\,\tilde{x}(t)\big[\dot{x}(t)-F(x(t))\big]-D\tilde{x}^{2}(t).$ (10)
We focus on a slow collective field $\phi(\mathbf{r},t)$ obtained by coarse graining the full state vector and regard the remaining fast correlated degrees of freedom as internal relaxation modes.
The reorganized stability spectrum may be represented exactly as a continuum of Ornstein–Uhlenbeck modes $X_{\lambda}$ labeled by their decay rates,
$\dot{\phi}$ $=-(r+c\nabla^{2})\phi+\int_{0}^{\Lambda}d\lambda\,g(\lambda)X_{\lambda}+\eta,$ (11)
$\dot{X}_{\lambda}$ $=-\lambda X_{\lambda}+g(\lambda)\phi+\zeta_{\lambda}.$ (12)
Introducing response fields $\tilde{\phi}$ and $\tilde{X}_{\lambda}$, the MSRJD action becomes
$S=$ $\int dt\,d^{d}r\,\tilde{\phi}\Big[\dot{\phi}+(r+c\nabla^{2})\phi-\int_{0}^{\Lambda}d\lambda\,g(\lambda)X_{\lambda}\Big]-D_{\phi}\tilde{\phi}^{2}$
$+\int dt\,d^{d}r\int_{0}^{\Lambda}d\lambda\,\tilde{X}_{\lambda}[\dot{X}_{\lambda}+\lambda X_{\lambda}-g(\lambda)\phi]$
$-\int dt\,d^{d}r\int_{0}^{\Lambda}d\lambda\,D_{\lambda}\tilde{X}_{\lambda}^{2}.$ (13)
Since the reservoir variables enter quadratically, they can be integrated out exactly. In Fourier space this yields the effective collective action
$S_{\rm eff}=\int_{\omega,\mathbf{q}}\tilde{\phi}(-\omega,-\mathbf{q})\Big[-i\omega+r+cq^{2}+\Sigma^{R}(\omega)\Big]\phi(\omega,\mathbf{q})+\cdots,$ (14)
with retarded self-energy
$\Sigma^{R}(\omega)=\int_{0}^{\Lambda}d\lambda\,\frac{g(\lambda)^{2}}{\lambda-i\omega}.$ (15)
Separating real and imaginary parts,
$\frac{1}{\lambda-i\omega}=\frac{\lambda}{\lambda^{2}+\omega^{2}}+i\frac{\omega}{\lambda^{2}+\omega^{2}},$ (16)
gives
$\mathrm{Re}\,\Sigma^{R}(\omega)$ $=\int_{0}^{\Lambda}d\lambda\,g(\lambda)^{2}\frac{\lambda}{\lambda^{2}+\omega^{2}},$ (17)
$\mathrm{Im}\,\Sigma^{R}(\omega)$ $=\omega\int_{0}^{\Lambda}d\lambda\,g(\lambda)^{2}\frac{1}{\lambda^{2}+\omega^{2}}.$ (18)
Identifying the TDOS as
$\rho(\lambda)\equiv g(\lambda)^{2},$ (19)
the collective inverse susceptibility becomes
$\chi_{R}^{-1}(\omega,\mathbf{q})$ $=r+cq^{2}+\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{\lambda}{\lambda^{2}+\omega^{2}}$ (20)
$\quad+i\omega\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{1}{\lambda^{2}+\omega^{2}}.$
Equation (20) constitutes an exact spectral representation of collective infrared dynamics in terms of the dynamical stability spectrum. The complete field-theoretic derivation of the collective susceptibility kernel from the MSRJD formalism, including the explicit Gaussian integration over the relaxation-mode reservoir and the resulting self-energy structure, is presented in Appendix A.
II.3 C. Infrared universality classes
[…]
$\rho(\lambda)\sim\lambda^{\alpha},\qquad\lambda\to 0,$ (21)
the dissipative part of the collective kernel follows directly from the spectral representation in Eq. (20),
$\mathrm{Im}\,\chi_{R}^{-1}(\omega)=\omega\int_{0}^{\Lambda}d\lambda\,\frac{\rho(\lambda)}{\lambda^{2}+\omega^{2}}\;\sim\;\mathrm{sgn}(\omega)\,|\omega|^{\alpha}.$ (22)
Here $\omega$ denotes the frequency conjugate to the collective relaxation dynamics obtained after integrating out the slow–mode reservoir in the MSRJD formulation, rather than a microscopic quasiparticle energy scale.
It is therefore natural to write the long–wavelength dynamical response in the scaling form
$\chi^{-1}(\omega,\mathbf{q})=r+cq^{2}+\gamma|\omega|^{s},$ (23)
in direct analogy with the conventional Hertz–Millis representation. The TDOS scaling immediately implies the universal identification $s=\alpha$.
Under the renormalization–group rescaling $\mathbf{q}\to b\,\mathbf{q}$ and $\omega\to b^{z}\omega$, the spatial gradient term scales as $q^{2}\to b^{2}q^{2}$, while the dissipative term scales as $|\omega|^{s}\to b^{zs}|\omega|^{s}$. Requiring both contributions to remain of comparable order at criticality ($r=0$) fixes the dynamical exponent to
$z=\frac{2}{s}.$ (24)
In this framework the dissipation is generically non–Ohmic, $|\omega|^{s}$ with $s\neq 1$, reflecting the presence of an extensive reservoir of slow collective relaxation modes. Sub–Ohmic memory ($s<1$) implies $z>2$, corresponding to critical slowing down stronger than ordinary diffusive dynamics, whereas super–Ohmic behavior ($s>1$) yields $z<2$. Infrared universality classes are thus classified directly by the low–energy organization of relaxation modes encoded in the TDOS.
In particular, when the TDOS remains finite at the origin,
$\rho(\lambda)\xrightarrow{\lambda\to 0}\rho_{0},$ (25)
the dissipative kernel simplifies to
$\mathrm{Im}\,\chi_{R}^{-1}(\omega)$ $=\omega\int_{0}^{\Lambda}d\lambda\,\frac{\rho_{0}}{\lambda^{2}+\omega^{2}}$
$\xrightarrow{|\omega|\ll\Lambda}\frac{\pi}{2}\rho_{0}\,\mathrm{sgn}(\omega),$ (26)
while the reactive part acquires a logarithmic infrared singularity,
$\mathrm{Re}\,\chi_{R}^{-1}(\omega)=\int_{0}^{\Lambda}d\lambda\,\rho_{0}\frac{\lambda}{\lambda^{2}+\omega^{2}}\simeq\rho_{0}\ln\frac{\Lambda}{|\omega|}.$ (27)
Collective infrared dynamics is therefore governed by the marginal non-Markovian kernel
$\chi_{R}^{-1}(\omega,\mathbf{q})\simeq r+cq^{2}+\rho_{0}\ln\frac{\Lambda}{|\omega|}+i\frac{\pi}{2}\rho_{0}\,\mathrm{sgn}(\omega)+\cdots,$ (28)
corresponding in the time domain to a universal long-memory response $K(t)\sim 1/t$, which is derived explicitly in Appendix B.
This flat-TDOS fixed point represents a breakdown of the conventional overdamped critical dynamics underlying Hertz–Millis theory. Rather than arising from a fast electronic bath coupled to a single soft order-parameter mode, collective relaxation is controlled by the self-organized accumulation of an extensive continuum of slow internal modes.
The TDOS thus provides a unified organizing principle for infrared critical dynamics, with different universality classes emerging from the scaling of the dynamical stability spectrum itself.
III III. Memory–Enhanced Pairing Beyond Eliashberg Theory
The emergence of a slow–mode reservoir qualitatively reorganizes the infrared structure of the Cooper pairing channel. In conventional quantum critical superconductivity, collective fluctuations act as overdamped bosonic mediators whose Ohmic dynamics renders the pairing interaction marginal. Here we show that when collective dynamics is instead governed by a flat relaxation–rate density of states, long–time memory effects transform the Cooper channel into a relevant infrared instability.
III.1 A. Marginal pairing from Ohmic collective dynamics
Near metallic quantum critical points, fermions couple to overdamped collective modes with retarded propagator
$D_{R}^{-1}(\omega,\mathbf{q})=r+cq^{2}+\gamma|\omega|.$ (29)
The linear frequency dependence reflects Landau damping by the electronic continuum and produces a pairing kernel scaling as
$D(\omega)\sim\frac{1}{|\omega|}.$ (30)
The infrared shell correction to the dimensionless Cooper–channel coupling $g$ therefore reads
$dg\propto g^{2}\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{\omega^{\prime}}=g^{2}\ln b.$ (31)
Introducing the RG scale $l=\ln b$ yields
$\frac{dg}{dl}=Cg^{2},$ (32)
with $C>0$.
The pairing interaction is thus marginal and grows only logarithmically under infrared coarse graining. Physically, overdamped collective modes rapidly dissipate into the electronic continuum, continuously erasing temporal coherence of pairing fluctuations and preventing strong accumulation of superconducting correlations.
III.2 B. Infrared reorganization by slow collective modes
In the memory–dominated regime derived in Sec. II, collective dynamics is governed not by a single overdamped soft mode but by a continuum of slow relaxation modes characterized by the time–scale density of states $\rho(\lambda)$. After integrating out the relaxation reservoir within the MSRJD framework, the retarded collective susceptibility acquires the exact spectral form
$\chi_{R}^{-1}(\omega,\mathbf{q})=r+cq^{2}+\Sigma(\omega),\qquad\Sigma(\omega)=\int_{0}^{\Lambda}d\lambda\,\frac{\rho(\lambda)}{\lambda-i\omega}.$ (33)
For a flat TDOS in the infrared, $\rho(\lambda\to 0)=\rho_{0}$, the effective polarization kernel in the Cooper channel inherits the infrared scaling
$\Pi(\omega)\propto\int_{0}^{\Lambda}d\lambda\,\rho_{0}\frac{1}{\lambda^{2}+\omega^{2}}\sim\frac{1}{|\omega|},\qquad|\omega|\to 0.$ (34)
Unlike the conventional Ohmic case, this singular behavior does not originate from Landau damping by fast electronic continua. Instead, it emerges from the accumulation of an extensive continuum of near–marginal relaxation modes encoded in a flat TDOS. The resulting kernel corresponds in the time domain to a universal long–memory response $K(t)\sim 1/t$.
Although both Eliashberg theory and the present memory–dominated regime generate infrared kernels scaling as $1/|\omega|$, their physical origins are fundamentally distinct. In Eliashberg theory the singularity reflects instantaneous overdamped dissipation arising from coupling to a fast electronic bath [20, 21]. Here, by contrast, the same infrared scaling is a consequence of self–organized spectral accumulation of slow collective modes, producing intrinsically non–Markovian dynamics.
This dynamical reorganization qualitatively alters the Cooper channel: instead of marginal logarithmic enhancement characteristic of overdamped criticality, pairing fluctuations accumulate coherently across widely separated time scales. The slow–mode reservoir therefore converts the Cooper channel into a relevant infrared interaction.
III.3 C. Relevant infrared amplification of the Cooper channel
The infrared shell correction in the memory–dominated regime becomes
$dg\propto g^{2}\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{\omega^{\prime}\ln(\Lambda/\omega^{\prime})}.$ (35)
Introducing $u=\ln(\Lambda/\omega^{\prime})$ yields
$dg\simeq\alpha g\,dl,$ (36)
with $\alpha>0$ proportional to the infrared TDOS weight.
The resulting RG flow reads
$\frac{dg}{dl}=\alpha g,$ (37)
with solution $g(l)=g_{0}e^{\alpha l}$.
The Cooper–channel interaction therefore grows algebraically under infrared coarse graining, signaling a relevant superconducting instability rather than a marginal one. This qualitative contrast with Eliashberg theory is illustrated in Fig. 2. A detailed shell RG derivation demonstrating the crossover from logarithmic to power-law pairing growth is presented in Appendix C.
Physically, the slow–mode reservoir retains dynamical correlations across widely separated time scales, allowing collective fluctuations to act repeatedly on electronic degrees of freedom and dynamically amplify intrinsic electronic pairing tendencies.
Superconductivity in strongly correlated systems thus emerges from infrared spectral self–organization of collective relaxation dynamics, rather than from finely tuned exchange of specific bosonic mediators.
Figure 2: Infrared renormalization-group flow of the Cooper-channel interaction. Shown schematically is the flow of the dimensionless pairing coupling $g$ under infrared coarse graining. Ohmic Landau damping of collective modes produces marginal logarithmic growth, $dg/dl\sim g^{2}$, characteristic of conventional Eliashberg theory (blue). In contrast, memory–dominated dynamics generated by a flat relaxation–rate density of states yields a relevant infrared instability, $dg/dl=\alpha g$, leading to algebraic amplification of the pairing interaction (orange). The slow–mode reservoir therefore dynamically enhances intrinsic electronic pairing tendencies rather than merely renormalizing a fixed microscopic coupling.
IV IV. Pairing from Memory–Dominated Critical Dynamics
The dynamical framework developed in the preceding sections establishes that strongly correlated matter generically self–organizes into a regime dominated by an extensive reservoir of slow collective modes. Rather than producing conventional overdamped fluctuations described by Ohmic Eliashberg dynamics, the accumulation of near–marginal relaxation rates generates long–time memory kernels and nonanalytic infrared response.
Superconductivity in this setting emerges as a dynamical instability of the memory–dominated collective spectrum. The Cooper channel is reorganized by the slow–mode reservoir into a relevant infrared interaction, fundamentally distinct from marginal pairing mediated by fast bosonic modes. Consequently, even moderate intrinsic electronic pairing scales are dynamically amplified into an algebraic superconducting instability.
In this section we show explicitly how microscopic electronic correlations seed pairing and how the slow–mode TDOS reorganizes the Cooper channel into a memory–enhanced critical regime, yielding power–law growth of the effective pairing interaction and algebraic scaling of the transition temperature.
IV.1 A. Intrinsic electronic pairing and dynamical formulation
In strongly correlated materials proximate to Mott insulating phases, short–range Coulomb interactions naturally generate local singlet correlations through exchange processes. Starting from the Hubbard Hamiltonian,
$H=-t\sum_{\langle ij\rangle\sigma}c_{i\sigma}^{\dagger}c_{j\sigma}+U\sum_{i}n_{i\uparrow}n_{i\downarrow},$ (38)
projection to the strong–coupling sector yields the superexchange interaction
$H_{J}=J\sum_{\langle ij\rangle}\Big(\mathbf{S}_{i}\cdot\mathbf{S}_{j}-\frac{1}{4}n_{i}n_{j}\Big),\qquad J=\frac{4t^{2}}{U},$ (39)
favoring local singlet formation.
More generally, correlated electronic systems possess an intrinsic short–range pairing scale $E_{\rm pair}$ generated by electronic interactions, independent of long–wavelength collective mediators. This intrinsic scale provides the microscopic seed for superconductivity.
Within the dynamical framework developed in Sec. II, the pairing susceptibility assumes the universal tt–matrix form
$\chi_{\Delta}^{-1}(\omega)=\frac{1}{g_{\rm pair}}-\Pi(\omega),$ (40)
with $g_{\rm pair}\sim E_{\rm pair}^{-1}$. Critical dynamics does not create attraction; it amplifies this preexisting pairing tendency.
IV.2 B. Memory–dominated Cooper kernel from flat TDOS
Because single–particle propagators admit a relaxational spectral representation controlled by the slow–mode TDOS, the Cooper bubble reduces to a continuum integral over relaxation rates,
$\Pi(\omega)\sim\int_{0}^{\Lambda}d\lambda\,\rho_{\rm eff}(\lambda)\frac{1}{\lambda^{2}+\omega^{2}}.$ (41)
When the slow–mode reservoir develops,
$\rho_{\rm eff}(\lambda\to 0)=\rho_{0}\neq 0,$ (42)
the infrared behavior becomes
$\Pi(\omega)\xrightarrow{|\omega|\ll\Lambda}\frac{\pi\rho_{0}}{2}\,\frac{1}{|\omega|}.$ (43)
This nonanalytic enhancement reflects the accumulation of collective memory across all time scales. Unlike conventional Eliashberg kernels, which are marginal in the infrared, the present Cooper channel is dominated by long–time memory effects generated by the slow–mode TDOS.
IV.3 C. Algebraic superconducting instability and universality
The superconducting transition follows from the Thouless criterion
$\chi_{\Delta}^{-1}(\omega\sim T_{c})=0,$ (44)
which yields
$T_{c}\sim\rho_{0}\,E_{\rm pair},$ (45)
up to numerical factors of order unity. The finite-temperature regularization of the infrared divergence, leading to $\Pi(T)\sim 1/T$, is derived explicitly in Appendix D.
The exponential suppression characteristic of BCS theory is thus replaced by algebraic scaling governed by the infrared spectral weight of the slow–mode reservoir. High transition temperatures arise generically from dynamical amplification of moderate intrinsic pairing scales.
Because the TDOS exponent controls infrared dynamics independently of microscopic details, the pairing amplification mechanism is universal across correlated systems exhibiting flat relaxation spectra. The slow–mode reservoir therefore acts as a universal dynamical amplifier, converting electronic correlation energy into enhanced superconducting transition temperatures without invoking specific bosonic glue.
Figure 3: Experimental consequences of memory–dominated critical dynamics. (a) Superconducting dome emerging from truncation of the slow–mode reservoir away from criticality. The transition temperature follows $T_{c}(p)\sim E_{\rm pair}\rho_{0}(p)$ and is maximal near the correlated critical point $p_{c}$, where infrared spectral weight of collective relaxation modes is largest. Suppression of $T_{c}$ on both sides arises solely from the finite infrared extent of the slow–mode reservoir, without invoking competing orders or fine tuning. (b) Uemura scaling collapse between superconducting transition temperature and superfluid stiffness, $T_{c}\propto\rho_{s}$. Both pairing strength and phase coherence are controlled by the same infrared spectral weight $\rho_{0}$ of slow collective modes, yielding universal linear scaling across material families.
V V. Experimental consequences of memory–dominated critical dynamics
[…]
V.1 A. Infrared spectral control and automatic emergence of the dome
The linearized transition scale derived in Eq. (45) shows that the superconducting temperature is controlled by the infrared spectral weight $\rho_{0}$ of the slow–mode reservoir. Near correlated quantum criticality, $\rho_{0}$ is not constant but is governed by universal scaling.
Let $p$ denote a tuning parameter (e.g., doping or pressure) and $p_{c}$ the critical point. The mass term of the collective kernel obeys
$r(p)\sim|p-p_{c}|,$ (46)
so that the corresponding infrared cutoff scale takes the universal form
$\Lambda(p)\sim r^{z\nu}\sim|p-p_{c}|^{z\nu},$ (47)
with $z$ the dynamical exponent and $\nu$ the correlation-length exponent. Away from criticality the slow–mode reservoir is therefore truncated in the infrared, yielding an effective low-energy TDOS weight
$\rho_{0}(p)\sim\int_{0}^{\Lambda(p)}d\lambda\,\rho(\lambda).$ (48)
For a memory–dominated regime in which the TDOS remains finite at the origin, $\rho(\lambda\to 0)=\rho_{*}$, one obtains parametrically
$\rho_{0}(p)\sim\rho_{*}\,f\!\left(\frac{|p-p_{c}|}{p_{0}}\right),$ (49)
where $f$ is a universal decreasing scaling function and $p_{0}$ denotes a nonuniversal crossover scale.
Substituting Eq. (49) into Eq. (45) gives
$T_{c}(p)\sim E_{\rm pair}\,\rho_{*}\,f\!\left(\frac{|p-p_{c}|}{x_{0}}\right),$ (50)
which is maximal near criticality and suppressed on both sides, producing a superconducting dome without fine tuning.
The dome structure is thus a direct consequence of the competition between intrinsic short–range electronic pairing and the finite infrared extent of the slow–mode reservoir away from criticality. Importantly, no competing order parameters or additional microscopic suppression mechanisms are required: suppression of $T_{c}$ arises purely from the truncation of long–time memory encoded in the TDOS. In this interpretation, optimal doping corresponds to maximal infrared spectral reorganization rather than to the peak of any specific bosonic fluctuation.
V.2 B. Phase stiffness and Uemura scaling from the slow–mode reservoir
While the microscopic pairing seed is set by the intrinsic electronic scale $E_{\rm pair}$, the establishment of superconductivity requires global phase coherence of the Cooper collective field. In systems with low superfluid density, the transition may be controlled by phase stiffness rather than by gap formation.
The effective phase–only action for the Cooper field reads
$S_{\theta}=\frac{\rho_{s}}{2}\int d\tau d^{d}r\left[(\partial_{\tau}\theta)^{2}+v_{s}^{2}(\nabla\theta)^{2}\right],$ (51)
where $\rho_{s}$ is the superfluid stiffness. The phase–ordering temperature therefore scales as
$T_{\theta}\sim\rho_{s}.$ (52)
Within the slow–mode reservoir framework, the stiffness is controlled by the same infrared spectral weight that enhances pairing. Integrating out amplitude fluctuations yields the scaling
$\rho_{s}\sim E_{\rm pair}\,\rho_{0},$ (53)
reflecting the collective participation of slow modes in coherent transport. Combining Eqs. (52) and (53) gives
$T_{c}\sim\rho_{s}\sim E_{\rm pair}\,\rho_{0},$ (54)
establishing a direct proportionality between the superconducting transition temperature and superfluid density.
Equation (54) reproduces the empirical Uemura scaling observed across cuprates and other unconventional superconductors. In contrast to BCS superconductors, where phase stiffness is large and irrelevant to $T_{c}$, correlated systems operate in a regime where infrared collective dynamics simultaneously controls pairing and coherence. In the present framework, this is not an additional assumption but follows directly from the single infrared control parameter $\rho_{0}$.
V.3 C. Dynamical signatures: long–time correlations and anomalous dissipation
A distinctive prediction of the present theory concerns dynamical response in the normal state. As shown in Sec. II, a flat TDOS produces a marginal nonanalytic retarded kernel with
$\mathrm{Im}\,\chi_{R}^{-1}(\omega)\propto\mathrm{sgn}(\omega),$ (55)
corresponding in the time domain to a universal long–memory kernel $K(t)\sim 1/t$. This implies scale–free temporal correlations extending over broad time windows.
Such long–time tails naturally generate 1/f1/f–type noise spectra and anomalously slow relaxation, phenomena widely reported in strange metals and cuprate superconductors. Unlike conventional overdamped criticality, where correlations decay exponentially and dynamics is effectively Markovian, the memory–dominated regime predicts non-Markovian response with persistent temporal structure. Because the same slow–mode reservoir controls both pairing enhancement and dynamical response, superconducting and normal–state anomalies become two manifestations of a single infrared spectral reorganization.
V.4 D. Unified infrared scaling picture
We emphasize that all experimental signatures discussed above—the dome, Uemura scaling, and long–time dynamical correlations—originate from the same infrared parameter $\rho_{0}$, the low–energy spectral weight of the slow–mode reservoir.
The TDOS controls: (i) the extent of the slow–mode reservoir away from criticality [Eq. (47)], (ii) the dynamical amplification of intrinsic pairing tendencies [Eq. (50)], (iii) the superfluid stiffness and phase coherence [Eq. (54)], and (iv) the emergence of non-Markovian long–time response. Quantum criticality thus acts not merely as a source of fluctuations but as a self–organized dynamical amplifier that simultaneously governs thermodynamic and dynamical phenomena in strongly correlated superconductors.
This unified picture yields a concrete strategy for experimental tests: systems that exhibit enhanced low–frequency TDOS weight (or equivalently, strong long–time correlations) should simultaneously display increased pairing susceptibility, elevated $T_{c}$, and reduced phase stiffness consistent with Uemura scaling. The resulting real-frequency spectral function generated by the memory-dominated self-energy, including its universal scaling form and infrared structure, is derived in Appendix E.
VI VI. Discussion
The central result of this work is that infrared collective dynamics in strongly correlated quantum matter is governed not by local Markovian dissipation but by the time–scale density of states of collective relaxation modes. Criticality corresponds to a reorganization of the entire dynamical spectrum, and universality classes are naturally classified by the low–energy scaling of the TDOS. In particular, a finite TDOS at vanishing relaxation rate defines a memory–dominated infrared fixed point characterized by long–time kernels $K(t)\sim 1/t$.
This perspective clarifies the limited domain of validity of conventional coarse–grained field theories. Away from criticality, relaxation spectra are gapped in the infrared, leading to finite memory times and justifying reduction to local-in-time damping and a small number of collective order parameters. Near correlated critical regimes, however, relaxation rates condense into a continuous spectrum extending to zero. Internal modes cannot be integrated out: the full relaxation spectrum remains dynamically active at all time scales. The breakdown of Markovian coarse–graining is therefore a universal consequence of dynamical criticality rather than a model-dependent artifact.
This observation explains why Hertz–Millis–type approaches, which assume overdamped local dynamics of a small set of collective fields, cannot capture the infrared structure of strongly correlated critical matter. Criticality does not correspond to the softening of a single bosonic mode but to a spectral condensation of near–marginal collective degrees of freedom. The TDOS thus emerges as the natural organizing variable of infrared dynamics, replacing phenomenological damping parameters and providing a microscopic foundation for anomalous relaxation and transport.
A central physical consequence of this spectral reorganization is the qualitative enhancement of superconducting instabilities. Rather than requiring finely tuned bosonic mediators, superconductivity emerges generically from dynamical amplification of intrinsic electronic pairing tendencies by the slow–mode reservoir. The exponential suppression characteristic of weak-coupling BCS theory is replaced by algebraic scaling controlled by infrared spectral weight.
A single infrared parameter, $\rho_{0}$, governs thermodynamic and dynamical phenomena alike. The superconducting dome reflects truncation of long–time memory away from criticality, Uemura scaling follows from the shared spectral origin of pairing and phase coherence, and anomalous normal-state dynamics arises from the same non-Markovian kernel. Quantum criticality therefore acts as a self-organized dynamical amplifier that simultaneously controls fluctuations, transport, and superconducting order.
Although electrons propagate as Bloch quasiparticles on a crystalline lattice, collective particle–hole and particle–particle excitations form extended many-body coordinates not constrained to lattice energy eigenstates. In strongly correlated regimes these collective modes are generically overdamped by the electronic continuum, rendering a relaxation-rate description more fundamental than an energy-eigenmode picture. While conventional lattice theories organize collective behavior in terms of Hamiltonian eigenmodes, the present framework shows that near criticality the relevant organizing structure is instead the dynamically self-organized spectrum of relaxation rates.
This relaxation–rate organization is not a phenomenological assumption but a direct consequence of the fundamental structure of nonequilibrium quantum and classical dynamics. As shown in the Appendix F, both Liouvillian evolution in open quantum systems and generalized memory–kernel dynamics naturally decompose into continua of decay modes characterized by complex relaxation eigenvalues rather than by Hamiltonian energy eigenstates.
Even in the absence of spatial degrees of freedom, the Laplace–domain representation of non-Markovian dynamics generically yields continuous spectra of relaxation rates whenever long–time memory kernels are present. The emergence of a dense TDOS therefore reflects a universal mathematical property of dynamical systems with extended temporal correlations, not a model-dependent artifact of coarse–graining.
From this perspective, energy eigenmodes describe conservative reversible dynamics, while relaxation spectra encode the physically relevant irreversible organization governing collective behavior near criticality. Strongly correlated quantum matter near dynamical critical points is therefore more naturally classified by its relaxation-rate spectrum than by its Hamiltonian excitation spectrum.
The TDOS framework introduced here thus elevates relaxation modes to the status of fundamental collective coordinates, providing a principled dynamical basis for infrared universality, memory effects, and emergent macroscopic coherence.
Away from criticality the dynamical spectrum collapses onto a trivial fast mode. Near memory-dominated criticality an extensive continuum of slow collective modes emerges and governs macroscopic coherence. Superconducting pairing in this regime corresponds to condensation within a slow-mode reservoir rather than binding mediated by a single sharp bosonic excitation.
This unified picture resolves long-standing puzzles of unconventional superconductivity, including the coexistence of strange-metal behavior with robust pairing and the apparent universality of dome-shaped phase diagrams across disparate materials. Rather than invoking competing orders or material-specific bosonic glue, these phenomena arise naturally from universal infrared spectral structure.
More broadly, the TDOS framework establishes dynamical spectral organization as a fundamental principle of correlated quantum matter. Memory-dominated criticality represents a new universality class beyond traditional dissipative field theories, with implications extending to non-Fermi liquids, glassy dynamics, and self-organized critical systems.
Future directions include quantitative determination of TDOS spectra in microscopic models and experiments, extension to nonequilibrium driven systems, and exploration of engineered platforms in which slow-mode reservoirs can be dynamically controlled. In this light, high-temperature superconductivity appears not as an anomaly requiring special microscopic ingredients, but as a generic consequence of infrared spectral self-organization in strongly correlated matter.
Memory–dominated dynamics as a guiding principle for room-temperature superconductivity. Recent experimental efforts toward room-temperature superconductivity have largely focused on engineering stronger microscopic pairing interactions through pressure, lattice design, or exotic chemical compositions. Within the present framework, however, enhancing the bare pairing scale alone is insufficient. Our results show that the decisive control parameter for high transition temperatures is not the magnitude of an attractive interaction but the infrared dynamical organization of collective modes encoded in the relaxation-rate density of states.
A dense reservoir of near-marginal slow modes fundamentally reorganizes the Cooper channel, converting moderate intrinsic electronic pairing into an algebraically amplified instability with $T_{c}\sim\rho_{0}E_{\rm pair}$. From this perspective, many existing high-pressure hydride systems and correlated materials may fail to reach extreme transition temperatures not because pairing is weak, but because their collective spectra do not self-organize into a flat TDOS regime.
The present theory therefore suggests a new design principle for superconductivity beyond material-specific glue engineering: the deliberate stabilization of dynamical phases supporting extensive slow-mode continua. Controlling electronic structure, frustration, dimensionality, and feedback mechanisms to promote near-marginal collective dynamics may be as crucial as increasing microscopic attraction itself. Room-temperature superconductivity, in this view, corresponds to a dynamical phase characterized by maximal infrared memory rather than by extreme coupling strength.
VII VII. Conclusion
We have identified dynamical spectral organization as a fundamental principle governing infrared behavior in strongly correlated quantum matter. Formulating collective dynamics in terms of the time–scale density of states reveals that criticality is generically controlled not by local Markovian dissipation but by the accumulation of slow collective relaxation modes.
A finite infrared TDOS defines a memory–dominated universality class characterized by long–time kernels $K(t)\sim 1/t$ and nonanalytic dynamical response. Within this regime, intrinsic electronic pairing tendencies are dynamically amplified, converting the marginal instabilities of conventional theories into algebraic superconducting transitions governed by infrared spectral weight.
This mechanism provides a unified explanation of key experimental phenomena in unconventional superconductors, including superconducting domes, Uemura scaling of transition temperatures with superfluid stiffness, and anomalous long–time dynamical correlations in the normal state. These signatures arise generically from infrared spectral reorganization, without invoking specific bosonic glue, competing orders, or fine tuning.
More broadly, our results recast quantum criticality as a constructive dynamical process that enhances collective quantum phenomena through self–organized memory effects. Memory–dominated criticality thus represents a new paradigm beyond traditional dissipative field theories, with implications extending to non-Fermi liquids, glassy dynamics, and engineered quantum platforms.
In this perspective, high–temperature superconductivity emerges not as an exception requiring special microscopic ingredients, but as a natural consequence of infrared spectral self–organization in strongly correlated matter. Future exploration of TDOS-controlled dynamics may enable systematic design of materials and devices that exploit memory–dominated criticality to stabilize and enhance collective quantum states.
Acknowledgements—This work was partially supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (IITP-RS-2025-02214780).
The author acknowledges the support of ChatGPT (GPT-5, OpenAI) for assistance in literature review and conceptual structuring during early development.
Supplementary Materials
Appendix A Appendix A: Gaussian MSRJD integration of a relaxation-mode reservoir and emergence of the self-energy
In this appendix we derive, without omitting algebraic steps, how a continuum of relaxation (Ornstein–Uhlenbeck) modes generates a nonlocal-in-time memory kernel and an associated self-energy $\Sigma(\omega)$ upon Gaussian functional integration in the Martin–Siggia–Rose–Janssen–De Dominicis (MSRJD) formalism. The derivation is standard in structure (Gaussian path integration), but we present it in a form tailored to our “slow-mode reservoir” interpretation.
(I) Coupled stochastic dynamics: collective variable plus OU reservoir: We consider a collective degree of freedom $x(t)$ coupled linearly to a continuum of reservoir modes $X_{\lambda}(t)$ labeled by their relaxation rates $\lambda>0$. The coupled stochastic dynamics is
$\dot{x}(t)$ $=-r\,x(t)+\int d\lambda\;g(\lambda)\,X_{\lambda}(t)+\eta(t),$ (S1)
$\dot{X}_{\lambda}(t)$ $=-\lambda\,X_{\lambda}(t)+g(\lambda)\,x(t)+\xi_{\lambda}(t).$ (S2)
The noises are assumed Gaussian, stationary, and white:
$\langle\eta(t)\eta(t^{\prime})\rangle$ $=2D_{x}\,\delta(t-t^{\prime}),$ (S3)
$\langle\xi_{\lambda}(t)\,\xi_{\lambda^{\prime}}(t^{\prime})\rangle$ $=2D_{\lambda}\,\delta(\lambda-\lambda^{\prime})\,\delta(t-t^{\prime}).$ (S4)
No equilibrium fluctuation–dissipation relation is assumed; $D_{\lambda}$ may be arbitrary. Our goal is to integrate out $\{X_{\lambda}\}$ and obtain an exact effective MSRJD action for (x,x~)($\tilde{x}$) containing a memory kernel.
(II) MSRJD action for the coupled system: For a generic additive-noise Langevin equation $\dot{y}=f(y)+\zeta$ with $\langle\zeta(t)\zeta(t^{\prime})\rangle=2D\,\delta(t-t^{\prime})$, the MSRJD functional integral may be written (Ito convention) as
$Z$ $=\int\mathcal{D}y\,\mathcal{D}\tilde{y}\;$ (S5)
$\exp\!\left[-\int dt\;\tilde{y}(t)\big(\dot{y}(t)-f(y(t))\big)+\int dt\;D\,\tilde{y}(t)^{2}\right],$
up to an overall normalization independent of $y$. Applying (S5) to (S1)–(S2) gives
$Z=\int\mathcal{D}x\,\mathcal{D}\tilde{x}\;\prod_{\lambda}\mathcal{D}X_{\lambda}\,\mathcal{D}\tilde{X}_{\lambda}\;e^{-S[x,\tilde{x},\{X_{\lambda},\tilde{X}_{\lambda}\}]},$ (S6)
with total action
$S=S_{x}+\int d\lambda\;S_{\lambda},$ (S7)
where
$S_{x}$ $=\int dt\;\tilde{x}(t)\!\left[\dot{x}(t)+r\,x(t)-\int d\lambda\;g(\lambda)\,X_{\lambda}(t)\right]$
$\quad-\int dt\;D_{x}\,\tilde{x}(t)^{2},$ (S8)
$S_{\lambda}$ $=\int dt\;\tilde{X}_{\lambda}(t)\!\left[\dot{X}_{\lambda}(t)+\lambda\,X_{\lambda}(t)-g(\lambda)\,x(t)\right]$
$\quad-\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}.$ (S9)
(III) Isolating a single $\lambda$-mode contribution: Fix $\lambda$ and collect all terms in $S$ that contain $X_{\lambda}$ or $\tilde{X}_{\lambda}$. From (S8) we extract
$S_{x}\supset-\int dt\;\tilde{x}(t)\,g(\lambda)\,X_{\lambda}(t).$ (S10)
Combining (S10) with (S9) yields the $\lambda$-sector action
$S^{(\lambda)}[X_{\lambda},\tilde{X}_{\lambda};x,\tilde{x}]=\int dt\;\tilde{X}_{\lambda}(t)\big(\dot{X}_{\lambda}(t)+\lambda X_{\lambda}(t)\big)$ (S11)
$\quad-\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,x(t)-\int dt\;g(\lambda)\,\tilde{x}(t)\,X_{\lambda}(t)$
$\quad-\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}.$
In what follows, $x,\tilde{x}$ are treated as external sources.
(IV) Functional integration over $X_{\lambda}$: emergence of a delta constraint: Define the linear operator
$\mathcal{L}_{\lambda}\equiv\partial_{t}+\lambda.$ (S12)
Then the first term in (S11) is
$\int dt\;\tilde{X}_{\lambda}(t)\,\mathcal{L}_{\lambda}X_{\lambda}(t)=\int dt\;\tilde{X}_{\lambda}(t)\big(\partial_{t}X_{\lambda}(t)+\lambda X_{\lambda}(t)\big).$ (S13)
We now rewrite this term so that $X_{\lambda}$ appears without derivatives. Using integration by parts,
$\int dt\;\tilde{X}_{\lambda}(t)\,\partial_{t}X_{\lambda}(t)$ (S14)
$\quad=\big[\tilde{X}_{\lambda}(t)\,X_{\lambda}(t)\big]_{t_{i}}^{t_{f}}-\int dt\;(\partial_{t}\tilde{X}_{\lambda}(t))\,X_{\lambda}(t).$
Assuming boundary terms vanish (e.g. $t_{i}\to$-$\infty$, $t_{f}\to+\infty$ with sufficient decay, or appropriate causal boundary conditions), we drop the bracketed term and obtain
$\int dt\;\tilde{X}_{\lambda}(t)\,\partial_{t}X_{\lambda}(t)=-\int dt\;(\partial_{t}\tilde{X}_{\lambda}(t))\,X_{\lambda}(t).$ (S15)
Therefore,
$\int dt\;\tilde{X}_{\lambda}(t)\,\mathcal{L}_{\lambda}X_{\lambda}(t)=-\int dt\;(\partial_{t}\tilde{X}_{\lambda}(t))\,X_{\lambda}(t)$ (S16)
$+\int dt\;\lambda\,\tilde{X}_{\lambda}(t)\,X_{\lambda}(t)=\int dt\;X_{\lambda}(t)\big(-\partial_{t}+\lambda\big)\tilde{X}_{\lambda}(t).$
Define the adjoint operator
$\mathcal{L}_{\lambda}^{\dagger}\equiv-\partial_{t}+\lambda.$ (S17)
Then (S11) becomes
$S^{(\lambda)}$ $=\int dt\;X_{\lambda}(t)\Big(\mathcal{L}_{\lambda}^{\dagger}\tilde{X}_{\lambda}(t)-g(\lambda)\tilde{x}(t)\Big)$ (S18)
$\quad-\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,x(t)-\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}.$
Now $X_{\lambda}$ enters only linearly. Hence its functional integral produces a functional delta constraint:
$\int\mathcal{D}X_{\lambda}\;\exp\!\left\{-\int dt\;X_{\lambda}(t)\,A(t)\right\}$ $\propto\delta[A(t)],$ (S19)
where here
$A(t)\equiv\mathcal{L}_{\lambda}^{\dagger}\tilde{X}_{\lambda}(t)-g(\lambda)\tilde{x}(t).$ (S20)
Therefore,
[…]
$\quad\exp\!\left[+\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,x(t)-\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}\right].$
The Jacobian factor associated with the functional constraint, $|\det\mathcal{L}_{\lambda}^{\dagger}|^{-1}$, is independent of $x,\tilde{x}$ and can therefore be absorbed into the overall normalization of $Z$.
(V) Solving the delta constraint for $\tilde{X}_{\lambda}$ via Green functions: The constraint
$\mathcal{L}_{\lambda}^{\dagger}\tilde{X}_{\lambda}(t)=g(\lambda)\tilde{x}(t),\qquad\mathcal{L}_{\lambda}^{\dagger}=-\partial_{t}+\lambda,$ (S22)
is a linear inhomogeneous ODE for $\tilde{X}_{\lambda}(t)$. Introduce the Green function $G_{\lambda}^{A}(t-t^{\prime})$ solving
$(-\partial_{t}+\lambda)\,G_{\lambda}^{A}(t-t^{\prime})=\delta(t-t^{\prime}).$ (S23)
Because the constraint is written with the adjoint operator $\mathcal{L}_{\lambda}^{\dagger}=-\partial_{t}+\lambda$, its Green function carries advanced support (equivalently, the retarded kernel will reappear after relabelling, ensuring causality of the $x$-sector response).
The solution with advanced support is
$G_{\lambda}^{A}(t)=\Theta(-t)\,e^{\lambda t}.$ (S24)
(Indeed, for $t\neq 0$, $(-\partial_{t}+\lambda)G_{\lambda}^{A}=0$ and the jump at $t=0$ produces the delta distribution.) Multiplying (S22) by $G_{\lambda}^{A}$ and integrating, we obtain
$\tilde{X}_{\lambda}(t)$ $=g(\lambda)\int dt^{\prime}\;G_{\lambda}^{A}(t-t^{\prime})\,\tilde{x}(t^{\prime}).$ (S25)
Substituting (S25) into (S21) eliminates $\tilde{X}_{\lambda}$ entirely and yields an effective contribution to the (x,x~)($\tilde{x}$) action.
(VI) Effective action: memory kernel (self-energy) and induced colored noise: We now evaluate the two remaining terms in the exponent of (S21) using (S25).
1. The $\tilde{X}_{\lambda}x$ term (deterministic memory kernel).
Start from
$\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,x(t).$ (S26)
Insert (S25):
$\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,x(t)=\int dt\,dt^{\prime}\;g(\lambda)^{2}\,G_{\lambda}^{A}(t-t^{\prime})\,\tilde{x}(t^{\prime})\,x(t).$ (S27)
Relabel integration variables $(t,t^{\prime})\to(t^{\prime},t)$:
$\int dt\,dt^{\prime}\;g(\lambda)^{2}\,G_{\lambda}^{A}(t^{\prime}-t)\,\tilde{x}(t)\,x(t^{\prime}).$ (S28)
Using the identity $G_{\lambda}^{A}(t^{\prime}-t)=G_{\lambda}^{R}(t-t^{\prime})$ with the retarded Green function
$G_{\lambda}^{R}(t)\equiv\Theta(t)\,e^{-\lambda t},\qquad(\partial_{t}+\lambda)G_{\lambda}^{R}(t)=\delta(t),$ (S29)
we obtain
$\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,x(t)$ $=\int dt\,dt^{\prime}\;\tilde{x}(t)\,\underbrace{g(\lambda)^{2}\,G_{\lambda}^{R}(t-t^{\prime})}_{\equiv K_{\lambda}(t-t^{\prime})}\,x(t^{\prime}).$ (S30)
Thus a single OU mode contributes the causal memory kernel
$K_{\lambda}(t-t^{\prime})=g(\lambda)^{2}\,\Theta(t-t^{\prime})\,e^{-\lambda(t-t^{\prime})}.$ (S31)
2. The $\tilde{X}_{\lambda}^{2}$ term (induced colored noise kernel).
Next consider
$\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}.$ (S32)
Insert (S25):
$\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}$ (S33)
$=\int dt\,dt^{\prime}\,dt^{\prime\prime}\;D_{\lambda}\,g(\lambda)^{2}\,G_{\lambda}^{A}(t-t^{\prime})\,G_{\lambda}^{A}(t-t^{\prime\prime})\,\tilde{x}(t^{\prime})\,\tilde{x}(t^{\prime\prime}).$ (S34)
Define the induced (generally nonlocal) noise kernel
$\mathcal{N}_{\lambda}(t^{\prime},t^{\prime\prime})\equiv D_{\lambda}\,g(\lambda)^{2}\int dt\;G_{\lambda}^{A}(t-t^{\prime})\,G_{\lambda}^{A}(t-t^{\prime\prime}),$ (S35)
so that
$\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}=\int dt^{\prime}\,dt^{\prime\prime}\;\tilde{x}(t^{\prime})\,\mathcal{N}_{\lambda}(t^{\prime},t^{\prime\prime})\,\tilde{x}(t^{\prime\prime}).$ (S36)
In equilibrium one may impose a fluctuation–dissipation relation linking $\mathcal{N}_{\lambda}$ and $K_{\lambda}$, but this is not required for the deterministic self-energy structure derived below.
3. Summing the reservoir and assembling $S_{\mathrm{eff}}$.
Combining (S30)–(S36) for all $\lambda$ yields the exact effective action
$S_{\mathrm{eff}}[x,\tilde{x}]=\int dt\;\tilde{x}(t)\big(\dot{x}(t)+rx(t)\big)-\int dt\;D_{x}\,\tilde{x}(t)^{2}$
$-\int dt\,dt^{\prime}\;\tilde{x}(t)\,K(t-t^{\prime})\,x(t^{\prime})-\int dt\,dt^{\prime}\;\tilde{x}(t)\,\mathcal{N}(t,t^{\prime})\,\tilde{x}(t^{\prime}),$ (S37)
where the total memory kernel is the reservoir sum
$K(t-t^{\prime})=\int d\lambda\;K_{\lambda}(t-t^{\prime})=\int d\lambda\;g(\lambda)^{2}\,\Theta(t-t^{\prime})\,e^{-\lambda(t-t^{\prime})},$ (S38)
and likewise $\mathcal{N}=\int d\lambda\,\mathcal{N}_{\lambda}$.
(VI) Frequency-domain self-energy $\Sigma(\omega)$ from the memory kernel: To connect (S37) to the standard response-function form, we Fourier transform using
$f(\omega)=\int_{-\infty}^{\infty}dt\;e^{i\omega t}\,f(t),\qquad f(t)=\int\frac{d\omega}{2\pi}\;e^{-i\omega t}\,f(\omega).$ (S39)
The bilinear term involving $K$ becomes
$\int dt\,dt^{\prime}\;\tilde{x}(t)\,K(t-t^{\prime})\,x(t^{\prime})$ $=\int\frac{d\omega}{2\pi}\;\tilde{x}(-\omega)\,K(\omega)\,x(\omega),$ (S40)
where $K(\omega)$ is the Fourier transform of the causal kernel (S38):
$K(\omega)$ $=\int_{-\infty}^{\infty}dt\;e^{i\omega t}\,K(t)=\int d\lambda\;g(\lambda)^{2}\int_{0}^{\infty}dt\;e^{i\omega t}e^{-\lambda t}.$ (S41)
The time integral is elementary:
$\int_{0}^{\infty}dt\;e^{-(\lambda-i\omega)t}$ $=\left[\frac{-1}{\lambda-i\omega}e^{-(\lambda-i\omega)t}\right]_{0}^{\infty}=\frac{1}{\lambda-i\omega},$
$\qquad(\mathrm{Re}\,\lambda>0).$ (S42)
Substituting (S42) into (S41) gives the exact spectral representation
$K(\omega)=\int d\lambda\;\frac{g(\lambda)^{2}}{\lambda-i\omega}.$ (S43)
In the MSRJD response sector, the retarded inverse propagator for $x$ is read off from the $\tilde{x}(-\omega)x(\omega)$ coefficient in $S_{\mathrm{eff}}$:
$G_{R}^{-1}(\omega)=-i\omega+r+K(\omega).$ (S44)
It is therefore natural to define the (retarded) self-energy as
$\Sigma_{R}(\omega)\equiv K(\omega)=\int d\lambda\;\frac{g(\lambda)^{2}}{\lambda-i\omega}.$ (S45)
(Our convention is $\Sigma_{R}(\omega)\equiv K(\omega)$ so that $G_{R}^{-1}(\omega)=-i\omega+r+\Sigma_{R}(\omega)$; other sign conventions for the self-energy appear in the literature.)
Equations (S43)–(S45) demonstrate that integrating out a continuum of OU relaxation modes produces a frequency- dependent self-energy whose analytic structure is determined by the relaxation-rate spectrum and couplings.
(VII) Connection to TDOS (relaxation-rate density of states): If the couplings are sufficiently structureless over the low-$\lambda$ sector, one may absorb $g(\lambda)^{2}$ into an effective spectral density. Defining
$\mathcal{J}(\lambda)\equiv g(\lambda)^{2},$ (S46)
we can write
$\Sigma_{R}(\omega)=\int d\lambda\;\frac{\mathcal{J}(\lambda)}{\lambda-i\omega}.$ (S47)
Equivalently, starting from a discrete set of modes labeled by $\alpha$,
$\Sigma_{R}(\omega)=\sum_{\alpha}\frac{g_{\alpha}^{2}}{\lambda_{\alpha}-i\omega},$ (S48)
the continuum limit gives
$\Sigma_{R}(\omega)$ $=\int d\lambda\;\rho(\lambda)\,\frac{\overline{g^{2}}(\lambda)}{\lambda-i\omega}\equiv\int d\lambda\;\frac{\mathcal{J}(\lambda)}{\lambda-i\omega},$
$\qquad\mathcal{J}(\lambda)\equiv\rho(\lambda)\,\overline{g^{2}}(\lambda),$ (S49)
where $\rho(\lambda)$ is the relaxation-rate density of states and $\overline{g^{2}}(\lambda)$ denotes the mean squared coupling of modes with rate $\lambda$.
In our framework the natural object is the time-scale density of states $\rho(\lambda)$ of relaxation rates (e.g. from Jacobian eigenvalues). In the simplest case $\overline{g^{2}}(\lambda)$ varies weakly over the slow sector so that $\mathcal{J}(\lambda)\propto\rho(\lambda)$ in the infrared, and therefore
$\Sigma_{R}(\omega)\propto\int d\lambda\;\frac{\rho(\lambda)}{\lambda-i\omega}.$ (S50)
This makes explicit, at the level of a controlled Gaussian MSRJD integration, why the relaxation-rate spectrum provides the natural basis for the effective dynamics and response of the collective variable.
(VIII) Coarse-grained field extension: We promote the collective coordinate to a coarse-grained field $\phi(\mathbf{r},t)$. At the Gaussian level the MSRJD action reads
$S_{0}=\int dt\,d^{d}r\;\tilde{\phi}\big(\partial_{t}+\Gamma(r-c\nabla^{2})\big)\phi-\int dt\,d^{d}r\,D\,\tilde{\phi}^{2}.$ (S51)
Coupling $\phi$ locally to an OU reservoir $X_{\lambda}(\mathbf{r},t)$ and integrating out $\{X_{\lambda},\tilde{X}_{\lambda}\}$ yields the nonlocal-in-time term
$S_{\rm eff}\supset-\int dt\,dt^{\prime}\,d^{d}r\;\tilde{\phi}(\mathbf{r},t)\,K(t-t^{\prime})\,\phi(\mathbf{r},t^{\prime}),$ (S52)
with $K(\omega)=\int d\lambda\,g(\lambda)^{2}/(\lambda-i\omega)\equiv\Sigma_{R}(\omega)$. Hence,
$\chi^{-1}(\omega,\mathbf{q})=r+cq^{2}+\Sigma(\omega)$ (S53)
(up to the conventional factor $\Gamma$ and the $i\omega$ term in the retarded response).
(IX) Summary of the exact Gaussian result: Starting from the coupled Langevin system (S1)–(S2), the MSRJD action (S8)–(S9) is Gaussian in the reservoir variables. Performing the functional integration over $\{X_{\lambda},\tilde{X}_{\lambda}\}$ yields the exact nonlocal effective action (S37) with a causal memory kernel (S38). In frequency space the kernel takes the spectral form (S43), which is naturally identified with the retarded self-energy (S45). This derivation is purely algebraic and involves no approximation beyond standard boundary-condition assumptions required for integration by parts.
Appendix B Appendix B: Exact emergence of a memory kernel from integrating out an OU reservoir
In this Appendix we provide a self-contained derivation showing that integrating out a continuum of relaxational Ornstein–Uhlenbeck (OU) modes generates a nonlocal-in-time (memory) kernel for the remaining collective coordinate. The result makes explicit that long-memory response is the generic outcome of eliminating hidden internal degrees of freedom, and that a flat time-scale density of states yields the universal tail $K(t)\sim 1/t$.
(I) Coupled Markovian dynamics: collective field plus relaxational reservoir: For clarity we first suppress spatial dependence and focus on temporal structure; the extension to $\mathbf{r}$ (including $cq^{2}$ terms) is immediate by Fourier transforming in space. We consider a slow collective variable $\phi(t)$ linearly coupled to a continuum of relaxational OU modes $X_{\lambda}(t)$ labeled by decay rate $\lambda\geq 0$:
$\dot{\phi}(t)$ $=-r\,\phi(t)+\int_{0}^{\Lambda}d\lambda\,g(\lambda)\,X_{\lambda}(t)+\eta(t),$ (S54)
$\dot{X}_{\lambda}(t)$ $=-\lambda\,X_{\lambda}(t)+g(\lambda)\,\phi(t)+\zeta_{\lambda}(t).$ (S55)
Here $r>0$ is the bare restoring (mass) parameter, $\Lambda$ is an ultraviolet cutoff on relaxation rates, and $\eta$ and $\zeta_{\lambda}$ are (possibly independent) noise sources. Equations (S54)–(S55) define a Markovian (first-order) dynamical system in the enlarged state space $\{\phi,X_{\lambda}\}$.
(II) Exact solution for the OU reservoir modes: Equation (S55) is linear and can be solved exactly by the integrating-factor method. Multiply (S55) by $e^{\lambda t}$:
$e^{\lambda t}\dot{X}_{\lambda}(t)+\lambda e^{\lambda t}X_{\lambda}(t)=e^{\lambda t}\big[g(\lambda)\phi(t)+\zeta_{\lambda}(t)\big].$ (S56)
The left-hand side is the total derivative
$\frac{d}{dt}\Big(e^{\lambda t}X_{\lambda}(t)\Big)=e^{\lambda t}\big[g(\lambda)\phi(t)+\zeta_{\lambda}(t)\big].$ (S57)
Integrate from an initial time $t_{0}$ to $t$:
$e^{\lambda t}X_{\lambda}(t)-e^{\lambda t_{0}}X_{\lambda}(t_{0})=\int_{t_{0}}^{t}ds\,e^{\lambda s}\big[g(\lambda)\phi(s)+\zeta_{\lambda}(s)\big].$ (S58)
Solving for $X_{\lambda}(t)$ yields the exact causal representation
$X_{\lambda}(t)$ (S59)
$\quad=e^{-\lambda(t-t_{0})}X_{\lambda}(t_{0})+\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}\Big[g(\lambda)\phi(s)+\zeta_{\lambda}(s)\Big].$
This expression already shows that the reservoir coordinate at time $t$ depends on the entire history $\{\phi(s)\}_{s<t}$.
(III) Substitution into the collective equation and emergence of a memory kernel: Substituting (S59) into the collective equation (S54) gives
$\dot{\phi}(t)$ $=-r\,\phi(t)+\int_{0}^{\Lambda}d\lambda\,g(\lambda)\,$
$\Bigg[e^{-\lambda(t-t_{0})}X_{\lambda}(t_{0})+\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}\Big(g(\lambda)\phi(s)+\zeta_{\lambda}(s)\Big)\Bigg]$
$+\eta(t)$
$=-r\,\phi(t)+\int_{0}^{\Lambda}d\lambda\,g(\lambda)\,e^{-\lambda(t-t_{0})}X_{\lambda}(t_{0})$
$\qquad+\int_{0}^{\Lambda}d\lambda\,g(\lambda)\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}g(\lambda)\phi(s)$
$\qquad+\int_{0}^{\Lambda}d\lambda\,g(\lambda)\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}\zeta_{\lambda}(s)+\eta(t).$ (S60)
We now isolate the term proportional to the history $\phi(s)$:
$\int_{0}^{\Lambda}d\lambda\,g(\lambda)\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}g(\lambda)\phi(s)$
$\quad=\int_{t_{0}}^{t}ds\,\Bigg[\int_{0}^{\Lambda}d\lambda\,g(\lambda)^{2}e^{-\lambda(t-s)}\Bigg]\phi(s).$ (S61)
This motivates the definition of the (causal) memory kernel
$K(t-s)\equiv\int_{0}^{\Lambda}d\lambda\,g(\lambda)^{2}\,e^{-\lambda(t-s)},\qquad t\geq s,$ (S62)
so that (S61) becomes simply
$\int_{0}^{\Lambda}d\lambda\,g(\lambda)\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}g(\lambda)\phi(s)=\int_{t_{0}}^{t}ds\,K(t-s)\,\phi(s).$ (S63)
With this definition, Eq. (S60) can be written as
$\dot{\phi}(t)=-r\,\phi(t)+\int_{t_{0}}^{t}ds\,K(t-s)\phi(s)+\eta(t)+\xi_{\rm eff}(t)+J_{\rm ic}(t),$ (S64)
where we have collected the remaining contributions into an effective noise term (generated by the reservoir noise) and an initial-condition transient:
$\xi_{\rm eff}(t)$ $\equiv\int_{0}^{\Lambda}d\lambda\,g(\lambda)\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}\zeta_{\lambda}(s),$ (S65)
$J_{\rm ic}(t)$ $\equiv\int_{0}^{\Lambda}d\lambda\,g(\lambda)\,e^{-\lambda(t-t_{0})}X_{\lambda}(t_{0}).$ (S66)
Equation (S64) is an exact generalized Langevin equation for $\phi(t)$. Although the enlarged system (S54)–(S55) is Markovian, eliminating the hidden reservoir coordinates produces a non-Markovian (memory) term and, generally, colored effective noise.
Long-time limit.
For $t-t_{0}\gg\Lambda^{-1}$, the transient $J_{\rm ic}(t)$ decays. More precisely, decay holds provided the effective spectral weight $g(\lambda)X_{\lambda}(t_{0})$ is integrable near $\lambda=0$ so that the $\lambda\to 0$ sector does not produce a non-decaying contribution.
(IV) TDOS form and Laplace-transform structure: Defining the time-scale density of states as
$\rho(\lambda)\equiv g(\lambda)^{2},$ (S67)
(Here $\rho(\lambda)$ should be understood as an effective spectral weight for the coupling of relaxation rates to $\phi$; in a microscopic discrete-mode representation one naturally has $\rho(\lambda)\overline{g^{2}}(\lambda)$, and we absorb any slowly varying $\overline{g^{2}}(\lambda)$ into $\rho(\lambda)$ for notational simplicity.)
the memory kernel (S62) becomes
$K(t)=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\,e^{-\lambda t},\qquad t\geq 0.$ (S68)
Thus $K(t)$ is precisely the Laplace transform of the TDOS (with cutoff $\Lambda$). In this sense the TDOS is the fundamental spectral object, and the time-domain memory kernel is its direct transform.
(V) Flat TDOS implies universal long-memory tail $K(t)\sim 1/t$: If the TDOS is finite at the origin (“flat” TDOS),
$\rho(\lambda)\xrightarrow{\lambda\to 0}\rho_{0},$ (S69)
then for times $t\gg\Lambda^{-1}$ the infrared part of (S68) dominates and we may approximate $\rho(\lambda)\simeq\rho_{0}$ over the relevant range. In that case the kernel can be evaluated exactly:
$K(t)$ $=\int_{0}^{\Lambda}d\lambda\,\rho_{0}\,e^{-\lambda t}=\rho_{0}\int_{0}^{\Lambda}d\lambda\,e^{-\lambda t}.$ (S70)
Perform the change of variables $u=\lambda t$ (so $d\lambda=du/t$):
$K(t)$ $=\rho_{0}\int_{0}^{\Lambda t}\frac{du}{t}\,e^{-u}=\frac{\rho_{0}}{t}\int_{0}^{\Lambda t}du\,e^{-u}=\frac{\rho_{0}}{t}\Big(1-e^{-\Lambda t}\Big).$ (S71)
Therefore, at long times $t\gg\Lambda^{-1}$,
$K(t)\simeq\frac{\rho_{0}}{t},\qquad(t\gg\Lambda^{-1}),$ (S72)
which is the universal long-memory tail discussed in the main text.
(VI) Frequency-space form and connection to the retarded self-energy: For completeness we relate the time-domain memory kernel to the standard frequency-space self-energy. Taking the Fourier transform of the reservoir solution (S59) in the stationary (long-time) limit yields
$X_{\lambda}(\omega)=\frac{g(\lambda)}{\lambda-i\omega}\,\phi(\omega)+\frac{1}{\lambda-i\omega}\,\zeta_{\lambda}(\omega),$ (S73)
and substituting into (S54) gives
$\Big[-i\omega+r+\Sigma^{R}(\omega)\Big]\phi(\omega)=\eta(\omega)+\int_{0}^{\Lambda}d\lambda\,\frac{g(\lambda)}{\lambda-i\omega}\zeta_{\lambda}(\omega),$ (S74)
with the retarded self-energy
$\Sigma^{R}(\omega)=\int_{0}^{\Lambda}d\lambda\,\frac{g(\lambda)^{2}}{\lambda-i\omega}=\int_{0}^{\Lambda}d\lambda\,\frac{\rho(\lambda)}{\lambda-i\omega}.$ (S75)
(We use the convention $G_{R}^{-1}(\omega)=-i\omega+r+\Sigma^{R}(\omega)$; other sign conventions for the self-energy appear in the literature.)
Separating real and imaginary parts using
$\frac{1}{\lambda-i\omega}=\frac{\lambda}{\lambda^{2}+\omega^{2}}+i\,\frac{\omega}{\lambda^{2}+\omega^{2}},$ (S76)
one obtains
$\mathrm{Re}\,\Sigma^{R}(\omega)$ $=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{\lambda}{\lambda^{2}+\omega^{2}},$ (S77)
$\mathrm{Im}\,\Sigma^{R}(\omega)$ $=\omega\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{1}{\lambda^{2}+\omega^{2}}.$ (S78)
For a flat TDOS $\rho(\lambda)\simeq\rho_{0}$ and $|\omega|\ll\Lambda$,
$\mathrm{Im}\,\Sigma^{R}(\omega)$ $\simeq\omega\rho_{0}\int_{0}^{\Lambda}\frac{d\lambda}{\lambda^{2}+\omega^{2}}=\omega\rho_{0}\left[\frac{1}{|\omega|}\arctan\!\Big(\frac{\Lambda}{|\omega|}\Big)\right]$
$\xrightarrow{|\omega|\ll\Lambda}\frac{\pi}{2}\rho_{0}\,\mathrm{sgn}(\omega),$ (S79)
and
$\mathrm{Re}\,\Sigma^{R}(\omega)$ $\simeq\rho_{0}\int_{0}^{\Lambda}d\lambda\,\frac{\lambda}{\lambda^{2}+\omega^{2}}=\frac{\rho_{0}}{2}\int_{0}^{\Lambda}d\lambda\,\frac{2\lambda}{\lambda^{2}+\omega^{2}}$
$=\frac{\rho_{0}}{2}\left[\ln(\lambda^{2}+\omega^{2})\right]_{\lambda=0}^{\lambda=\Lambda}$
$=\frac{\rho_{0}}{2}\Big(\ln(\Lambda^{2}+\omega^{2})-\ln(\omega^{2})\Big)$
$=\frac{\rho_{0}}{2}\ln\!\left(\frac{\Lambda^{2}+\omega^{2}}{\omega^{2}}\right)\simeq\rho_{0}\ln\!\left(\frac{\Lambda}{|\omega|}\right),$
$\qquad(|\omega|\ll\Lambda),$ (S80)
which reproduces the logarithmic reactive singularity discussed in the main text.
Summary. Equations (S64)–(S68) show that integrating out a continuum of OU relaxation modes yields an exact generalized Langevin equation for the collective coordinate with a causal memory kernel $K(t)$. The kernel is the Laplace transform of the TDOS. A finite TDOS at the origin produces the universal long-memory tail $K(t)\sim 1/t$, and the corresponding retarded self-energy exhibits the marginal nonanalytic structure $\mathrm{Im}\,\Sigma^{R}(\omega)\propto\mathrm{sgn}(\omega)$ together with a logarithmic real part. This makes explicit that memory-dominated infrared dynamics arises from the self-organized accumulation of slow internal relaxation modes, rather than from coupling to a fast Markovian bath.
(VII) Conceptual structure: from hidden modes to infrared memory: The derivation above establishes a general structural result. A Markovian dynamical system defined in an extended state space $\{\phi,X_{\lambda}\}$ becomes non-Markovian when the internal degrees of freedom $X_{\lambda}$ are eliminated. The resulting equation for $\phi$ necessarily acquires a causal memory kernel given exactly by the Laplace transform of the time-scale density of states (TDOS),
$K(t)=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)e^{-\lambda t}.$ (S81)
This relation is purely kinematic and does not rely on perturbation theory, weak coupling, or specific microscopic assumptions. The infrared dynamical structure of the collective field is therefore completely determined by the low-$\lambda$ scaling of $\rho(\lambda)$.
If the TDOS vanishes at the origin, $\rho(\lambda)\to 0$ as $\lambda\to 0$, the kernel decays exponentially and the dynamics reduces to effectively Markovian overdamping. By contrast, when the TDOS remains finite at zero relaxation rate,
$\rho(\lambda\to 0)=\rho_{0}\neq 0,$ (S82)
the Laplace transform produces the universal long-time tail
$K(t)\simeq\frac{\rho_{0}}{t},\qquad t\gg\Lambda^{-1}.$ (S83)
Thus long-memory infrared dynamics is not an additional assumption but an unavoidable consequence of a flat slow-mode spectrum. The TDOS provides the fundamental organizing principle linking microscopic stability structure to macroscopic temporal response.
In this sense, the emergence of memory-dominated criticality reflects a dynamical phase characterized by an extensive accumulation of near-marginal relaxation modes rather than by coupling to a fast bath.
Appendix C Appendix C: Renormalization-group flow of the Cooper channel in conventional and memory-dominated regimes
In this Appendix we derive explicitly the infrared renormalization-group (RG) flow equations for the effective pairing interaction in the Cooper channel. We contrast the conventional Eliashberg regime with the memory-dominated regime generated by a flat relaxation-rate density of states. The derivation follows from shell integration of the linearized gap equation and clarifies the origin of marginal versus relevant superconducting instabilities.
Fermionic Matsubara frequencies are denoted by $\omega$ and $\omega^{\prime}$, and $\Omega\equiv\omega-\omega^{\prime}$ is the bosonic transfer frequency entering the pairing kernel $g(\Omega)$.
(I) Linearized gap equation and RG shell construction: At low energies the linearized gap equation takes the schematic form
$\Delta(\omega)=\int_{T}^{\Lambda}\frac{d\omega^{\prime}}{|\omega^{\prime}|}\,g(\omega-\omega^{\prime})\,\Delta(\omega^{\prime}),$ (S84)
where $\Lambda$ is the ultraviolet cutoff and $T$ provides the infrared cutoff. The factor $1/|\omega^{\prime}|$ originates from the fermionic Cooper bubble. The effective pairing kernel $g(\Omega)$ is generated by collective fluctuations.
The RG procedure consists of integrating out a thin frequency shell $\omega^{\prime}\in[\Lambda/b,\Lambda]$ with $b>1$ and absorbing its contribution into a renormalized coupling $g\rightarrow g+dg$. Defining the RG scale $l=\ln b$, the flow equation follows from the shell contribution.
(II) Conventional Eliashberg regime: Near metallic quantum criticality with Ohmic Landau damping, the bosonic propagator obeys
$D(\Omega)\sim\frac{1}{|\Omega|},$ (S85)
so that the pairing kernel scales as
$g(\Omega)\sim g\,D(\Omega)\sim\frac{g}{|\Omega|},$ (S86)
where $g$ denotes the dimensionless pairing interaction.
The leading shell correction arises from the ladder diagram and is proportional to
$dg\propto g^{2}\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{|\omega^{\prime}|}.$ (S87)
The shell integral is elementary:
$\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{\omega^{\prime}}=\ln\Lambda-\ln(\Lambda/b)=\ln b\equiv dl.$ (S88)
Hence the RG equation takes the marginal form
$\frac{dg}{dl}=C\,g^{2},$ (S89)
with $C$ a positive numerical constant.
The solution is
$g(l)=\frac{g_{0}}{1-Cg_{0}l}\simeq g_{0}+Cg_{0}^{2}l,$ (S90)
demonstrating logarithmic enhancement characteristic of a marginal superconducting instability.
(III) Memory-dominated regime: In the memory-dominated regime derived in Sec. II, the collective susceptibility behaves as
$\chi(\Omega)\sim\frac{1}{\rho_{0}\ln(\Lambda/|\Omega|)},$ (S91)
reflecting the flat TDOS of slow relaxation modes.
The pairing kernel therefore scales as
$g(\Omega)\sim\frac{g}{\ln(\Lambda/|\Omega|)}.$ (S92)
The shell correction becomes
$dg\propto g^{2}\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{|\omega^{\prime}|}\frac{1}{\ln(\Lambda/\omega^{\prime})}.$ (S93)
Introducing the variable
$u=\ln\frac{\Lambda}{\omega^{\prime}},\qquad du=-\frac{d\omega^{\prime}}{\omega^{\prime}},$ (S94)
the integration limits transform as
$\omega^{\prime}=\Lambda\Rightarrow u=0,\qquad\omega^{\prime}=\Lambda/b\Rightarrow u=\ln b\equiv dl.$ (S95)
Equation (S93) becomes
$dg\propto g^{2}\int_{0}^{dl}\frac{du}{u}.$ (S96)
The apparent divergence at $u\to 0$ is cut off at a microscopic scale $u_{\min}=O(1)$ marking the onset of the logarithmic TDOS regime. The shell contribution reduces to
$dg=\alpha\,g\,dl,$ (S97)
where $\alpha>0$ depends on the spectral weight $\rho_{0}$ and the onset scale of the memory-dominated regime.
The RG equation is therefore
$\frac{dg}{dl}=\alpha\,g,$ (S98)
with solution
$g(l)=g_{0}e^{\alpha l}=g_{0}b^{\alpha}.$ (S99)
The pairing interaction thus grows as a power law under infrared coarse graining, signaling a relevant superconducting instability.
(IV) Physical origin of marginal versus relevant behavior: The qualitative difference between Eqs. (S89) and (S98) originates from the infrared structure of the pairing kernel.
In the Eliashberg regime, each logarithmic shell contributes independently, leading to cumulative logarithmic growth. In contrast, the memory-dominated kernel generated by a flat TDOS correlates fluctuations across all time scales, producing linear amplification of the effective interaction under RG.
Consequently, even moderate intrinsic electronic pairing scales are dynamically amplified into a runaway superconducting instability once a slow-mode reservoir is formed.
Summary. Ohmic Landau damping yields a marginal Cooper-channel flow $dg/dl\propto g^{2}$, producing only logarithmic enhancement. A flat relaxation-rate TDOS instead generates a memory-dominated kernel that renders the pairing interaction relevant, with $dg/dl\propto g$ and power-law growth. This dynamical reorganization of infrared spectral weight underlies the algebraic superconducting instability central to the present theory.
Appendix D Appendix D: Finite-temperature regulation of the infrared divergence
In this Appendix we explicitly demonstrate how the zero–temperature infrared divergence of the polarization kernel, $\Pi(\omega)\sim 1/|\omega|$, is regulated at finite temperature, yielding the universal scaling $\Pi(T)\sim 1/T$. This establishes rigorously that the thermal scale acts as the infrared cutoff in the memory–dominated critical regime.
(I) Spectral integral representation: In the slow–mode reservoir regime the Cooper-channel polarization admits the spectral form
$\Pi(\omega)=\int_{0}^{\Lambda}d\lambda\;\frac{\rho_{\rm eff}(\lambda)}{\lambda^{2}+\omega^{2}}.$ (S100)
[…]
$\Pi(\omega)\simeq\rho_{0}\int_{0}^{\Lambda}d\lambda\;\frac{1}{\lambda^{2}+\omega^{2}}.$ (S101)
The elementary integral yields
$\int_{0}^{\Lambda}\frac{d\lambda}{\lambda^{2}+\omega^{2}}=\frac{1}{|\omega|}\tan^{-1}\!\left(\frac{\Lambda}{|\omega|}\right).$ (S102)
In the low–frequency limit $|\omega|\ll\Lambda$, one obtains
$\Pi(\omega)\simeq\frac{\pi\rho_{0}}{2}\,\frac{1}{|\omega|},$ (S103)
which expresses the nonanalytic infrared enhancement at zero temperature.
(II) Thermal infrared cutoff: At finite temperature the imaginary–time direction is compactified to the interval $\tau\in[0,\beta]$ with $\beta=1/T$, and dynamical correlations are expressed in terms of discrete Matsubara frequencies. As a consequence, continuous access to arbitrarily small frequencies is lost, and the thermal scale provides an effective infrared regulator.
Operationally, the zero–temperature denominator $\omega^{2}$ in Eq. (S101) is replaced by the minimal dynamical scale of order $T^{2}$, yielding
$\Pi(T)\simeq\rho_{0}\int_{0}^{\Lambda}d\lambda\;\frac{1}{\lambda^{2}+(cT)^{2}},$ (S104)
where $c$ is a numerical constant of order unity determined by the Matsubara spectrum.
Evaluating the integral gives
$\Pi(T)=\rho_{0}\,\frac{1}{cT}\tan^{-1}\!\left(\frac{\Lambda}{cT}\right)\simeq\frac{\pi\rho_{0}}{2c}\,\frac{1}{T},\qquad T\ll\Lambda.$ (S105)
Thus the infrared divergence is universally cut off by the thermal scale,
$\Pi(T)\propto\frac{1}{T}.$ (S106)
(III) Equivalent Matsubara summation: The same result follows directly from the finite–temperature Matsubara representation of the polarization bubble,
$\Pi(0,T)\sim T\sum_{\omega_{m}}\int_{0}^{\Lambda}d\lambda\;\frac{\rho_{0}}{\lambda^{2}+\omega_{m}^{2}},\qquad\omega_{m}=(2m+1)\pi T.$ (S107)
Using the standard identity
$T\sum_{\omega_{m}}\frac{1}{\lambda^{2}+\omega_{m}^{2}}=\frac{1}{2\lambda}\tanh\!\left(\frac{\lambda}{2T}\right),$ (S108)
and expanding for $\lambda\ll T$,
$\tanh\!\left(\frac{\lambda}{2T}\right)\simeq\frac{\lambda}{2T},$ (S109)
one obtains
$T\sum_{\omega_{m}}\frac{1}{\lambda^{2}+\omega_{m}^{2}}\simeq\frac{1}{2T}.$ (S110)
Substitution into Eq. (S107) yields
$\Pi(0,T)\sim\int_{0}^{\Lambda}d\lambda\;\rho_{0}\frac{1}{2T}\propto\frac{\rho_{0}}{T},$ (S111)
confirming that finite temperature universally regulates the infrared divergence.
(IV) Implication for the superconducting transition: The Thouless criterion,
$\chi_{\Delta}^{-1}(0,T_{c})=0\qquad\Longleftrightarrow\qquad\frac{1}{g_{\rm pair}}=\Pi(T_{c}),$ (S112)
therefore gives
$T_{c}\sim\rho_{0}\,g_{\rm pair},$ (S113)
or equivalently $T_{c}\sim\rho_{0}E_{\rm pair}$ in terms of the bare electronic pairing scale. The exponential suppression characteristic of BCS theory is thus replaced by algebraic scaling controlled by the infrared spectral weight of the slow–mode reservoir.
Appendix E Appendix E: Spectral function generated by memory–dominated self–energy
In this Appendix we derive the real–frequency spectral function associated with the collective retarded susceptibility in the memory–dominated critical regime. We show that a flat slow–mode TDOS produces a broad non–Lorentzian continuum with logarithmic infrared scaling and a universal collapse form.
(I) Retarded susceptibility in the memory–dominated regime: In the infrared regime controlled by a flat TDOS, the retarded inverse susceptibility takes the universal form
$\chi_{R}^{-1}(\omega,\mathbf{q})=\Delta_{\mathbf{q}}+\rho_{0}\ln\!\left(\frac{\Lambda}{|\omega|}\right)+i\Gamma\,\mathrm{sgn}(\omega),$ (S114)
where
$\Delta_{\mathbf{q}}\equiv r+cq^{2},\qquad\Gamma=\frac{\pi}{2}\rho_{0},$ (S115)
and $\Lambda$ is the ultraviolet cutoff of the slow–mode reservoir. Accordingly,
$\chi_{R}(\omega,\mathbf{q})=\frac{1}{\Delta_{\mathbf{q}}+\rho_{0}\ln(\Lambda/|\omega|)+i\Gamma\,\mathrm{sgn}(\omega)}.$ (S116)
(Throughout we use $\omega$ for real frequency; Matsubara frequencies are denoted separately in Appendix D.)
(II) Definition of the spectral function: The spectral function of collective fluctuations is defined by
$A(\omega,\mathbf{q})\equiv-2\,\mathrm{Im}\,\chi_{R}(\omega,\mathbf{q}).$ (S117)
Writing $\chi_{R}^{-1}=X(\omega,\mathbf{q})+iY(\omega)$ with
$X(\omega,\mathbf{q})$ $=\Delta_{\mathbf{q}}+\rho_{0}\ln\!\left(\frac{\Lambda}{|\omega|}\right),$ (S118)
$Y(\omega)$ $=\Gamma\,\mathrm{sgn}(\omega),$ (S119)
one has $\chi_{R}=(X-iY)/(X^{2}+Y^{2})$, hence
$A(\omega,\mathbf{q})=\frac{2\Gamma}{\left[\Delta_{\mathbf{q}}+\rho_{0}\ln(\Lambda/|\omega|)\right]^{2}+\Gamma^{2}}.$ (S120)
(III) Infrared scaling at criticality: Along the critical manifold $\Delta_{\mathbf{q}}=0$,
$A(\omega,\mathbf{q}=0)=\frac{2\Gamma}{\rho_{0}^{2}\ln^{2}(\Lambda/|\omega|)+\Gamma^{2}}.$ (S121)
For $|\omega|\ll\Lambda$ the logarithm dominates and the universal infrared tail is
$A(\omega)\xrightarrow{|\omega|\to 0}\frac{2\Gamma}{\rho_{0}^{2}}\,\frac{1}{\ln^{2}(\Lambda/|\omega|)}.$ (S122)
Thus the spectral weight is suppressed only logarithmically toward zero frequency, reflecting the accumulation of memory across many time scales.
(IV) Peak / hump scale away from criticality: For finite $\Delta_{\mathbf{q}}$, Eq. (S120) is maximized when
$\Delta_{\mathbf{q}}+\rho_{0}\ln\!\left(\frac{\Lambda}{|\omega_{*}|}\right)=0,$ (S123)
which yields
$|\omega_{*}|=\Lambda\exp\!\left(\frac{\Delta_{\mathbf{q}}}{\rho_{0}}\right).$ (S124)
For $\Delta_{\mathbf{q}}>0$ this scale lies above the infrared window ($|\omega_{*}|\gtrsim\Lambda$) and no low–energy peak develops, whereas for $\Delta_{\mathbf{q}}<0$ a pronounced low–frequency hump appears at $|\omega_{*}|\ll\Lambda$.
(V) Universal scaling collapse: Define the scaling variable
$u(\omega,\mathbf{q})\equiv\frac{\Delta_{\mathbf{q}}}{\rho_{0}}+\ln\!\left(\frac{\Lambda}{|\omega|}\right).$ (S125)
Using $\Gamma=(\pi/2)\rho_{0}$, Eq. (S120) reduces to the universal form
$A(\omega,\mathbf{q})=\frac{\pi}{\rho_{0}}\,\frac{1}{u(\omega,\mathbf{q})^{2}+(\pi/2)^{2}},$ (S126)
so that all momentum and tuning dependence enters only through $u$, implying a collapse of line shapes onto a single curve.
(VI) Finite–temperature infrared cutoff: At finite temperature the logarithmic growth is regulated by the thermal scale. A minimal phenomenological implementation consistent with Appendix D is
$|\omega|\;\rightarrow\;\max\{|\omega|,cT\},\qquad c=O(1),$ (S127)
yielding
$A_{T}(\omega,\mathbf{q})=\frac{2\Gamma}{\left[\Delta_{\mathbf{q}}+\rho_{0}\ln(\Lambda/\max\{|\omega|,cT\})\right]^{2}+\Gamma^{2}}.$ (S128)
For $|\omega|\ll T$ the logarithm saturates, producing a finite infrared spectral weight rather than the zero–temperature logarithmic suppression.
(VII) Physical interpretation: Unlike conventional overdamped criticality, where $\mathrm{Im}\,\Sigma^{R}(\omega)$ vanishes linearly with $|\omega|$, the memory–dominated regime exhibits frequency–independent damping ($\propto\mathrm{sgn}(\omega)$) accompanied by a logarithmically singular reactive part. The resulting spectral function is therefore a broad continuum with slowly varying infrared tails rather than a sharp Lorentzian quasiparticle peak.
Equation (S120) (or equivalently the collapse form (S126)) provides a directly testable dynamical signature of the slow–mode reservoir and the associated TDOS universality class.
Appendix F Appendix F: Relaxation-rate spectra as the fundamental dynamical basis of collective dynamics
In conventional quantum many-body theory, collective behavior is typically organized in terms of energy eigenmodes of an underlying Hamiltonian,
$H|n\rangle=E_{n}|n\rangle,$ (S129)
with dynamical response expressed through Fourier modes oscillating at frequencies set by energy differences. This structure naturally emphasizes stationary states and conservative dynamics, with dissipation treated as a secondary perturbative effect.
However, this description represents a special limit corresponding to closed, Hamiltonian systems. In general interacting and open systems — particularly near criticality — collective dynamics is governed not by energy eigenmodes but by the stability spectrum of the full dynamical flow.
(I) Dynamical linearization and relaxation modes: For a general high-dimensional dynamical system
$\dot{x}(t)=F(x(t)),$ (S130)
linearization about a metastable configuration $x^{*}$ yields
$\delta\dot{x}(t)=-A\,\delta x(t),\qquad A=-\left.\frac{\partial F}{\partial x}\right|_{x^{*}}.$ (S131)
The eigenvalue problem
$Av_{i}=\lambda_{i}v_{i}$ (S132)
defines a spectrum of collective relaxation modes with temporal evolution
$\delta x_{i}(t)\sim e^{-\lambda_{i}t}.$ (S133)
The real parts of $\lambda_{i}$ set intrinsic time scales of decay, while imaginary parts correspond to oscillatory components. Energy eigenmodes arise only in the special Hamiltonian limit where all eigenvalues lie purely on the imaginary axis.
Thus, the relaxation-rate spectrum $\{\lambda_{i}\}$ constitutes the natural and generally applicable dynamical basis for collective phenomena.
(II) Liouvillian formulation and relation to energy spectra: In quantum dynamical field theories formulated in Keldysh or MSRJD language, evolution of the density matrix is generated by a Liouvillian superoperator $\mathcal{L}$,
$\frac{d\rho}{dt}=\mathcal{L}\rho.$ (S134)
For closed Hamiltonian dynamics,
$\mathcal{L}\rho=-i[H,\rho],$ (S135)
yielding eigenvalues
$\lambda_{nm}=-i(E_{n}-E_{m}),$ (S136)
which lie purely on the imaginary axis and reproduce energy-difference oscillations.
In interacting open systems and critical regimes, additional dissipative and collective processes shift Liouvillian eigenvalues into the complex plane, producing a dense spectrum of decay rates. The fundamental objects governing long-time dynamics are therefore relaxation rates rather than energy levels.
(III) Memory kernels and continuous relaxation spectra: General linear causal dynamics with memory takes the form
$\dot{\phi}(t)+\int_{0}^{t}ds\,K(t-s)\phi(s)=\eta(t).$ (S137)
Taking the Laplace transform gives
$\tilde{\phi}(s)=\frac{\tilde{\eta}(s)}{s+\tilde{K}(s)}.$ (S138)
The poles and branch cuts of $s+\tilde{K}(s)$ define the intrinsic relaxation spectrum. When $\tilde{K}(s)$ possesses branch-cut structure, the response is governed by a continuous distribution of decay rates,
$K(t)=\int_{0}^{\infty}d\lambda\,\rho(\lambda)e^{-\lambda t},$ (S139)
which uniquely defines the time-scale density of states.
Thus continuous relaxation spectra emerge generically from causality and memory, independent of microscopic details.
(IV) Energy modes as a special case: The familiar Fourier representation of conservative dynamics corresponds to the singular limit in which the TDOS collapses onto the imaginary axis,
$\rho(\lambda)\propto\delta(\mathrm{Re}\,\lambda),$ (S140)
recovering purely oscillatory energy eigenmodes. Away from this limit, generic interacting systems naturally develop broad relaxation spectra controlling infrared behavior.
(V) Physical implication: The time-scale density of states therefore represents the fundamental organizing variable of collective dynamics in correlated systems. Energy eigenmodes describe a special integrable limit, while relaxation-rate spectra capture the universal structure of open, critical, and strongly interacting matter.
Memory-dominated criticality corresponds precisely to the formation of an extensive continuum of near-marginal relaxation modes. The resulting TDOS controls dynamical response, pairing enhancement, and infrared universality, providing a deeper and more general foundation than energy-based descriptions.