### v2

# arXiv:2602.22626v2[cond-mat.str-el] 02 Mar 2026

Abstract
[…]
I I. Introduction
High–temperature superconductivity remains one of the central unsolved problems of strongly correlated quantum matter [1, 2]. Despite decades of intensive theoretical and experimental effort, no consensus has emerged on the microscopic mechanism responsible for robust pairing far above conventional phonon scales. A central difficulty lies in the intimate proximity of superconductivity to quantum critical regimes, where strong collective fluctuations dominate low–energy dynamics.
Within the prevailing theoretical paradigm, quantum critical fluctuations are assumed to be overdamped by fast electronic continua [3, 4]. This framework, formalized in Hertz–Millis theory, predicts a universal Ohmic Landau damping kernel of the form $|\omega|$ governing collective dynamics. While remarkably successful in weakly correlated metals, this assumption has increasingly come into tension with experiments in strongly correlated systems, including cuprate superconductors and strange metals, where anomalous temporal correlations, non-Markovian response, and $1/f$-like noise are ubiquitously observed [5, 6, 7, 8, 9].
These observations raise a fundamental question: Is overdamped Markovian critical dynamics truly universal, or does it rely on implicit assumptions about the structure of dynamical spectra that break down in strongly correlated regimes?
A key limitation of conventional approaches is their formulation in terms of a small number of collective order-parameter fields [10, 11]. Such descriptions implicitly assume that the infrared behavior can be captured by a finite set of soft bosonic modes with local-in-time damping. However, strongly correlated quantum matter is fundamentally governed by Liouvillian evolution of the density matrix, where the full spectrum of relaxation eigenvalues determines long-time dynamics [12, 13, 14]. In more general, criticality need not correspond to softening of a single excitation mode but may instead involve reorganization of the entire relaxation-rate spectrum.
Rather than reducing the dynamics to a few collective coordinates, one must consider the possibility that the Liouvillian spectrum develops an extended continuum of near-marginal decay modes. Such spectral accumulation naturally generates a broad distribution of relaxation time scales and invalidates the assumption of Markovian coarse-graining near criticality. The relevant organizing structure of the infrared theory is then not a small set of overdamped bosons, but the time–scale density of states (TDOS) of collective relaxation modes.
In this work we show that infrared collective dynamics is generically controlled by the TDOS of correlated internal decay channels. By formulating a dynamical coarse–graining theory within the Martin–Siggia–Rose–Janssen–De Dominicis (MSRJD) framework [15, 16, 17], we derive an exact spectral representation of the collective susceptibility in terms of the relaxation–rate density, without assuming a priori a finite set of soft modes.
We demonstrate that a TDOS scaling $\rho(\lambda)\sim\lambda^{\alpha}$ directly classifies universality classes of quantum critical dynamics. While the conventional Hertz–Millis kernel emerges only for $\alpha=1$, a finite TDOS at the origin ($\alpha=0$) produces a marginal nonanalytic retarded kernel proportional to $\mathrm{sgn}(\omega)$, corresponding in the time domain to a universal long–memory response $K(t)\sim 1/t$. This memory–dominated infrared dynamics represents a breakdown of Markovian overdamped criticality and defines a distinct universality class governed by the structure of the relaxation spectrum.
Crucially, we show that such flat relaxation-rate spectra arise naturally in strongly correlated collective systems, including dynamical manifolds explored in our previous work [18, 19, 20]. The resulting long-memory kernel dynamically amplifies intrinsic electronic pairing tendencies, providing a route to strong superconductivity without invoking finely tuned bosonic mediators.
Figure 1 summarizes the resulting dynamical universality map of correlated quantum matter. The infrared structure of the relaxation spectrum determines collective kernels in the time domain and controls physical consequences including pairing enhancement, superconducting domes, Uemura scaling, and anomalous dynamical response [1, 21, 22]. Charge, spin, and pairing channels reorganize within a common slow-mode reservoir, with superconductivity corresponding to condensation in the particle–particle sector of a shared infrared spectral continuum.
Our results establish the relaxation-rate spectrum as a fundamental organizing principle of quantum critical dynamics and reveal memory–dominated criticality as a generic route toward high–temperature superconductivity. This framework reconciles anomalous dynamical phenomena observed in strange metals with robust pairing and provides a unified spectral perspective on collective behavior across strongly correlated systems [5, 7, 10].
The remainder of this paper is organized as follows. In Section II we derive the exact TDOS representation of the collective susceptibility and classify infrared universality classes. In Sections III and IV we analyze the resulting memory–dominated pairing kernel and its implications for superconducting instabilities. In Section V we discuss experimental consequences including superconducting domes, Uemura scaling, and dynamical anomalies. We conclude in Section VI with a summary and outlook.
Figure 1: Dynamical universality map of correlated quantum matter. (a) Infrared time–scale density of states $\rho(\lambda)$ defining distinct dynamical universality classes. Vanishing TDOS yields conventional Hertz–Millis and phonon dynamics, marginal spectra produce Ohmic damping, while a flat TDOS defines memory–dominated criticality. (b) Corresponding long–time kernels in the time domain, ranging from Markovian dissipation to non-Markovian power–law memory. (c) Physical consequences of memory–dominated criticality, including algebraic pairing enhancement, superconducting domes, Uemura scaling, strange metal transport, and long–time noise correlations.
II II. Emergence of a Slow-Mode Reservoir from Correlated Dynamical Spectra
Strongly correlated electronic matter is intrinsically a quantum many-body system. At long times and coarse-grained scales, its dynamics is governed not by isolated quasiparticles, but by collective relaxation channels generated by the full quantum dynamical evolution of the density matrix [12, 13, 14],
$\frac{d\hat{\varrho}}{dt}=\mathcal{L}\hat{\varrho}$, (1)
where $\mathcal{L}$ denotes the Liouvillian generator of the effective (open or coarse-grained) dynamics.
The eigenmodes of $\mathcal{L}$,
$\mathcal{L}R_{n}=\lambda_{n}R_{n},$ (2)
define the fundamental decay channels of the interacting system. For stable dynamics, $\mathrm{Re}\,\lambda_{n}\leq 0$, and we define the collective relaxation rates as
$\lambda\equiv-\mathrm{Re}\,\lambda_{n}\geq 0.$ (3)
Infrared dynamics is therefore organized by the distribution of these decay scales rather than by Hamiltonian excitation energies.
Observable collective responses—charge, spin, pairing, current, and phase-coherence fluctuations—arise from projections of the full relaxation spectrum onto experimentally accessible operators. The resulting low-energy theory is thus controlled by the statistical structure of relaxation rates. We define the TDOS as the infrared distribution of Liouvillian relaxation eigenvalues,
$\rho(\lambda)=\sum_{n}\delta\!\left(\lambda+\mathrm{Re}\,\lambda_{n}\right),$ (4)
or its continuum limit.
The spectral structure introduced here should therefore be understood at the level of the full dynamical generator rather than solely in terms of Hamiltonian energy eigenmodes. Energy spectra characterize the conservative limit in which all Liouvillian eigenvalues lie on the imaginary axis, whereas generic interacting or coarse-grained systems develop complex spectra with nonzero decay rates. A detailed discussion of this relaxation-rate (Liouvillian) formulation and its connection to memory kernels is provided in Appendix A.
II.1 A. Dynamical stability spectrum and time-scale density of states
The infrared structure of correlated matter is encoded in the spectrum of collective relaxation rates. Rather than softening a single order parameter, correlated critical regimes reorganize the entire decay spectrum, producing a dense accumulation of near-marginal modes.
We characterize this structure by introducing the TDOS, which measures the density of collective decay channels at relaxation rate $\lambda$. This object plays a role analogous to the energy density of states, but in relaxation-rate space. It determines how many collective modes decay on a given time scale.
In weakly interacting systems, the number of slow modes remains finite, and the TDOS vanishes as $\lambda\to 0$, reflecting the presence of well-defined quasiparticles and isolated collective excitations.
By contrast, in strongly correlated critical regimes, the relaxation spectrum generically develops an extensive accumulation of near-marginal decay channels,
$\rho(\lambda\to 0)=\rho_{0}\neq 0,$ (5)
forming a slow-mode reservoir that dominates infrared collective dynamics.
This accumulation signifies a dynamical approach to marginal stability along a macroscopic manifold of collective directions, rather than along a single order-parameter coordinate.
II.2 B. Functional representation of the relaxation spectrum
To compute collective response functions, we employ a functional representation of the dissipative dynamics. The dissipative quantum dynamics of interacting electrons may be formulated in the real–time Keldysh functional integral framework [23, 24]. Upon projection onto slow collective operators and in the regime where decoherence suppresses quantum interference between forward and backward contours, the theory reduces to a classical MSRJD functional integral describing the infrared stochastic dynamics.
Let $\phi(\mathbf{r},t)$ denote a coarse–grained collective bosonic field representing charge, spin, pairing, or phase fluctuations. Projecting the full quantum dynamics onto this operator sector yields an exact decomposition in terms of relaxation eigenmodes.
Because the dissipative spectrum consists of decay rates $\lambda\geq 0$, the projected dynamics may be represented as a continuum of relaxational modes labeled by their decay scales,
$\dot{\phi}$ $=-(r+c\nabla^{2})\phi+\int_{0}^{\Lambda}d\lambda\,g(\lambda)X_{\lambda}+\eta,$ (6)
$\dot{X}_{\lambda}$ $=-\lambda X_{\lambda}+g(\lambda)\phi+\zeta_{\lambda},$ (7)
where $X_{\lambda}$ represents a decay mode with relaxation rate $\lambda$, and $g(\lambda)$ encodes its overlap with $\phi$.
Introducing response fields $\tilde{\phi}$ and $\tilde{X}_{\lambda}$, the MSRJD action reads
$S=$ $\int dt\,d^{d}r\,\tilde{\phi}\Big[\dot{\phi}+(r+c\nabla^{2})\phi-\int_{0}^{\Lambda}d\lambda\,g(\lambda)X_{\lambda}\Big]-D_{\phi}\tilde{\phi}^{2}$
$+\int dt\,d^{d}r\int_{0}^{\Lambda}d\lambda\,\tilde{X}_{\lambda}\big[\dot{X}_{\lambda}+\lambda X_{\lambda}-g(\lambda)\phi\big]$
$-\int dt\,d^{d}r\int_{0}^{\Lambda}d\lambda\,D_{\lambda}\tilde{X}_{\lambda}^{2}.$ (8)
Because the reservoir fields $X_{\lambda}$ enter quadratically, they can be integrated out exactly. Performing the Gaussian functional integration generates a nonlocal-in-time memory kernel in the effective equation of motion,
$\partial_{t}\phi(t)=-(r+c\nabla^{2})\phi(t)-\int_{0}^{t}dt^{\prime}\,K(t-t^{\prime})\,\phi(t^{\prime})+\eta_{\rm eff}(t),$ (9)
where the memory kernel is given exactly by
$K(t)=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\,e^{-\lambda t},$ (10)
with the time-scale density of states
$\rho(\lambda)\equiv g(\lambda)^{2}.$ (11)
Equation (10) makes explicit that the slow-mode reservoir is entirely controlled by the distribution of relaxation rates. The long-time structure of $K(t)$ therefore encodes the infrared organization of decay modes in the interacting system.
Fourier transforming the kernel yields the retarded self-energy,
$\Sigma^{R}(\omega)=\int_{0}^{\Lambda}d\lambda\frac{\rho(\lambda)}{\lambda-i\omega},$ (12)
so that the effective action becomes
$S_{\rm eff}=\int_{\omega,\mathbf{q}}\tilde{\phi}(-\omega,-\mathbf{q})\Big[-i\omega+r+cq^{2}+\Sigma^{R}(\omega)\Big]\phi(\omega,\mathbf{q})+\cdots.$ (13)
Separating real and imaginary parts,
$\mathrm{Re}\,\Sigma^{R}(\omega)$ $=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{\lambda}{\lambda^{2}+\omega^{2}},$ (14)
$\mathrm{Im}\,\Sigma^{R}(\omega)$ $=\omega\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{1}{\lambda^{2}+\omega^{2}}.$ (15)
The collective inverse susceptibility therefore assumes the exact spectral representation
$\chi_{R}^{-1}(\omega,\mathbf{q})$ $=r+cq^{2}+\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{\lambda}{\lambda^{2}+\omega^{2}}$ (16)
$\quad+i\omega\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{1}{\lambda^{2}+\omega^{2}}.$
Equation (16) provides an exact infrared response fully controlled by the relaxation-rate density. The slow-mode reservoir thus appears as a memory-dominated fixed point, whose long-time kernel structure governs the dynamical amplification mechanism discussed below.
The explicit Gaussian integration and derivation of the memory kernel are presented in Appendix B.
II.3 C. Infrared universality classes
[…]
$\rho(\lambda)\sim\lambda^{\alpha},\qquad\lambda\to 0,$ (17)
the dissipative part of the collective kernel follows directly from the spectral representation in Eq. (16),
$\mathrm{Im}\,\chi_{R}^{-1}(\omega)=\omega\int_{0}^{\Lambda}d\lambda\,\frac{\rho(\lambda)}{\lambda^{2}+\omega^{2}}\;\sim\;\mathrm{sgn}(\omega)\,|\omega|^{\alpha}.$ (18)
Here $\omega$ denotes the collective relaxation frequency arising from the projected Liouvillian dynamics, rather than a microscopic quasiparticle excitation energy.
It is therefore natural to write the long–wavelength dynamical response in the scaling form
$\chi^{-1}(\omega,\mathbf{q})=r+cq^{2}+\gamma|\omega|^{s},$ (19)
in direct analogy with the conventional Hertz–Millis representation [3, 4]. The TDOS scaling immediately implies the universal identification $s=\alpha$.
Under the renormalization–group rescaling $\mathbf{q}\to b\,\mathbf{q}$ and $\omega\to b^{z}\omega$, the spatial gradient term scales as $q^{2}\to b^{2}q^{2}$, while the dissipative term scales as $|\omega|^{s}\to b^{zs}|\omega|^{s}$. Requiring both contributions to remain of comparable order at criticality ($r=0$) fixes the dynamical exponent to
$z=\frac{2}{s}.$ (20)
In this framework the dissipation is generically non–Ohmic, $|\omega|^{s}$ with $s\neq 1$, reflecting the organization of the underlying relaxation spectrum. Sub–Ohmic memory ($s<1$) implies $z>2$, corresponding to critical slowing down stronger than ordinary diffusive dynamics, whereas super–Ohmic behavior ($s>1$) yields $z<2$. Infrared universality classes are thus classified directly by the low–energy structure of the relaxation-rate density.
In particular, when the TDOS remains finite at the origin,
$\rho(\lambda)\xrightarrow{\lambda\to 0}\rho_{0},$ (21)
the dissipative kernel simplifies to
$\mathrm{Im}\,\chi_{R}^{-1}(\omega)$ $=\omega\int_{0}^{\Lambda}d\lambda\,\frac{\rho_{0}}{\lambda^{2}+\omega^{2}}$
$\xrightarrow{|\omega|\ll\Lambda}\frac{\pi}{2}\rho_{0}\,\mathrm{sgn}(\omega),$ (22)
while the reactive part acquires a logarithmic infrared singularity,
$\mathrm{Re}\,\chi_{R}^{-1}(\omega)=\int_{0}^{\Lambda}d\lambda\,\rho_{0}\frac{\lambda}{\lambda^{2}+\omega^{2}}\simeq\rho_{0}\ln\frac{\Lambda}{|\omega|}.$ (23)
Collective infrared dynamics is therefore governed by the marginal non-Markovian kernel
$\chi_{R}^{-1}(\omega,\mathbf{q})\simeq r+cq^{2}+\rho_{0}\ln\frac{\Lambda}{|\omega|}+i\frac{\pi}{2}\rho_{0}\,\mathrm{sgn}(\omega)+\cdots,$ (24)
corresponding in the time domain to a universal long-memory response $K(t)\sim 1/t$, derived explicitly in Appendix C.
This flat-TDOS fixed point represents a breakdown of the conventional overdamped critical dynamics underlying Hertz–Millis theory. Rather than arising from a fast electronic bath coupled to a single soft order-parameter mode, collective relaxation is controlled by the self-organized accumulation of an extensive continuum of slow decay modes.
III III. Memory–Enhanced Pairing Beyond Eliashberg Theory
The emergence of a slow–mode reservoir reorganizes the infrared structure of the Cooper pairing channel. In conventional quantum critical superconductivity, collective fluctuations act as overdamped bosonic mediators whose Ohmic dynamics renders the pairing interaction marginal. Here we show that when collective dynamics is instead governed by a flat relaxation–rate density of states, long–time memory effects transform the Cooper channel into a relevant infrared instability.
III.1 A. Marginal pairing from Ohmic collective dynamics
Near metallic quantum critical points, fermions couple to overdamped collective modes with retarded propagator
$D_{R}^{-1}(\omega,\mathbf{q})=r+cq^{2}+\gamma|\omega|.$ (25)
The linear frequency dependence reflects Landau damping by the electronic continuum and produces a pairing kernel scaling as
$D(\omega)\sim\frac{1}{|\omega|}.$ (26)
The infrared shell correction to the dimensionless Cooper–channel coupling $g$ therefore reads
$dg\propto g^{2}\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{\omega^{\prime}}=g^{2}\ln b.$ (27)
Introducing the RG scale $l=\ln b$ yields
$\frac{dg}{dl}=Cg^{2},$ (28)
with $C>0$.
The pairing interaction is thus marginal and grows only logarithmically under infrared coarse graining. Physically, overdamped collective modes dissipate into the electronic continuum, continuously erasing temporal coherence of pairing fluctuations and preventing strong accumulation of superconducting correlations.
III.2 B. Infrared reorganization by slow collective modes
In the memory–dominated regime derived in Sec. II, collective dynamics is governed not by a single overdamped soft mode but by a continuum of slow relaxation modes characterized by the time–scale density of states $\rho(\lambda)$. After integrating out the relaxation reservoir, the retarded collective susceptibility takes the exact spectral form
$\chi_{R}^{-1}(\omega,\mathbf{q})=r+cq^{2}+\Sigma(\omega),\qquad\Sigma(\omega)=\int_{0}^{\Lambda}d\lambda\,\frac{\rho(\lambda)}{\lambda-i\omega}.$ (29)
For a flat TDOS in the infrared, $\rho(\lambda\to 0)=\rho_{0}$, the effective polarization kernel in the Cooper channel inherits the infrared scaling
$\Pi(\omega)\propto\int_{0}^{\Lambda}d\lambda\,\rho_{0}\frac{1}{\lambda^{2}+\omega^{2}}\sim\frac{1}{|\omega|},\qquad|\omega|\to 0.$ (30)
Unlike the conventional Ohmic case, this singular behavior does not originate from Landau damping by fast electronic continua. Instead, it emerges from the accumulation of an extensive continuum of near–marginal relaxation modes encoded in a flat TDOS. The resulting kernel corresponds in the time domain to a universal long–memory response $K(t)\sim 1/t$.
Although both Eliashberg theory and the present memory–dominated regime generate infrared kernels scaling as $1/|\omega|$, their physical origins are fundamentally distinct. In Eliashberg theory the singularity reflects overdamped dissipation arising from coupling to a fast electronic bath [25, 26]. Here, by contrast, the same infrared scaling results from spectral accumulation of slow collective modes, producing intrinsically non–Markovian dynamics.
This dynamical reorganization qualitatively alters the Cooper channel: instead of marginal logarithmic enhancement characteristic of overdamped criticality, pairing fluctuations accumulate coherently across widely separated time scales. The slow–mode reservoir therefore converts the Cooper channel into a relevant infrared interaction.
III.3 C. Relevant infrared amplification of the Cooper channel
The infrared shell correction in the memory–dominated regime becomes
$dg\propto g^{2}\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{\omega^{\prime}\ln(\Lambda/\omega^{\prime})}.$ (31)
Introducing $u=\ln(\Lambda/\omega^{\prime})$ yields
$dg\simeq\alpha g\,dl,$ (32)
with $\alpha>0$ proportional to the infrared TDOS weight.
The resulting RG flow reads
$\frac{dg}{dl}=\alpha g,$ (33)
with solution $g(l)=g_{0}e^{\alpha l}$.
The Cooper–channel interaction therefore grows algebraically under infrared coarse graining, signaling a relevant superconducting instability rather than a marginal one. This qualitative contrast with Eliashberg theory is illustrated in Fig. 2. A detailed shell RG derivation demonstrating the crossover from logarithmic to power-law pairing growth is presented in Appendix D.
Physically, the slow–mode reservoir retains dynamical correlations across widely separated time scales, allowing collective fluctuations to act repeatedly on electronic degrees of freedom and amplify intrinsic electronic pairing tendencies.
Superconductivity in strongly correlated systems thus emerges from infrared spectral organization of collective relaxation dynamics, rather than from finely tuned exchange of specific bosonic mediators.
Figure 2: Infrared renormalization-group flow of the Cooper-channel interaction. Shown schematically is the flow of the dimensionless pairing coupling $g$ under infrared coarse graining. Ohmic Landau damping of collective modes produces marginal logarithmic growth, $dg/dl\sim g^{2}$, characteristic of conventional Eliashberg theory (blue). In contrast, memory–dominated dynamics generated by a flat relaxation–rate density of states yields a relevant infrared instability, $dg/dl=\alpha g$, leading to algebraic amplification of the pairing interaction (orange). The slow–mode reservoir therefore dynamically enhances intrinsic electronic pairing tendencies rather than merely renormalizing a fixed microscopic coupling.
IV IV. Pairing from Memory–Dominated Critical Dynamics
The dynamical framework developed in the preceding sections establishes that strongly correlated matter generically self–organizes into a regime dominated by an extensive reservoir of slow collective modes. Rather than producing conventional overdamped fluctuations described by Ohmic Eliashberg dynamics, the accumulation of near–marginal relaxation rates generates long–time memory kernels and nonanalytic infrared response.
Superconductivity in this setting emerges as a dynamical instability of the memory–dominated collective spectrum. The Cooper channel is reorganized by the slow–mode reservoir into a relevant infrared interaction, fundamentally distinct from marginal pairing mediated by fast bosonic modes. Consequently, even moderate intrinsic electronic pairing scales are dynamically amplified into an algebraic superconducting instability.
In this section we demonstrate explicitly how microscopic electronic correlations seed pairing and how the slow–mode TDOS reorganizes the Cooper channel into a memory–enhanced critical regime, ultimately yielding algebraic scaling of the transition temperature and robust phase coherence.
IV.1 A. Intrinsic electronic pairing and dynamical formulation
In strongly correlated materials proximate to Mott insulating phases, short–range Coulomb interactions naturally generate local singlet correlations through exchange processes. Starting from the Hubbard Hamiltonian,
$H=-t\sum_{\langle ij\rangle\sigma}c_{i\sigma}^{\dagger}c_{j\sigma}+U\sum_{i}n_{i\uparrow}n_{i\downarrow},$ (34)
projection to the strong–coupling sector yields the superexchange interaction
$H_{J}=J\sum_{\langle ij\rangle}\Big(\mathbf{S}_{i}\cdot\mathbf{S}_{j}-\frac{1}{4}n_{i}n_{j}\Big),\qquad J=\frac{4t^{2}}{U},$ (35)
favoring local singlet formation.
More generally, correlated electronic systems possess an intrinsic short–range pairing scale $E_{\rm pair}$ generated purely by electronic interactions, independent of long–wavelength collective mediators. This intrinsic scale provides the microscopic seed for superconductivity.
Within the dynamical framework developed in Sec. II, the pairing susceptibility assumes the universal tt–matrix form
$\chi_{\Delta}^{-1}(\omega)=\frac{1}{g_{\rm pair}}-\Pi(\omega),$ (36)
with $g_{\rm pair}\sim E_{\rm pair}^{-1}$. Critical dynamics does not create attraction; it amplifies a preexisting electronic pairing tendency through infrared spectral reorganization.
IV.2 B. Memory–dominated Cooper kernel from flat TDOS
Because single–particle propagators admit a relaxational spectral representation governed by the slow–mode TDOS, the Cooper bubble reduces to a continuum integral over relaxation rates,
$\Pi(\omega)\sim\int_{0}^{\Lambda}d\lambda\,\rho_{\rm eff}(\lambda)\frac{1}{\lambda^{2}+\omega^{2}}.$ (37)
When the slow–mode reservoir develops,
$\rho_{\rm eff}(\lambda\to 0)=\rho_{0}\neq 0,$ (38)
the infrared behavior becomes
$\Pi(\omega)\xrightarrow{|\omega|\ll\Lambda}\frac{\pi\rho_{0}}{2}\,\frac{1}{|\omega|}.$ (39)
This nonanalytic enhancement reflects the accumulation of collective memory across widely separated time scales. Unlike conventional Eliashberg kernels, which are marginal in the infrared, the present Cooper channel becomes relevant due to the flat TDOS. The slow–mode reservoir therefore reorganizes the pairing problem into a memory–dominated critical regime.
IV.3 C. Algebraic Instability and Transition Scale
The superconducting transition follows from the Thouless criterion
$\chi_{\Delta}^{-1}(\omega\sim T_{c})=0,$ (40)
which yields
$T_{c}\sim\rho_{0}\,E_{\rm pair},$ (41)
up to numerical factors of order unity. The finite-temperature regularization of the infrared divergence, leading to $\Pi(T)\sim 1/T$, is derived explicitly in Appendix E.
The exponential suppression characteristic of BCS theory is thus replaced by algebraic scaling governed by the infrared spectral weight of the slow–mode reservoir. High transition temperatures arise generically from dynamical amplification of moderate intrinsic pairing scales.
It is important to emphasize that superconductivity requires not only pair formation but also global phase coherence. The transition temperature is therefore ultimately bounded by the smaller of the pairing amplitude and the superfluid phase stiffness,
$k_{B}T_{c}\sim\min\left(\Delta,\,\rho_{s}\right),$ (42)
with $\rho_{s}\sim n_{s}/m^{*}$. The detailed scaling of the phase stiffness within the slow–mode framework is discussed in Sec. V.
The key result of the present section is that the pairing scale itself is controlled by a single infrared parameter, the zero–frequency TDOS $\rho_{0}$. The slow–mode reservoir therefore acts as a universal dynamical amplifier: it converts intrinsic electronic correlation energy into an algebraically enhanced instability of the Cooper channel, independent of any specific bosonic mediator.
In this sense, the TDOS–controlled fixed point defines a dynamical stability class distinct from both weak–coupling BCS superconductors and conventional Eliashberg scenarios. The emergence of superconductivity is governed not by the exchange of a particular glue, but by infrared spectral organization of collective relaxation modes.
Figure 3: Experimental consequences of memory–dominated critical dynamics. (a) Superconducting dome emerging from truncation of the slow–mode reservoir away from criticality. The transition temperature follows $T_{c}(p)\sim E_{\rm pair}\rho_{0}(p)$ and is maximal near the correlated critical point $p_{c}$, where infrared spectral weight of collective relaxation modes is largest. Suppression of $T_{c}$ on both sides arises solely from the finite infrared extent of the slow–mode reservoir, without invoking competing orders or fine tuning. (b) Uemura scaling collapse between superconducting transition temperature and superfluid stiffness, $T_{c}\propto\rho_{s}$. Both pairing strength and phase coherence are controlled by the same infrared spectral weight $\rho_{0}$ of slow collective modes, yielding universal linear scaling across material families.
V V. Experimental consequences of memory–dominated critical dynamics
[…]
V.1 A. Infrared spectral control and automatic emergence of the dome
The linearized transition scale derived in Eq. (41) shows that superconductivity is governed by the infrared spectral weight $\rho_{0}$ of the slow–mode reservoir. We interpret this regime not as a finely tuned quantum critical point, but as a self–organized dynamical fixed point of the relaxation spectrum.
In this phase, feedback between electronic correlations and collective relaxation continuously regenerates an extensive near–marginal sector, leading to a scale–free infrared TDOS,
$\rho(\lambda\to 0)=\rho_{*}>0,$ (43)
without the need for fine tuning of microscopic parameters. The flat infrared spectrum therefore characterizes a dynamical phase rather than a measure–zero critical point.
External control parameters such as doping or pressure primarily modulate the effective infrared cutoff scale of this reservoir. Let $p$ denote a tuning parameter and $p_{\star}$ the point at which the infrared cutoff is minimal. The collective mass term behaves as
$r(p)\sim|p-p_{\star}|,$ (44)
so that the associated cutoff obeys
$\Lambda(p)\sim r^{z\nu}\sim|p-p_{\star}|^{z\nu}.$ (45)
Away from $p_{\star}$, the slow–mode reservoir remains present but is truncated in the deep infrared, yielding
$\rho_{0}(p)\sim\rho_{*}\,f\!\left(\frac{|p-p_{\star}|}{p_{0}}\right),$ (46)
with $f$ a decreasing crossover function.
Substituting into Eq. (41) gives
$T_{c}(p)\sim E_{\rm pair}\,\rho_{*}\,f\!\left(\frac{|p-p_{\star}|}{p_{0}}\right),$ (47)
which naturally produces a superconducting dome. Importantly, the suppression of $T_{c}$ away from $p_{\star}$ does not arise from the disappearance of the dynamical fixed point, but from the progressive truncation of its infrared extent.
In this interpretation, optimal doping does not correspond to a conventional symmetry–breaking quantum critical point. Rather, it marks the regime in which the self–organized infrared spectral phase is most fully exposed.
V.2 B. Phase stiffness and Uemura scaling from the slow–mode reservoir
While Sec. IV established that the pairing scale is algebraically enhanced by the infrared TDOS, the observable transition temperature is ultimately limited by global phase coherence. In systems with reduced superfluid density, the ordering temperature may be controlled by phase stiffness rather than by gap formation.
The long-wavelength phase dynamics of the Cooper field is described by the effective action
$S_{\theta}=\frac{\rho_{s}}{2}\int d\tau d^{d}r\left[(\partial_{\tau}\theta)^{2}+v_{s}^{2}(\nabla\theta)^{2}\right],$ (48)
where $\rho_{s}$ denotes the superfluid stiffness. The corresponding phase-ordering temperature scales as
$T_{\theta}\sim\rho_{s}.$ (49)
Within the slow–mode reservoir framework, the same infrared spectral weight that enhances pairing also governs coherent charge transport. The superfluid stiffness measures the energy cost of imposing a spatial phase twist, and is therefore proportional to the density of coherent paired carriers that remain dynamically correlated over long time scales.
In the memory–dominated regime, long–time coherence is controlled by the infrared TDOS weight $\rho_{0}$. Because the pairing amplitude itself scales as $\Delta\sim E_{\rm pair}\,\rho_{0}$ [Eq. (39)], the density of phase-coherent pairs is likewise proportional to the same infrared spectral weight. Integrating out amplitude fluctuations in the presence of the slow–mode reservoir therefore yields
$\rho_{s}\sim E_{\rm pair}\,\rho_{0},$ (50)
up to nonuniversal numerical factors.
Both pairing amplitude and phase rigidity are thus controlled by a single infrared parameter that quantifies the extent of the slow–mode reservoir.
Combining Eqs. (49) and (50) gives
$T_{c}\sim\rho_{s}\sim E_{\rm pair}\,\rho_{0},$ (51)
establishing a direct proportionality between the superconducting transition temperature and superfluid density.
This relation reproduces the empirical Uemura scaling observed across cuprates and other unconventional superconductors, where $T_{c}$ tracks the zero-temperature superfluid density. In contrast to weak-coupling BCS superconductors—where phase stiffness is large and essentially decoupled from gap formation—correlated systems operate in a regime in which infrared collective dynamics simultaneously controls pairing and coherence.
Within the present framework, this proportionality is not imposed phenomenologically. It follows directly from the TDOS-controlled infrared fixed point, where a single spectral parameter $\rho_{0}$ governs both the Cooper instability and phase stiffness. The observed scaling of $T_{c}$ with superfluid density thus emerges as a natural consequence of memory-dominated critical dynamics.
V.3 C. Dynamical signatures: long–time correlations and anomalous dissipation
A distinctive prediction of the present theory concerns dynamical response in the normal state. As shown in Sec. II, a flat TDOS produces a marginal nonanalytic retarded kernel with
$\mathrm{Im}\,\chi_{R}^{-1}(\omega)\propto\mathrm{sgn}(\omega),$ (52)
corresponding in the time domain to a universal long–memory kernel $K(t)\sim 1/t$. This implies scale–free temporal correlations extending over broad time windows.
Such long–time tails can produce 1/f1/f–type noise spectra and anomalously slow relaxation, phenomena widely reported in strange metals and cuprate superconductors. Unlike conventional overdamped criticality, where correlations decay exponentially and dynamics is effectively Markovian, the memory–dominated regime predicts non-Markovian response with persistent temporal structure. Because the same slow–mode reservoir controls both pairing enhancement and dynamical response, superconducting and normal–state anomalies become two manifestations of a single infrared spectral reorganization.
V.4 D. Unified infrared scaling picture
We emphasize that all experimental signatures discussed above—the dome, Uemura scaling, and long–time dynamical correlations—originate from the same infrared parameter $\rho_{0}$, the low–energy spectral weight of the slow–mode reservoir.
The TDOS controls: (i) the infrared extent of the slow–mode reservoir, (ii) the dynamical amplification of intrinsic pairing tendencies, (iii) the superfluid stiffness and phase coherence, and (iv) the emergence of non-Markovian long–time response. Quantum criticality thus acts not merely as a source of fluctuations but as a dynamical amplifier governed by infrared spectral reorganization that simultaneously governs thermodynamic and dynamical phenomena in strongly correlated superconductors.
This unified picture yields a concrete strategy for experimental tests: systems that exhibit enhanced low–frequency TDOS weight (or equivalently, strong long–time correlations) are expected to display increased pairing susceptibility, elevated $T_{c}$, and reduced phase stiffness consistent with Uemura scaling. The resulting real-frequency spectral function generated by the memory-dominated self-energy, including its universal scaling form and infrared structure, is derived in Appendix F.
VI VI. Discussion
The central result of this work is that infrared collective dynamics in strongly correlated quantum matter is organized by the structure of its relaxation-rate spectrum rather than solely by softening of a single order parameter. The time–scale density of states provides a quantitative measure of this organization. When the TDOS remains finite as $\lambda\to 0$, the system enters a memory–dominated regime characterized by long–time kernels $K(t)\sim 1/t$ and nonanalytic infrared response.
This viewpoint clarifies the limits of conventional coarse-grained field theories. Away from criticality, the relaxation spectrum is gapped in the infrared, leading to finite memory times and justifying reduction to local-in-time damping and a small number of collective fields. Near correlated critical regimes, however, relaxation rates condense into a continuous spectrum extending toward zero. In this regime, internal modes cannot be integrated out: the full decay spectrum remains dynamically active across time scales. The breakdown of Markovian coarse-graining is therefore not a model-dependent artifact but a universal consequence of dynamical criticality.
This observation highlights the structural difference from Hertz–Millis–type approaches. Conventional quantum critical theories assume overdamped local dynamics of a small set of bosonic modes, with dissipation controlled by Landau damping. In contrast, the present framework allows for spectral condensation of near-marginal relaxation channels. Criticality is reinterpreted not as single-mode softening but as reorganization of the entire decay spectrum. The TDOS thus replaces phenomenological damping coefficients with a microscopic spectral quantity governing infrared universality.
A direct physical consequence of this spectral reorganization is qualitative enhancement of superconducting instabilities. Superconductivity need not rely on finely tuned bosonic mediators. Instead, intrinsic short-range electronic pairing tendencies are dynamically amplified by the slow-mode reservoir. The exponential suppression characteristic of weak-coupling BCS theory is replaced by algebraic scaling controlled by infrared spectral weight. In this sense, superconducting order corresponds to condensation within an extended slow-mode manifold rather than binding via a sharp bosonic excitation.
A single infrared parameter, $\rho_{0}$, governs both thermodynamic and dynamical phenomena. The superconducting dome reflects truncation of long-time memory away from the optimal spectral reorganization point; Uemura scaling follows from the shared spectral origin of pairing and phase coherence; anomalous normal-state dynamics arises from the same non-Markovian kernel. Quantum criticality therefore acts as a dynamical amplifier linking transport, coherence, and pairing within a unified spectral framework.
The relaxation-rate organization emphasized here is not merely a phenomenological assumption. Both Liouvillian evolution in open quantum systems and generalized memory-kernel dynamics naturally decompose into continua of decay modes characterized by complex relaxation eigenvalues. Whenever long-time correlations are present, Laplace-domain representations generically produce continuous relaxation spectra. A dense TDOS therefore reflects a universal mathematical property of dynamical systems with extended temporal correlations, rather than a model-specific construction.
From this perspective, Hamiltonian energy eigenmodes describe conservative reversible dynamics, whereas relaxation spectra encode the irreversible organization governing collective behavior near criticality. Strongly correlated quantum matter near dynamical critical points may thus be more naturally classified by its relaxation-rate spectrum than by its excitation spectrum alone.
More broadly, the TDOS framework elevates dynamical spectral organization to a fundamental organizing principle of correlated quantum matter. Memory-dominated criticality represents a universality class distinct from traditional dissipative field theories, with potential implications for non-Fermi liquids, glassy systems, and self-organized critical dynamics.
Memory–dominated dynamics as a guiding principle for room-temperature superconductivity. Most contemporary efforts toward room-temperature superconductivity focus on strengthening microscopic pairing interactions through pressure, lattice engineering, or exotic chemical composition. Within the present framework, however, enhancing the bare pairing scale alone is insufficient. The decisive control parameter for high transition temperatures is the infrared dynamical organization of collective modes encoded in the relaxation-rate density of states.
A dense reservoir of near-marginal slow modes reorganizes the Cooper channel, converting moderate intrinsic electronic pairing into an algebraically amplified instability with $T_{c}\sim\rho_{0}E_{\rm pair}$. From this viewpoint, certain high-pressure hydrides or correlated materials may fail to reach extreme transition temperatures not because pairing is weak, but because their collective spectra do not self-organize into a flat TDOS regime.
The present theory therefore suggests a complementary design principle for superconductivity beyond material-specific glue engineering: stabilization of dynamical phases that support extensive slow-mode continua [27]. Control of electronic frustration, dimensionality, feedback mechanisms, or proximity to dynamical criticality may be as crucial as increasing microscopic attraction itself. Room-temperature superconductivity, in this perspective, corresponds to a dynamical phase characterized by maximal infrared memory rather than simply by extreme coupling strength.
VII VII. Conclusion
We have identified dynamical spectral organization as a fundamental principle governing infrared behavior in strongly correlated quantum matter. By formulating collective dynamics in terms of the time–scale density of states, we show that criticality is generically controlled not by local Markovian dissipation but by the accumulation of slow collective relaxation modes.
A finite infrared TDOS defines a memory–dominated universality class characterized by long–time kernels $K(t)\sim 1/t$ and nonanalytic dynamical response. Within this regime, intrinsic electronic pairing tendencies are dynamically amplified, converting marginal logarithmic instabilities into algebraic superconducting transitions governed by infrared spectral weight.
This mechanism provides a unified account of key experimental phenomena in unconventional superconductors, including superconducting domes, Uemura scaling of transition temperatures with superfluid stiffness, and anomalous long–time dynamical correlations in the normal state. These signatures arise naturally from infrared spectral reorganization, without requiring specific bosonic mediators or material-dependent fine tuning.
More broadly, the results recast quantum criticality as a dynamical process that enhances collective quantum phenomena through extended temporal correlations. Memory–dominated criticality therefore represents a universality class distinct from traditional dissipative field theories, with implications for non-Fermi liquids, glassy dynamics, and engineered quantum platforms.
In this perspective, high–temperature superconductivity appears as a generic consequence of infrared spectral organization in strongly correlated matter. Systematic exploration of TDOS-controlled dynamics may provide a pathway toward stabilizing and enhancing collective quantum states in future materials and devices.
Acknowledgements—This work was partially supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (IITP-RS-2025-02214780).
The author acknowledges the support of ChatGPT (GPT-5, OpenAI) for assistance in literature review and conceptual structuring during early development.
Supplementary Materials
Appendix A Appendix A: Relaxation-rate spectra as the fundamental dynamical basis of collective dynamics
In conventional quantum many-body theory, collective behavior is often organized in terms of energy eigenmodes of an underlying Hamiltonian,
$H|n\rangle=E_{n}|n\rangle,$ (S1)
with dynamical response expressed through oscillatory Fourier components at frequencies set by energy differences. This viewpoint is natural for closed systems whose dynamics is unitary and time-reversal invariant, and it emphasizes stationary states and conservative motion.
However, energy eigenmodes provide only a special organizing basis. In generic interacting settings—especially in open, driven, or internally coarse-grained systems—the physically relevant long-time organization is not set by Hamiltonian energies but by the stability spectrum of the effective dynamical generator. In this work, the central spectral object is therefore the distribution of relaxation rates governing irreversible decay and memory.
(I) Dynamical linearization and relaxation modes: For a general high-dimensional (possibly coarse-grained) dynamical system,
$\dot{x}(t)=F(x(t)),$ (S2)
linearization about a metastable configuration $x^{*}$ yields
$\delta\dot{x}(t)=-A\,\delta x(t),\qquad A=-\left.\frac{\partial F}{\partial x}\right|_{x^{*}}.$ (S3)
The eigenvalue problem
$Av_{i}=\lambda_{i}v_{i}$ (S4)
defines a spectrum of collective relaxation modes with temporal evolution
$\delta x_{i}(t)\sim e^{-\lambda_{i}t}.$ (S5)
The real parts of $\lambda_{i}$ set decay time scales, while imaginary parts encode oscillatory components. Purely oscillatory energy-mode dynamics corresponds to the singular limit $\mathrm{Re}\,\lambda_{i}=0$ for all modes, which is realized only in strictly Hamiltonian, non-dissipative evolution.
Thus the relaxation-rate spectrum $\{\lambda_{i}\}$ constitutes the natural dynamical basis for generic collective phenomena, including metastability, critical slowing down, and non-Markovian memory.
(II) Liouvillian formulation (closed, open, and non-Markovian dynamics): A conceptually complete formulation of nonequilibrium quantum dynamics is given in terms of the density matrix $\hat{\varrho}(t)$ and its dynamical generator. For a closed system, $\hat{\varrho}(t)$ evolves by the von Neumann equation,
$\frac{d\hat{\varrho}}{dt}=-i[H,\hat{\varrho}]\equiv\mathcal{L}_{H}\hat{\varrho}$, (S6)
where $\mathcal{L}_{H}$ is the Hamiltonian Liouvillian superoperator. If $H|n\rangle=E_{n}|n\rangle$, then the operator basis $|n\rangle\langle m|$ diagonalizes $\mathcal{L}_{H}$:
$\mathcal{L}_{H}\big(|n\rangle\langle m|\big)=-i(E_{n}-E_{m})\,|n\rangle\langle m|,$ (S7)
so the Liouvillian eigenvalues are
$\lambda_{nm}=-i(E_{n}-E_{m}),$ (S8)
lying purely on the imaginary axis and reproducing undamped oscillations.
For open quantum systems, the reduced density matrix typically obeys a master equation with a dissipative contribution. In the Markovian limit, the evolution takes Lindblad form,
[…]
where the jump operators $L_{\mu}$ encode decoherence and dissipation. The resulting Liouvillian $\mathcal{L}$ is generically non-Hermitian as a superoperator and therefore possesses complex eigenvalues
$\mathcal{L}R_{\alpha}=\lambda_{\alpha}R_{\alpha},\qquad\mathrm{Re}\,\lambda_{\alpha}\leq 0,$ (S10)
with decay rates set by $\mathrm{Re}\,\lambda_{\alpha}$. The density matrix admits the modal expansion
$\hat{\varrho}(t)=\hat{\varrho}_{\rm ss}+\sum_{\alpha\neq 0}c_{\alpha}\,e^{\lambda_{\alpha}t}\,R_{\alpha},$ (S11)
where $\hat{\varrho}_{\rm ss}$ is the stationary state ($\lambda_{0}=0$) and the long-time approach is controlled by the Liouvillian spectral gap and any near-marginal continuum.
More generally, when the environment or eliminated internal coordinates retain memory, the reduced dynamics becomes non-Markovian and takes a memory-kernel form (Nakajima–Zwanzig structure),
$\frac{d\hat{\varrho}}{dt}=\int_{0}^{t}ds\;\mathcal{K}(t-s)\,\hat{\varrho}(s),$ (S12)
where $\mathcal{K}(t)$ is a causal superoperator kernel. In Laplace space,
$\tilde{\hat{\varrho}}(s)=\big[s-\tilde{\mathcal{K}}(s)\big]^{-1}\hat{\varrho}(0),$ (S13)
so the poles and branch cuts of $s-\tilde{\mathcal{K}}(s)$ define the intrinsic relaxation spectrum. In this sense, even for quantum systems the relevant organizing structure for long-time dynamics is generically a relaxation-rate spectrum (possibly continuous), not an energy spectrum.
(III) Memory kernels and continuous relaxation spectra: The same spectral logic appears already at the level of linear causal dynamics for a collective coordinate $\phi(t)$:
$\dot{\phi}(t)+\int_{0}^{t}ds\,K(t-s)\phi(s)=\eta(t).$ (S14)
Taking the Laplace transform gives
$\tilde{\phi}(s)=\frac{\tilde{\eta}(s)}{s+\tilde{K}(s)}.$ (S15)
The analytic structure of $s+\tilde{K}(s)$ determines the intrinsic relaxation content: isolated poles correspond to discrete decay modes, while branch cuts encode continua of decay rates. Whenever long-time memory is present, the causal kernel admits a spectral representation of the form
$K(t)=\int_{0}^{\infty}d\lambda\,\rho(\lambda)e^{-\lambda t},$ (S16)
which defines the time-scale spectrum. A flat infrared TDOS produces the universal tail $K(t)\sim 1/t$.
Thus continuous relaxation spectra arise generically from causality and memory, independent of microscopic Hamiltonian structure.
(IV) Energy modes as a special limit: The familiar Fourier-mode picture of conservative dynamics corresponds to the singular limit in which the relevant spectrum collapses onto the imaginary axis. In the closed-system Liouvillian case this appears as $\lambda_{nm}=-i(E_{n}-E_{m})$, while in the memory-kernel representation it corresponds to eliminating all decay rates (no weight at $\mathrm{Re}\,\lambda>0$), recovering purely oscillatory behavior. Away from this limit, generic interacting and/or open systems naturally develop broad spectra of relaxation rates controlling infrared dynamics.
(V) Physical implication: The TDOS—the infrared distribution of collective relaxation rates— is therefore the fundamental organizing variable of collective dynamics in correlated systems. Energy eigenmodes describe an integrable, conservative limit, whereas relaxation-rate spectra capture the universal structure of irreversible, critical, and strongly interacting matter.
Memory-dominated criticality corresponds precisely to the formation of an extensive continuum of near-marginal relaxation modes. The resulting TDOS controls dynamical response, pairing enhancement, and infrared universality, providing a deeper and more generally applicable foundation than energy-based descriptions.
Appendix B Appendix B: Gaussian MSRJD integration of a relaxation-mode reservoir and emergence of the self-energy
In this appendix we derive how a continuum of relaxation (Ornstein–Uhlenbeck) modes generates a nonlocal-in-time memory kernel and an associated self-energy $\Sigma(\omega)$ upon Gaussian functional integration in the Martin–Siggia–Rose–Janssen–De Dominicis (MSRJD) formalism. The derivation is standard in structure (Gaussian path integration), but we present it in a form tailored to our “slow-mode reservoir” interpretation.
(I) Coupled stochastic dynamics: We consider a collective degree of freedom $\phi(t)$ coupled linearly to a continuum of reservoir modes $X_{\lambda}(t)$ labeled by their relaxation rates $\lambda>0$. The coupled stochastic dynamics is
$\dot{\phi}(t)$ $=-r\,\phi(t)+\int d\lambda\;g(\lambda)\,X_{\lambda}(t)+\eta(t),$ (S17)
$\dot{X}_{\lambda}(t)$ $=-\lambda\,X_{\lambda}(t)+g(\lambda)\,\phi(t)+\xi_{\lambda}(t).$ (S18)
The noises are assumed Gaussian, stationary, and white:
$\langle\eta(t)\eta(t^{\prime})\rangle$ $=2D_{\phi}\,\delta(t-t^{\prime}),$ (S19)
$\langle\xi_{\lambda}(t)\,\xi_{\lambda^{\prime}}(t^{\prime})\rangle$ $=2D_{\lambda}\,\delta(\lambda-\lambda^{\prime})\,\delta(t-t^{\prime}).$ (S20)
[…]
(II) MSRJD action for the coupled system: For a generic additive-noise Langevin equation $\dot{y}=f(y)+\zeta$ with $\langle\zeta(t)\zeta(t^{\prime})\rangle=2D\,\delta(t-t^{\prime})$, the MSRJD functional integral may be written (Ito convention) as
$Z$ $=\int\mathcal{D}y\,\mathcal{D}\tilde{y}\;$ (S21)
$\exp\!\left[-\int dt\;\tilde{y}(t)\big(\dot{y}(t)-f(y(t))\big)+\int dt\;D\,\tilde{y}(t)^{2}\right],$
up to an overall normalization independent of $y$. Applying (S21) to (S17)–(S18) gives
$Z=\int\mathcal{D}\phi\,\mathcal{D}\tilde{\phi}\;\prod_{\lambda}\mathcal{D}X_{\lambda}\,\mathcal{D}\tilde{X}_{\lambda}\;e^{-S[\phi,\tilde{\phi},\{X_{\lambda},\tilde{X}_{\lambda}\}]},$ (S22)
with total action
$S=S_{\phi}+\int d\lambda\;S_{\lambda},$ (S23)
where
$S_{\phi}$ $=\int dt\;\tilde{\phi}(t)\!\left[\dot{\phi}(t)+r\,\phi(t)-\int d\lambda\;g(\lambda)\,X_{\lambda}(t)\right]$
$\quad-\int dt\;D_{\phi}\,\tilde{\phi}(t)^{2},$ (S24)
$S_{\lambda}$ $=\int dt\;\tilde{X}_{\lambda}(t)\!\left[\dot{X}_{\lambda}(t)+\lambda\,X_{\lambda}(t)-g(\lambda)\,\phi(t)\right]$
$\quad-\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}.$ (S25)
(III) Isolating a single $\lambda$-mode contribution: Fix $\lambda$ and collect all terms in $S$ that contain $X_{\lambda}$ or $\tilde{X}_{\lambda}$. From (S24) we extract
$S_{\phi}\supset-\int dt\;\tilde{\phi}(t)\,g(\lambda)\,X_{\lambda}(t).$ (S26)
Combining (S26) with (S25) yields the $\lambda$-sector action
$S^{(\lambda)}[X_{\lambda},\tilde{X}_{\lambda};\phi,\tilde{\phi}]=\int dt\;\tilde{X}_{\lambda}(t)\big(\dot{X}_{\lambda}(t)+\lambda X_{\lambda}(t)\big)$ (S27)
$\quad-\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,\phi(t)-\int dt\;g(\lambda)\,\tilde{\phi}(t)\,X_{\lambda}(t)$
$\quad-\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}.$
In what follows, $\phi,\tilde{\phi}$ are treated as external sources.
(IV) Functional integration over $X_{\lambda}$: Define the linear operator
$\mathcal{L}_{\lambda}\equiv\partial_{t}+\lambda.$ (S28)
Then the first term in (S27) is
$\int dt\;\tilde{X}_{\lambda}(t)\,\mathcal{L}_{\lambda}X_{\lambda}(t)=\int dt\;\tilde{X}_{\lambda}(t)\big(\partial_{t}X_{\lambda}(t)+\lambda X_{\lambda}(t)\big).$ (S29)
We now rewrite this term so that $X_{\lambda}$ appears without derivatives. Using integration by parts,
$\int dt\;\tilde{X}_{\lambda}(t)\,\partial_{t}X_{\lambda}(t)$ (S30)
$\quad=\big[\tilde{X}_{\lambda}(t)\,X_{\lambda}(t)\big]_{t_{i}}^{t_{f}}-\int dt\;(\partial_{t}\tilde{X}_{\lambda}(t))\,X_{\lambda}(t).$
Assuming boundary terms vanish (e.g. $t_{i}\to$-$\infty$, $t_{f}\to+\infty$ with sufficient decay, or appropriate causal boundary conditions), we drop the bracketed term and obtain
$\int dt\;\tilde{X}_{\lambda}(t)\,\partial_{t}X_{\lambda}(t)=-\int dt\;(\partial_{t}\tilde{X}_{\lambda}(t))\,X_{\lambda}(t).$ (S31)
Therefore,
$\int dt\;\tilde{X}_{\lambda}(t)\,\mathcal{L}_{\lambda}X_{\lambda}(t)=-\int dt\;(\partial_{t}\tilde{X}_{\lambda}(t))\,X_{\lambda}(t)$ (S32)
$+\int dt\;\lambda\,\tilde{X}_{\lambda}(t)\,X_{\lambda}(t)=\int dt\;X_{\lambda}(t)\big(-\partial_{t}+\lambda\big)\tilde{X}_{\lambda}(t).$
Define the adjoint operator
$\mathcal{L}_{\lambda}^{\dagger}\equiv-\partial_{t}+\lambda.$ (S33)
Then (S27) becomes
$S^{(\lambda)}$ $=\int dt\;X_{\lambda}(t)\Big(\mathcal{L}_{\lambda}^{\dagger}\tilde{X}_{\lambda}(t)-g(\lambda)\tilde{\phi}(t)\Big)$ (S34)
$\quad-\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,\phi(t)-\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}.$
Now $X_{\lambda}$ enters only linearly. Hence its functional integral produces a functional delta constraint:
$\int\mathcal{D}X_{\lambda}\;\exp\!\left\{-\int dt\;X_{\lambda}(t)\,A(t)\right\}$ $\propto\delta[A(t)],$ (S35)
where here
$A(t)\equiv\mathcal{L}_{\lambda}^{\dagger}\tilde{X}_{\lambda}(t)-g(\lambda)\tilde{\phi}(t).$ (S36)
Therefore,
[…]
$\quad\exp\!\left[+\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,\phi(t)-\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}\right].$
The Jacobian factor associated with the functional constraint, $|\det\mathcal{L}_{\lambda}^{\dagger}|^{-1}$, is independent of $\phi,\tilde{\phi}$ and can therefore be absorbed into the overall normalization of $Z$.
(V) Solving the delta constraint for $\tilde{X}_{\lambda}$ via Green functions: The constraint
$\mathcal{L}_{\lambda}^{\dagger}\tilde{X}_{\lambda}(t)=g(\lambda)\tilde{\phi}(t),\qquad\mathcal{L}_{\lambda}^{\dagger}=-\partial_{t}+\lambda,$ (S38)
is a linear inhomogeneous ODE for $\tilde{X}_{\lambda}(t)$. Introduce the Green function $G_{\lambda}^{A}(t-t^{\prime})$ solving
$(-\partial_{t}+\lambda)\,G_{\lambda}^{A}(t-t^{\prime})=\delta(t-t^{\prime}).$ (S39)
Because the constraint is written with the adjoint operator $\mathcal{L}_{\lambda}^{\dagger}=-\partial_{t}+\lambda$, its Green function carries advanced support (equivalently, the retarded kernel will reappear after relabelling, ensuring causality of the $\phi$-sector response).
The solution with advanced support is
$G_{\lambda}^{A}(t)=\Theta(-t)\,e^{\lambda t}.$ (S40)
(Indeed, for $t\neq 0$, $(-\partial_{t}+\lambda)G_{\lambda}^{A}=0$ and the jump at $t=0$ produces the delta distribution.) Multiplying (S38) by $G_{\lambda}^{A}$ and integrating, we obtain
$\tilde{X}_{\lambda}(t)$ $=g(\lambda)\int dt^{\prime}\;G_{\lambda}^{A}(t-t^{\prime})\,\tilde{\phi}(t^{\prime}).$ (S41)
[…]
(VI) Effective action: memory kernel (self-energy) and induced colored noise: We now evaluate the two remaining terms in the exponent of (S37) using (S41).
1. The $\tilde{X}_{\lambda}\phi$ term (deterministic memory kernel).
Start from
$\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,\phi(t).$ (S42)
Insert (S41):
$\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,\phi(t)=\int dt\,dt^{\prime}\;g(\lambda)^{2}\,G_{\lambda}^{A}(t-t^{\prime})\,\tilde{\phi}(t^{\prime})\,\phi(t).$ (S43)
Relabel integration variables $(t,t^{\prime})\to(t^{\prime},t)$:
$\int dt\,dt^{\prime}\;g(\lambda)^{2}\,G_{\lambda}^{A}(t^{\prime}-t)\,\tilde{\phi}(t)\,\phi(t^{\prime}).$ (S44)
Using the identity $G_{\lambda}^{A}(t^{\prime}-t)=G_{\lambda}^{R}(t-t^{\prime})$ with the retarded Green function
$G_{\lambda}^{R}(t)\equiv\Theta(t)\,e^{-\lambda t},\qquad(\partial_{t}+\lambda)G_{\lambda}^{R}(t)=\delta(t),$ (S45)
we obtain
$\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,\phi(t)$ $=\int dt\,dt^{\prime}\;\tilde{\phi}(t)\,\underbrace{g(\lambda)^{2}\,G_{\lambda}^{R}(t-t^{\prime})}_{\equiv K_{\lambda}(t-t^{\prime})}\,\phi(t^{\prime}).$ (S46)
Thus a single OU mode contributes the causal memory kernel
$K_{\lambda}(t-t^{\prime})=g(\lambda)^{2}\,\Theta(t-t^{\prime})\,e^{-\lambda(t-t^{\prime})}.$ (S47)
2. The $\tilde{X}_{\lambda}^{2}$ term (induced colored noise kernel).
Next consider
$\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}.$ (S48)
Insert (S41):
$\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}$
$=\int dt\,dt^{\prime}\,dt^{\prime\prime}\;D_{\lambda}\,g(\lambda)^{2}\,G_{\lambda}^{A}(t-t^{\prime})\,G_{\lambda}^{A}(t-t^{\prime\prime})\,\tilde{\phi}(t^{\prime})\,\tilde{\phi}(t^{\prime\prime}).$ (S49)
Define the induced (generally nonlocal) noise kernel
$\mathcal{N}_{\lambda}(t^{\prime},t^{\prime\prime})\equiv D_{\lambda}\,g(\lambda)^{2}\int dt\;G_{\lambda}^{A}(t-t^{\prime})\,G_{\lambda}^{A}(t-t^{\prime\prime}),$ (S50)
so that
$\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}=\int dt^{\prime}\,dt^{\prime\prime}\;\tilde{\phi}(t^{\prime})\,\mathcal{N}_{\lambda}(t^{\prime},t^{\prime\prime})\,\tilde{\phi}(t^{\prime\prime}).$ (S51)
In equilibrium one may impose a fluctuation–dissipation relation linking $\mathcal{N}_{\lambda}$ and $K_{\lambda}$, but this is not required for the deterministic self-energy structure derived below.
3. Summing the reservoir and assembling $S_{\mathrm{eff}}$.
Combining (S46)–(S51) for all $\lambda$ yields the exact effective action
$S_{\mathrm{eff}}[\phi,\tilde{\phi}]=\int dt\;\tilde{\phi}(t)\big(\dot{\phi}(t)+r\phi(t)\big)-\int dt\;D_{\phi}\,\tilde{\phi}(t)^{2}$
$-\int dt\,dt^{\prime}\;\tilde{\phi}(t)\,K(t-t^{\prime})\,\phi(t^{\prime})-\int dt\,dt^{\prime}\;\tilde{\phi}(t)\,\mathcal{N}(t,t^{\prime})\,\tilde{\phi}(t^{\prime}),$ (S52)
where the total memory kernel is the reservoir sum
$K(t-t^{\prime})=\int d\lambda\;K_{\lambda}(t-t^{\prime})=\int d\lambda\;g(\lambda)^{2}\,\Theta(t-t^{\prime})\,e^{-\lambda(t-t^{\prime})},$ (S53)
and likewise $\mathcal{N}=\int d\lambda\,\mathcal{N}_{\lambda}$.
(VI) Frequency-domain self-energy $\Sigma(\omega)$ from the memory kernel: To connect (S52) to the standard response-function form, we Fourier transform using
$f(\omega)=\int_{-\infty}^{\infty}dt\;e^{i\omega t}\,f(t),\qquad f(t)=\int\frac{d\omega}{2\pi}\;e^{-i\omega t}\,f(\omega).$ (S54)
The bilinear term involving $K$ becomes
$\int dt\,dt^{\prime}\;\tilde{\phi}(t)\,K(t-t^{\prime})\,\phi(t^{\prime})$ $=\int\frac{d\omega}{2\pi}\;\tilde{\phi}(-\omega)\,K(\omega)\,\phi(\omega),$ (S55)
where $K(\omega)$ is the Fourier transform of the causal kernel (S53):
$K(\omega)$ $=\int_{-\infty}^{\infty}dt\;e^{i\omega t}\,K(t)=\int d\lambda\;g(\lambda)^{2}\int_{0}^{\infty}dt\;e^{i\omega t}e^{-\lambda t}.$ (S56)
The time integral is elementary:
$\int_{0}^{\infty}dt\;e^{-(\lambda-i\omega)t}$ $=\left[\frac{-1}{\lambda-i\omega}e^{-(\lambda-i\omega)t}\right]_{0}^{\infty}=\frac{1}{\lambda-i\omega},$
$\qquad(\mathrm{Re}\,\lambda>0).$ (S57)
Substituting (S57) into (S56) gives the exact spectral representation
$K(\omega)=\int d\lambda\;\frac{g(\lambda)^{2}}{\lambda-i\omega}.$ (S58)
In the MSRJD response sector, the retarded inverse propagator for $\phi$ is read off from the $\tilde{\phi}(-\omega)\phi(\omega)$ coefficient in $S_{\mathrm{eff}}$:
$G_{R}^{-1}(\omega)=-i\omega+r+K(\omega).$ (S59)
It is therefore natural to define the (retarded) self-energy as
$\Sigma_{R}(\omega)\equiv K(\omega)=\int d\lambda\;\frac{g(\lambda)^{2}}{\lambda-i\omega}.$ (S60)
(Our convention is $\Sigma_{R}(\omega)\equiv K(\omega)$ so that $G_{R}^{-1}(\omega)=-i\omega+r+\Sigma_{R}(\omega)$; other sign conventions for the self-energy appear in the literature.)
Equations (S58)–(S60) demonstrate that integrating out a continuum of OU relaxation modes produces a frequency- dependent self-energy whose analytic structure is determined by the relaxation-rate spectrum and couplings.
(VII) Connection to TDOS (relaxation-rate density of states): If the couplings are sufficiently structureless over the low-$\lambda$ sector, one may absorb $g(\lambda)^{2}$ into an effective spectral density. Defining
$\mathcal{J}(\lambda)\equiv g(\lambda)^{2},$ (S61)
we can write
$\Sigma_{R}(\omega)=\int d\lambda\;\frac{\mathcal{J}(\lambda)}{\lambda-i\omega}.$ (S62)
Equivalently, starting from a discrete set of modes labeled by $\alpha$,
$\Sigma_{R}(\omega)=\sum_{\alpha}\frac{g_{\alpha}^{2}}{\lambda_{\alpha}-i\omega},$ (S63)
the continuum limit gives
$\Sigma_{R}(\omega)$ $=\int d\lambda\;\rho(\lambda)\,\frac{\overline{g^{2}}(\lambda)}{\lambda-i\omega}\equiv\int d\lambda\;\frac{\mathcal{J}(\lambda)}{\lambda-i\omega},$
$\qquad\mathcal{J}(\lambda)\equiv\rho(\lambda)\,\overline{g^{2}}(\lambda),$ (S64)
where $\rho(\lambda)$ is the relaxation-rate density of states and $\overline{g^{2}}(\lambda)$ denotes the mean squared coupling of modes with rate $\lambda$.
In our framework the natural object is the time-scale density of states $\rho(\lambda)$ of relaxation rates. In the simplest case $\overline{g^{2}}(\lambda)$ varies weakly over the slow sector so that $\mathcal{J}(\lambda)\propto\rho(\lambda)$ in the infrared, and therefore
$\Sigma_{R}(\omega)\propto\int d\lambda\;\frac{\rho(\lambda)}{\lambda-i\omega}.$ (S65)
This makes explicit, at the level of a controlled Gaussian MSRJD integration, why the relaxation-rate spectrum provides the natural basis for the effective dynamics and response of the collective field.
(VIII) Coarse-grained field extension: We promote the collective coordinate to a coarse-grained field $\phi(\mathbf{r},t)$. At the Gaussian level the MSRJD action reads
$S_{0}=\int dt\,d^{d}r\;\tilde{\phi}\big(\partial_{t}+\Gamma(r-c\nabla^{2})\big)\phi-\int dt\,d^{d}r\,D\,\tilde{\phi}^{2}.$ (S66)
Coupling $\phi$ locally to an OU reservoir $X_{\lambda}(\mathbf{r},t)$ and integrating out $\{X_{\lambda},\tilde{X}_{\lambda}\}$ yields the nonlocal-in-time term
$S_{\rm eff}\supset-\int dt\,dt^{\prime}\,d^{d}r\;\tilde{\phi}(\mathbf{r},t)\,K(t-t^{\prime})\,\phi(\mathbf{r},t^{\prime}),$ (S67)
with $K(\omega)=\int d\lambda\,g(\lambda)^{2}/(\lambda-i\omega)\equiv\Sigma_{R}(\omega)$. Hence,
$\chi^{-1}(\omega,\mathbf{q})=r+cq^{2}+\Sigma(\omega)$ (S68)
(up to the conventional factor $\Gamma$ and the $i\omega$ term in the retarded response).
(IX) Summary of the exact Gaussian result: Starting from the coupled Langevin system (S17)–(S18), the MSRJD action (S24)–(S25) is Gaussian in the reservoir variables. Performing the functional integration over $\{X_{\lambda},\tilde{X}_{\lambda}\}$ yields the exact nonlocal effective action (S52) with a causal memory kernel (S53). In frequency space the kernel takes the spectral form (S58), which is naturally identified with the retarded self-energy (S60). This derivation is purely algebraic and involves no approximation beyond standard boundary-condition assumptions required for integration by parts.
(X) Relation to the Keldysh functional integral: We briefly clarify the formal relation between the MSRJD formulation employed in the present work and the real–time Keldysh functional integral for driven–dissipative quantum systems.
The MSRJD action for a stochastic dynamical variable $\phi(t)$ with memory kernel and Gaussian noise may be written as
$S_{\mathrm{MSRJD}}$ $=\int dt\,dt^{\prime}\,\tilde{\phi}(t)\left[\delta(t-t^{\prime})(\partial_{t}+r)-K_{R}(t-t^{\prime})\right]$ (S69)
$\quad\phi(t^{\prime})-\int dt\,dt^{\prime}\,\tilde{\phi}(t)\,\mathcal{N}(t,t^{\prime})\,\tilde{\phi}(t^{\prime}),$
where $\tilde{\phi}$ denotes the response field, $K_{R}$ is the retarded dynamical kernel (memory kernel), and $\mathcal{N}$ is the noise correlator introduced above.
The real–time Keldysh functional integral for an open quantum system is naturally formulated in terms of classical and quantum fields, $\phi_{\mathrm{cl}}$ and $\phi_{q}$, defined by the Keldysh rotation
$\phi_{\mathrm{cl}}=\frac{\phi_{+}+\phi_{-}}{2},\qquad\phi_{q}=\phi_{+}-\phi_{-},$ (S70)
where $\phi_{\pm}$ denote fields on the forward and backward branches of the time contour.
To quadratic order in the quantum field, the Keldysh action assumes the general structure
$S_{\mathrm{K}}=\int_{\omega}\begin{pmatrix}\phi_{\mathrm{cl}}&\phi_{q}\end{pmatrix}_{-\omega}\begin{pmatrix}0&K_{A}(\omega)\\ K_{R}(\omega)&iK_{K}(\omega)\end{pmatrix}\begin{pmatrix}\phi_{\mathrm{cl}}\\ \phi_{q}\end{pmatrix}_{\omega},$ (S71)
where $K_{R,A}$ are retarded/advanced kernels and $K_{K}$ is the Keldysh (noise) component. The retarded kernel encodes dissipative response, while the Keldysh component determines fluctuation amplitudes.
In the regime of strong decoherence or large occupation, where quantum interference effects are suppressed, the quantum field $\phi_{q}$ acts as a response field. Identifying
$\phi_{q}\;\leftrightarrow\;i\tilde{\phi}$, (S72)
and integrating out $\phi_{q}$ via a Hubbard–Stratonovich transformation, one recovers a stochastic Langevin equation with noise correlator set by $K_{K}$. The corresponding MSRJD action then takes the form (S69), with the correspondence
$K_{R}(\omega)\;\leftrightarrow\;K_{R}(\omega),\qquad\mathcal{N}(\omega)\;\leftrightarrow\;\tfrac{1}{2}K_{K}(\omega).$ (S73)
Thus the MSRJD formulation used in the present work can be viewed as the classical (or strongly dissipative) limit of the Keldysh functional integral, in which the relevant dynamical information is encoded in the retarded kernel and its associated spectral properties.
Importantly, our theory does not rely on coherent quantum interference or operator-valued field structure. Instead, the central object is the spectrum of the dynamical generator (equivalently, the retarded kernel), whose low-frequency density of states controls the infrared behavior and pairing amplification mechanism discussed in the main text.
Appendix C Appendix C: Exact emergence of a memory kernel from integrating out an OU reservoir
In this Appendix we provide a self-contained derivation showing that integrating out a continuum of relaxational Ornstein–Uhlenbeck (OU) modes generates a nonlocal-in-time (memory) kernel for the remaining collective coordinate. The result makes explicit that long-memory response is the generic outcome of eliminating hidden internal degrees of freedom, and that a flat time-scale density of states yields the universal tail $K(t)\sim 1/t$. This time-domain elimination is equivalent in content to the Gaussian MSRJD integration, but is presented here directly at the level of the coupled equations of motion.
(I) Coupled Markovian dynamics (collective field plus relaxational reservoir): For clarity we first suppress spatial dependence and focus on temporal structure; the extension to $\mathbf{r}$ (including $cq^{2}$ terms) is immediate by Fourier transforming in space. We consider a slow collective variable $\phi(t)$ linearly coupled to a continuum of relaxational OU modes $X_{\lambda}(t)$ labeled by decay rate $\lambda\geq 0$:
$\dot{\phi}(t)$ $=-r\,\phi(t)+\int_{0}^{\Lambda}d\lambda\,g(\lambda)\,X_{\lambda}(t)+\eta(t),$ (S74)
$\dot{X}_{\lambda}(t)$ $=-\lambda\,X_{\lambda}(t)+g(\lambda)\,\phi(t)+\zeta_{\lambda}(t).$ (S75)
Here $r>0$ is the bare restoring (mass) parameter, $\Lambda$ is an ultraviolet cutoff on relaxation rates, and $\eta$ and $\zeta_{\lambda}$ are (possibly independent) noise sources. For $\lambda>0$ each reservoir coordinate is exponentially stable; the sector $\lambda\to 0^{+}$ should be understood as the infrared limit of a continuous spectrum of increasingly slow relaxational modes rather than an exactly conserved degree of freedom. Equations (S74)–(S75) define a Markovian (first-order) dynamical system in the enlarged state space $\{\phi,X_{\lambda}\}$.
(II) Exact solution for the OU reservoir modes: Equation (S75) is linear and can be solved exactly by the integrating-factor method. Multiply (S75) by $e^{\lambda t}$:
$e^{\lambda t}\dot{X}_{\lambda}(t)+\lambda e^{\lambda t}X_{\lambda}(t)=e^{\lambda t}\big[g(\lambda)\phi(t)+\zeta_{\lambda}(t)\big].$ (S76)
The left-hand side is the total derivative
$\frac{d}{dt}\Big(e^{\lambda t}X_{\lambda}(t)\Big)=e^{\lambda t}\big[g(\lambda)\phi(t)+\zeta_{\lambda}(t)\big].$ (S77)
Integrate from an initial time $t_{0}$ to $t$:
$e^{\lambda t}X_{\lambda}(t)-e^{\lambda t_{0}}X_{\lambda}(t_{0})=\int_{t_{0}}^{t}ds\,e^{\lambda s}\big[g(\lambda)\phi(s)+\zeta_{\lambda}(s)\big].$ (S78)
Solving for $X_{\lambda}(t)$ yields the exact causal representation
$X_{\lambda}(t)$ $=e^{-\lambda(t-t_{0})}X_{\lambda}(t_{0})$ (S79)
$\quad+\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}\Big[g(\lambda)\phi(s)+\zeta_{\lambda}(s)\Big].$
This expression already shows that the reservoir coordinate at time $t$ depends on the entire history $\{\phi(s)\}_{s<t}$.
(III) Substitution into the collective equation and emergence of a memory kernel: Substituting (S79) into the collective equation (S74) gives
$\dot{\phi}(t)$ $=-r\,\phi(t)+\int_{0}^{\Lambda}d\lambda\,g(\lambda)\,$
$\Bigg[e^{-\lambda(t-t_{0})}X_{\lambda}(t_{0})+\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}\Big(g(\lambda)\phi(s)+\zeta_{\lambda}(s)\Big)\Bigg]$
$+\eta(t)$
This motivates the definition of the causal memory kernel
$K(t-s)\equiv\Theta(t-s)\int_{0}^{\Lambda}d\lambda\,g(\lambda)^{2}\,e^{-\lambda(t-s)},$ (S81)
so that (LABEL:eq:kernel_identification_step_app) becomes
$\int_{0}^{\Lambda}d\lambda\,g(\lambda)\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}g(\lambda)\phi(s)=\int_{t_{0}}^{t}ds\,K(t-s)\,\phi(s).$ (S82)
With this definition, Eq. (LABEL:eq:phi_substituted_app) can be written as
$\dot{\phi}(t)=-r\,\phi(t)+\int_{t_{0}}^{t}ds\,K(t-s)\phi(s)+\eta(t)+\xi_{\rm eff}(t)+J_{\rm ic}(t),$ (S83)
where we have collected the remaining contributions into an effective noise term (generated by the reservoir noise) and an initial-condition transient:
$\xi_{\rm eff}(t)$ $\equiv\int_{0}^{\Lambda}d\lambda\,g(\lambda)\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}\zeta_{\lambda}(s),$ (S84)
$J_{\rm ic}(t)$ $\equiv\int_{0}^{\Lambda}d\lambda\,g(\lambda)\,e^{-\lambda(t-t_{0})}X_{\lambda}(t_{0}).$ (S85)
Equation (S83) is an exact generalized Langevin equation for $\phi(t)$. Although the enlarged system (S74)–(S75) is Markovian, eliminating the hidden reservoir coordinates produces a non-Markovian (memory) term and, generally, colored effective noise.
Long-time limit.
For $t-t_{0}\gg\Lambda^{-1}$, the transient $J_{\rm ic}(t)$ decays. More precisely, decay holds provided the effective spectral weight $g(\lambda)X_{\lambda}(t_{0})$ is integrable near $\lambda=0$ so that the $\lambda\to 0^{+}$ sector does not produce a non-decaying contribution.
(IV) TDOS form and Laplace-transform structure: The memory kernel identified above depends only on the combination $g(\lambda)^{2}$ that weights the contribution of each relaxational mode to the collective dynamics. It is therefore natural to introduce the TDOS as the effective coupling-weighted spectral density of relaxation rates:
$\rho(\lambda)\equiv g(\lambda)^{2}.$ (S86)
More generally, in a microscopic description with discrete relaxational modes of density $\nu(\lambda)$ and coupling amplitudes $g_{\lambda}$, the kernel involves the combination $\nu(\lambda)|g_{\lambda}|^{2}$. For notational simplicity, we absorb any slowly varying factors into the definition of $\rho(\lambda)$. Thus $\rho(\lambda)$ should be understood as the effective spectral weight governing how strongly relaxation modes at scale $\lambda$ influence the collective coordinate $\phi$.
With this definition, the memory kernel becomes
$K(t)=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\,e^{-\lambda t},\qquad t\geq 0.$ (S87)
Equation (S87) shows that the memory kernel is precisely the Laplace transform of the TDOS (with ultraviolet cutoff $\Lambda$). In this sense the TDOS is the fundamental spectral object of the theory, while the time-domain memory kernel is its direct transform. The infrared dynamical behavior of the collective coordinate is therefore completely determined by the low-$\lambda$ structure of $\rho(\lambda)$.
(V) Flat TDOS implies universal long-memory tail $K(t)\sim 1/t$: If the TDOS is finite at the origin (“flat” TDOS),
$\rho(\lambda)\xrightarrow{\lambda\to 0}\rho_{0},$ (S88)
then for times $t\gg\Lambda^{-1}$ the infrared part of (S87) dominates and we may approximate $\rho(\lambda)\simeq\rho_{0}$ over the relevant range. In that case the kernel can be evaluated exactly:
$K(t)=\int_{0}^{\Lambda}d\lambda\,\rho_{0}\,e^{-\lambda t}.$ (S89)
Perform the change of variables $u=\lambda t$ (so $d\lambda=du/t$):
$K(t)$ $=\rho_{0}\int_{0}^{\Lambda t}\frac{du}{t}\,e^{-u}=\frac{\rho_{0}}{t}\Big(1-e^{-\Lambda t}\Big).$ (S90)
Therefore, at long times $t\gg\Lambda^{-1}$,
$K(t)\simeq\frac{\rho_{0}}{t},\qquad(t\gg\Lambda^{-1}),$ (S91)
which is the universal long-memory tail discussed in the main text.
(VI) Frequency-space form and connection to the retarded self-energy: For completeness we relate the time-domain memory kernel to the standard frequency-space self-energy. Taking the Fourier transform of the reservoir solution (S79) in the stationary (long-time) limit yields
$X_{\lambda}(\omega)=\frac{g(\lambda)}{\lambda-i\omega}\,\phi(\omega)+\frac{1}{\lambda-i\omega}\,\zeta_{\lambda}(\omega),$ (S92)
and substituting into (S74) gives
$\Big[-i\omega+r+\Sigma^{R}(\omega)\Big]\phi(\omega)=\eta(\omega)+\int_{0}^{\Lambda}d\lambda\,\frac{g(\lambda)}{\lambda-i\omega}\zeta_{\lambda}(\omega),$ (S93)
with the retarded self-energy
$\Sigma^{R}(\omega)=\int_{0}^{\Lambda}d\lambda\,\frac{g(\lambda)^{2}}{\lambda-i\omega}=\int_{0}^{\Lambda}d\lambda\,\frac{\rho(\lambda)}{\lambda-i\omega}.$ (S94)
We use the convention $G_{R}^{-1}(\omega)=-i\omega+r+\Sigma^{R}(\omega)$; other sign conventions for the self-energy appear in the literature.
Separating real and imaginary parts using
$\frac{1}{\lambda-i\omega}=\frac{\lambda}{\lambda^{2}+\omega^{2}}+i\,\frac{\omega}{\lambda^{2}+\omega^{2}},$ (S95)
one obtains
$\mathrm{Re}\,\Sigma^{R}(\omega)$ $=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{\lambda}{\lambda^{2}+\omega^{2}},$ (S96)
$\mathrm{Im}\,\Sigma^{R}(\omega)$ $=\omega\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{1}{\lambda^{2}+\omega^{2}}.$ (S97)
For a flat TDOS $\rho(\lambda)\simeq\rho_{0}$ and $|\omega|\ll\Lambda$,
$\mathrm{Im}\,\Sigma^{R}(\omega)$ $\simeq\omega\rho_{0}\int_{0}^{\Lambda}\frac{d\lambda}{\lambda^{2}+\omega^{2}}=\omega\rho_{0}\left[\frac{1}{|\omega|}\arctan\!\Big(\frac{\Lambda}{|\omega|}\Big)\right]$
$\xrightarrow{|\omega|\ll\Lambda}\frac{\pi}{2}\rho_{0}\,\mathrm{sgn}(\omega),$ (S98)
and
$\mathrm{Re}\,\Sigma^{R}(\omega)$ $\simeq\rho_{0}\int_{0}^{\Lambda}d\lambda\,\frac{\lambda}{\lambda^{2}+\omega^{2}}=\frac{\rho_{0}}{2}\int_{0}^{\Lambda}d\lambda\,\frac{2\lambda}{\lambda^{2}+\omega^{2}}$
$=\frac{\rho_{0}}{2}\left[\ln(\lambda^{2}+\omega^{2})\right]_{\lambda=0}^{\lambda=\Lambda}$
$=\frac{\rho_{0}}{2}\ln\!\left(\frac{\Lambda^{2}+\omega^{2}}{\omega^{2}}\right)\simeq\rho_{0}\ln\!\left(\frac{\Lambda}{|\omega|}\right),$
$\qquad(|\omega|\ll\Lambda),$ (S99)
which reproduces the logarithmic reactive singularity discussed in the main text.
Summary. Equations (S83)–(S87) show that integrating out a continuum of OU relaxation modes yields an exact generalized Langevin equation for the collective coordinate with a causal memory kernel $K(t)$. The kernel is the Laplace transform of the TDOS. A finite TDOS at the origin produces the universal long-memory tail $K(t)\sim 1/t$, and the corresponding retarded self-energy exhibits the marginal nonanalytic structure $\mathrm{Im}\,\Sigma^{R}(\omega)\propto\mathrm{sgn}(\omega)$ together with a logarithmic real part. This makes explicit that memory-dominated infrared dynamics arises from the self-organized accumulation of slow internal relaxation modes, rather than from coupling to a fast Markovian bath.
(VII) Conceptual structure (from hidden modes to infrared memory): The derivation above establishes a general structural result. A Markovian dynamical system defined in an extended state space $\{\phi,X_{\lambda}\}$ becomes non-Markovian when the internal degrees of freedom $X_{\lambda}$ are eliminated. The resulting equation for $\phi$ necessarily acquires a causal memory kernel given exactly by the Laplace transform of the time-scale density of states,
$K(t)=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)e^{-\lambda t}.$ (S100)
This relation is purely kinematic and does not rely on perturbation theory, weak coupling, or specific microscopic assumptions. The infrared dynamical structure of the collective field is therefore completely determined by the low-$\lambda$ scaling of $\rho(\lambda)$.
If the TDOS vanishes at the origin, $\rho(\lambda)\to 0$ as $\lambda\to 0$, the kernel decays exponentially and the dynamics reduces to effectively Markovian overdamping. By contrast, when the TDOS remains finite at zero relaxation rate,
$\rho(\lambda\to 0)=\rho_{0}\neq 0,$ (S101)
the Laplace transform produces the universal long-time tail
$K(t)\simeq\frac{\rho_{0}}{t}.$ (S102)
Thus long-memory infrared dynamics is not an additional assumption but an unavoidable consequence of a flat slow-mode spectrum. The TDOS provides the fundamental organizing principle linking microscopic stability structure to macroscopic temporal response.
In this sense, the emergence of memory-dominated criticality reflects a dynamical phase characterized by an extensive accumulation of near-marginal relaxation modes rather than by coupling to a fast bath.
Appendix D Appendix D: Renormalization-group flow of the Cooper channel in conventional and memory-dominated regimes
In this Appendix we derive explicitly the infrared renormalization-group (RG) flow equations for the effective pairing interaction in the Cooper channel. We contrast the conventional Eliashberg regime with the memory-dominated regime generated by a flat relaxation-rate density of states. The derivation follows from shell integration of the linearized gap equation and clarifies the origin of marginal versus relevant superconducting instabilities.
Fermionic Matsubara frequencies are denoted by $\omega$ and $\omega^{\prime}$, and $\Omega\equiv\omega-\omega^{\prime}$ is the bosonic transfer frequency entering the pairing kernel $g(\Omega)$.
(I) Linearized gap equation and RG shell construction: At low energies the linearized gap equation takes the schematic form
$\Delta(\omega)=\int_{T}^{\Lambda}\frac{d\omega^{\prime}}{|\omega^{\prime}|}\,g(\omega-\omega^{\prime})\,\Delta(\omega^{\prime}),$ (S103)
where $\Lambda$ is the ultraviolet cutoff and $T$ provides the infrared cutoff. The factor $1/|\omega^{\prime}|$ originates from the fermionic Cooper bubble. The effective pairing kernel $g(\Omega)$ is generated by collective fluctuations.
The RG procedure consists of integrating out a thin frequency shell $\omega^{\prime}\in[\Lambda/b,\Lambda]$ with $b>1$ and absorbing its contribution into a renormalized coupling $g\rightarrow g+dg$. Defining the RG scale $l=\ln b$, the flow equation follows from the shell contribution.
(II) Conventional Eliashberg regime: Near metallic quantum criticality with Ohmic Landau damping, the bosonic propagator obeys
$D(\Omega)\sim\frac{1}{|\Omega|},$ (S104)
so that the pairing kernel scales as
$g(\Omega)\sim g\,D(\Omega)\sim\frac{g}{|\Omega|},$ (S105)
where $g$ denotes the dimensionless pairing interaction.
The leading shell correction arises from the ladder diagram and is proportional to
$dg\propto g^{2}\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{|\omega^{\prime}|}.$ (S106)
The shell integral is elementary:
$\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{\omega^{\prime}}=\ln\Lambda-\ln(\Lambda/b)=\ln b\equiv dl.$ (S107)
Hence the RG equation takes the marginal form
$\frac{dg}{dl}=C\,g^{2},$ (S108)
with $C$ a positive numerical constant.
The solution is
$g(l)=\frac{g_{0}}{1-Cg_{0}l}\simeq g_{0}+Cg_{0}^{2}l,$ (S109)
demonstrating logarithmic enhancement characteristic of a marginal superconducting instability.
(III) Memory-dominated regime: In the memory-dominated regime derived in Sec. II, the collective susceptibility behaves as
$\chi(\Omega)\sim\frac{1}{\rho_{0}\ln(\Lambda/|\Omega|)},$ (S110)
reflecting the flat TDOS of slow relaxation modes.
The pairing kernel therefore scales as
$g(\Omega)\sim\frac{g}{\ln(\Lambda/|\Omega|)}.$ (S111)
The shell correction becomes
$dg\propto g^{2}\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{|\omega^{\prime}|}\frac{1}{\ln(\Lambda/\omega^{\prime})}.$ (S112)
Introducing the variable
$u=\ln\frac{\Lambda}{\omega^{\prime}},\qquad du=-\frac{d\omega^{\prime}}{\omega^{\prime}},$ (S113)
the integration limits transform as
$\omega^{\prime}=\Lambda\Rightarrow u=0,\qquad\omega^{\prime}=\Lambda/b\Rightarrow u=\ln b\equiv dl.$ (S114)
Equation (S112) becomes
$dg\propto g^{2}\int_{0}^{dl}\frac{du}{u}.$ (S115)
The apparent divergence at $u\to 0$ is cut off at a microscopic scale $u_{\min}=O(1)$ marking the onset of the logarithmic TDOS regime. The shell contribution reduces to
$dg=\alpha\,g\,dl,$ (S116)
where $\alpha>0$ depends on the spectral weight $\rho_{0}$ and the onset scale of the memory-dominated regime.
The RG equation is therefore
$\frac{dg}{dl}=\alpha\,g,$ (S117)
with solution
$g(l)=g_{0}e^{\alpha l}=g_{0}b^{\alpha}.$ (S118)
The pairing interaction thus grows as a power law under infrared coarse graining, signaling a relevant superconducting instability.
(IV) Physical origin of marginal versus relevant behavior: The qualitative difference between Eqs. (S108) and (S117) originates from the infrared structure of the pairing kernel.
In the Eliashberg regime, each logarithmic shell contributes independently, leading to cumulative logarithmic growth. In contrast, the memory-dominated kernel generated by a flat TDOS correlates fluctuations across all time scales, producing linear amplification of the effective interaction under RG.
Consequently, even moderate intrinsic electronic pairing scales are dynamically amplified into a runaway superconducting instability once a slow-mode reservoir is formed.
Summary. Ohmic Landau damping yields a marginal Cooper-channel flow $dg/dl\propto g^{2}$, producing only logarithmic enhancement. A flat relaxation-rate TDOS instead generates a memory-dominated kernel that renders the pairing interaction relevant, with $dg/dl\propto g$ and power-law growth. This dynamical reorganization of infrared spectral weight underlies the algebraic superconducting instability central to the present theory.
Appendix E Appendix E: Finite-temperature regulation of the infrared divergence
In this Appendix we explicitly demonstrate how the zero–temperature infrared divergence of the polarization kernel, $\Pi(\omega)\sim 1/|\omega|$, is regulated at finite temperature, yielding the universal scaling $\Pi(T)\sim 1/T$. This establishes rigorously that the thermal scale acts as the infrared cutoff in the memory–dominated critical regime.
(I) Spectral integral representation: In the slow–mode reservoir regime the Cooper-channel polarization admits the spectral form
$\Pi(\omega)=\int_{0}^{\Lambda}d\lambda\;\frac{\rho_{\rm eff}(\lambda)}{\lambda^{2}+\omega^{2}}.$ (S119)
[…]
$\Pi(\omega)\simeq\rho_{0}\int_{0}^{\Lambda}d\lambda\;\frac{1}{\lambda^{2}+\omega^{2}}.$ (S120)
The elementary integral yields
$\int_{0}^{\Lambda}\frac{d\lambda}{\lambda^{2}+\omega^{2}}=\frac{1}{|\omega|}\tan^{-1}\!\left(\frac{\Lambda}{|\omega|}\right).$ (S121)
In the low–frequency limit $|\omega|\ll\Lambda$, one obtains
$\Pi(\omega)\simeq\frac{\pi\rho_{0}}{2}\,\frac{1}{|\omega|},$ (S122)
which expresses the nonanalytic infrared enhancement at zero temperature.
(II) Thermal infrared cutoff: At finite temperature the imaginary–time direction is compactified to the interval $\tau\in[0,\beta]$ with $\beta=1/T$, and dynamical correlations are expressed in terms of discrete Matsubara frequencies. As a consequence, continuous access to arbitrarily small frequencies is lost, and the thermal scale provides an effective infrared regulator.
Operationally, the zero–temperature denominator $\omega^{2}$ in Eq. (S120) is replaced by the minimal dynamical scale of order $T^{2}$, yielding
$\Pi(T)\simeq\rho_{0}\int_{0}^{\Lambda}d\lambda\;\frac{1}{\lambda^{2}+(cT)^{2}},$ (S123)
where $c$ is a numerical constant of order unity determined by the Matsubara spectrum.
Evaluating the integral gives
$\Pi(T)=\rho_{0}\,\frac{1}{cT}\tan^{-1}\!\left(\frac{\Lambda}{cT}\right)\simeq\frac{\pi\rho_{0}}{2c}\,\frac{1}{T},\qquad T\ll\Lambda.$ (S124)
Thus the infrared divergence is universally cut off by the thermal scale,
$\Pi(T)\propto\frac{1}{T}.$ (S125)
(III) Equivalent Matsubara summation: The same result follows directly from the finite–temperature Matsubara representation of the polarization bubble,
$\Pi(0,T)\sim T\sum_{\omega_{m}}\int_{0}^{\Lambda}d\lambda\;\frac{\rho_{0}}{\lambda^{2}+\omega_{m}^{2}},\qquad\omega_{m}=(2m+1)\pi T.$ (S126)
Using the standard identity
$T\sum_{\omega_{m}}\frac{1}{\lambda^{2}+\omega_{m}^{2}}=\frac{1}{2\lambda}\tanh\!\left(\frac{\lambda}{2T}\right),$ (S127)
and expanding for $\lambda\ll T$,
$\tanh\!\left(\frac{\lambda}{2T}\right)\simeq\frac{\lambda}{2T},$ (S128)
one obtains
$T\sum_{\omega_{m}}\frac{1}{\lambda^{2}+\omega_{m}^{2}}\simeq\frac{1}{2T}.$ (S129)
Substitution into Eq. (S126) yields
$\Pi(0,T)\sim\int_{0}^{\Lambda}d\lambda\;\rho_{0}\frac{1}{2T}\propto\frac{\rho_{0}}{T},$ (S130)
confirming that finite temperature universally regulates the infrared divergence.
(IV) Implication for the superconducting transition: The Thouless criterion,
$\chi_{\Delta}^{-1}(0,T_{c})=0\qquad\Longleftrightarrow\qquad\frac{1}{g_{\rm pair}}=\Pi(T_{c}),$ (S131)
therefore gives
$T_{c}\sim\rho_{0}\,g_{\rm pair},$ (S132)
or equivalently $T_{c}\sim\rho_{0}E_{\rm pair}$ in terms of the bare electronic pairing scale. The exponential suppression characteristic of BCS theory is thus replaced by algebraic scaling controlled by the infrared spectral weight of the slow–mode reservoir.
Appendix F Appendix F: Spectral function generated by memory–dominated self–energy
In this Appendix we derive the real–frequency spectral function associated with the collective retarded susceptibility in the memory–dominated critical regime. We show that a flat slow–mode TDOS produces a broad non–Lorentzian continuum with logarithmic infrared scaling and a universal collapse form.
(I) Retarded susceptibility in the memory–dominated regime: In the infrared regime controlled by a flat TDOS, the retarded inverse susceptibility takes the universal form
$\chi_{R}^{-1}(\omega,\mathbf{q})=\Delta_{\mathbf{q}}+\rho_{0}\ln\!\left(\frac{\Lambda}{|\omega|}\right)+i\Gamma\,\mathrm{sgn}(\omega),$ (S133)
where
$\Delta_{\mathbf{q}}\equiv r+cq^{2},\qquad\Gamma=\frac{\pi}{2}\rho_{0},$ (S134)
and $\Lambda$ is the ultraviolet cutoff of the slow–mode reservoir. Accordingly,
$\chi_{R}(\omega,\mathbf{q})=\frac{1}{\Delta_{\mathbf{q}}+\rho_{0}\ln(\Lambda/|\omega|)+i\Gamma\,\mathrm{sgn}(\omega)}.$ (S135)
(Throughout we use $\omega$ for real frequency; Matsubara frequencies are denoted separately in Appendix D.)
(II) Definition of the spectral function: The spectral function of collective fluctuations is defined by
$A(\omega,\mathbf{q})\equiv-2\,\mathrm{Im}\,\chi_{R}(\omega,\mathbf{q}).$ (S136)
Writing $\chi_{R}^{-1}=X(\omega,\mathbf{q})+iY(\omega)$ with
$X(\omega,\mathbf{q})$ $=\Delta_{\mathbf{q}}+\rho_{0}\ln\!\left(\frac{\Lambda}{|\omega|}\right),$ (S137)
$Y(\omega)$ $=\Gamma\,\mathrm{sgn}(\omega),$ (S138)
one has $\chi_{R}=(X-iY)/(X^{2}+Y^{2})$, hence
$A(\omega,\mathbf{q})=\frac{2\Gamma}{\left[\Delta_{\mathbf{q}}+\rho_{0}\ln(\Lambda/|\omega|)\right]^{2}+\Gamma^{2}}.$ (S139)
(III) Infrared scaling at criticality: Along the critical manifold $\Delta_{\mathbf{q}}=0$,
$A(\omega,\mathbf{q}=0)=\frac{2\Gamma}{\rho_{0}^{2}\ln^{2}(\Lambda/|\omega|)+\Gamma^{2}}.$ (S140)
For $|\omega|\ll\Lambda$ the logarithm dominates and the universal infrared tail is
$A(\omega)\xrightarrow{|\omega|\to 0}\frac{2\Gamma}{\rho_{0}^{2}}\,\frac{1}{\ln^{2}(\Lambda/|\omega|)}.$ (S141)
Thus the spectral weight is suppressed only logarithmically toward zero frequency, reflecting the accumulation of memory across many time scales.
(IV) Peak / hump scale away from criticality: For finite $\Delta_{\mathbf{q}}$, Eq. (S139) is maximized when
$\Delta_{\mathbf{q}}+\rho_{0}\ln\!\left(\frac{\Lambda}{|\omega_{*}|}\right)=0,$ (S142)
which yields
$|\omega_{*}|=\Lambda\exp\!\left(\frac{\Delta_{\mathbf{q}}}{\rho_{0}}\right).$ (S143)
For $\Delta_{\mathbf{q}}>0$ this scale lies above the infrared window ($|\omega_{*}|\gtrsim\Lambda$) and no low–energy peak develops, whereas for $\Delta_{\mathbf{q}}<0$ a pronounced low–frequency hump appears at $|\omega_{*}|\ll\Lambda$.
(V) Universal scaling collapse: Define the scaling variable
$u(\omega,\mathbf{q})\equiv\frac{\Delta_{\mathbf{q}}}{\rho_{0}}+\ln\!\left(\frac{\Lambda}{|\omega|}\right).$ (S144)
Using $\Gamma=(\pi/2)\rho_{0}$, Eq. (S139) reduces to the universal form
$A(\omega,\mathbf{q})=\frac{\pi}{\rho_{0}}\,\frac{1}{u(\omega,\mathbf{q})^{2}+(\pi/2)^{2}},$ (S145)
so that all momentum and tuning dependence enters only through $u$, implying a collapse of line shapes onto a single curve.
(VI) Finite–temperature infrared cutoff: At finite temperature the logarithmic growth is regulated by the thermal scale. A minimal phenomenological implementation consistent with Appendix D is
$|\omega|\;\rightarrow\;\max\{|\omega|,cT\},\qquad c=O(1),$ (S146)
yielding
$A_{T}(\omega,\mathbf{q})=\frac{2\Gamma}{\left[\Delta_{\mathbf{q}}+\rho_{0}\ln(\Lambda/\max\{|\omega|,cT\})\right]^{2}+\Gamma^{2}}.$ (S147)
For $|\omega|\ll T$ the logarithm saturates, producing a finite infrared spectral weight rather than the zero–temperature logarithmic suppression.
(VII) Physical interpretation: Unlike conventional overdamped criticality, where $\mathrm{Im}\,\Sigma^{R}(\omega)$ vanishes linearly with $|\omega|$, the memory–dominated regime exhibits frequency–independent damping ($\propto\mathrm{sgn}(\omega)$) accompanied by a logarithmically singular reactive part. The resulting spectral function is therefore a broad continuum with slowly varying infrared tails rather than a sharp Lorentzian quasiparticle peak.
Equation (S139) (or equivalently the collapse form (S145)) provides a directly testable dynamical signature of the slow–mode reservoir and the associated TDOS universality class.