### v3

# Memory-Dominated Quantum Criticality as a Universal Route to High-Temperature Superconductivity

Abstract
[…]
I I. Introduction
High–temperature superconductivity remains one of the central unsolved problems of strongly correlated quantum matter [1, 2]. Despite decades of intensive theoretical and experimental effort, no consensus has emerged on the microscopic mechanism responsible for robust pairing far above conventional phonon scales. A central difficulty lies in the intimate proximity of superconductivity to quantum critical regimes, where strong collective fluctuations dominate low–energy dynamics.
Within the prevailing theoretical paradigm, quantum critical fluctuations are assumed to be overdamped by fast electronic continua [3, 4]. This framework, formalized in Hertz–Millis theory, predicts a universal Ohmic Landau damping kernel of the form $|\omega|$ governing collective dynamics. While remarkably successful in weakly correlated metals, this assumption has increasingly come into tension with experiments in strongly correlated systems, including cuprate superconductors and strange metals, where anomalous temporal correlations, non-Markovian response, and $1/f$-like noise are ubiquitously observed [5, 6, 7, 8, 9].
These observations raise a fundamental question: Is overdamped Markovian critical dynamics truly universal, or does it rely on implicit assumptions about the structure of dynamical spectra that break down in strongly correlated regimes?
A key limitation of many conventional approaches lies not in the use of collective fields themselves, but in the assumption that the infrared dynamics can be adequately described by a small number of soft bosonic modes with local-in-time damping [10, 11]. Such descriptions effectively close the theory within a finite set of order-parameter degrees of freedom. While this approximation is often adequate in weakly interacting systems, its validity becomes less clear in strongly correlated electronic matter, where experimentally relevant observables correspond to coarse-grained collective sectors that continuously exchange coherence and energy with a large number of internal degrees of freedom. In such situations the effective dynamics need not be governed by a finite set of collective modes, but may instead involve a broad spectrum of relaxation processes spanning many time scales.
Although the full many-body system evolves unitarily under a Hamiltonian, the effective dynamics of these collective sectors is more naturally described as that of an open subsystem. In this situation the appropriate dynamical generator is not simply the Hamiltonian itself but the effective Liouvillian governing relaxation and memory processes of the reduced density matrix [12, 13, 14]. The long-time behavior of the system is therefore controlled by the spectrum of relaxation eigenvalues of this Liouvillian operator. From this perspective, dynamical criticality need not correspond to the softening of a single excitation mode but may instead arise from a reorganization of the entire relaxation-rate spectrum.
Rather than reducing the dynamics to a few collective coordinates, one must therefore consider the possibility that the Liouvillian spectrum develops an extended continuum of near-marginal decay modes. Such spectral accumulation naturally generates a broad distribution of relaxation time scales and invalidates the assumption of Markovian coarse-graining near criticality. The relevant organizing structure of the infrared theory is then not a small set of overdamped bosons, but the time–scale density of states (TDOS) of collective relaxation modes.
In this work we show that infrared collective dynamics is generically controlled by the TDOS of correlated internal decay channels. By formulating a dynamical coarse–graining theory within the Martin–Siggia–Rose–Janssen–De Dominicis (MSRJD) framework [15, 16, 17], we derive an exact spectral representation of the collective susceptibility in terms of the relaxation–rate density, without assuming a priori a finite set of soft modes.
We demonstrate that the infrared scaling of the relaxation-rate density of states, $\rho(\lambda)\sim\lambda^{\alpha}$, naturally classifies universality classes of dynamical critical behavior. Within the relaxation-spectrum framework, this scaling directly determines the low-frequency structure of the retarded susceptibility. While the conventional Hertz–Millis kernel $\mathrm{Im}\,\chi^{-1}\propto|\omega|$ corresponds to the case $\alpha=1$, arising microscopically from particle–hole polarization of a Fermi liquid, a finite TDOS at the origin ($\alpha=0$) produces a marginal nonanalytic kernel proportional to $\mathrm{sgn}(\omega)$. In the time domain this corresponds to a universal long–memory response $K(t)\sim 1/t$. The latter regime reflects a fundamentally different physical mechanism: rather than overdamped order-parameter fluctuations generated by a fermionic bath, the infrared dynamics is governed by the collective organization of the relaxation spectrum itself. This memory-dominated regime therefore defines a distinct universality class controlled by the structure of the TDOS.
Crucially, we show that such flat relaxation-rate spectra arise naturally in strongly correlated collective systems, including dynamical manifolds explored in our previous work [18, 19, 20]. The resulting long-memory kernel dynamically amplifies intrinsic electronic pairing tendencies, providing a route to strong superconductivity without invoking finely tuned bosonic mediators.
Figure 1 summarizes the resulting dynamical universality map of correlated quantum matter. The infrared structure of the relaxation spectrum determines collective kernels in the time domain and controls physical consequences including pairing enhancement, superconducting domes, Uemura scaling, and anomalous dynamical response [1, 21, 22]. Charge, spin, and pairing channels reorganize within a common slow-mode reservoir, with superconductivity corresponding to condensation in the particle–particle sector of a shared infrared spectral continuum.
Our results establish the relaxation-rate spectrum as a fundamental organizing principle of quantum critical dynamics and reveal memory–dominated criticality as a generic route toward high–temperature superconductivity. This framework reconciles anomalous dynamical phenomena observed in strange metals with robust pairing and provides a unified spectral perspective on collective behavior across strongly correlated systems [5, 7, 10].
The remainder of this paper is organized as follows. In Section II we derive the exact TDOS representation of the collective susceptibility and classify infrared universality classes. In Sections III and IV we analyze the resulting memory–dominated pairing kernel and its implications for superconducting instabilities. In Section V we discuss experimental consequences including superconducting domes, Uemura scaling, and dynamical anomalies. We conclude in Section VI with a summary and outlook.
Figure 1: Dynamical universality map of correlated quantum matter. (a) Infrared time–scale density of states $\rho(\lambda)$ defining distinct dynamical universality classes. Vanishing TDOS yields conventional Hertz–Millis and phonon dynamics, marginal spectra produce Ohmic damping, while a flat TDOS defines memory–dominated criticality. (b) Corresponding long–time kernels in the time domain, ranging from Markovian dissipation to non-Markovian power–law memory. (c) Physical consequences of memory–dominated criticality, including algebraic pairing enhancement, superconducting domes, Uemura scaling, strange metal transport, and long–time noise correlations.
II II. Emergence of a Slow-Mode Reservoir from Correlated Dynamical Spectra
Strongly correlated electronic matter is intrinsically a quantum many-body system. At long times and coarse-grained scales, its dynamics is governed not by isolated quasiparticles, but by collective relaxation channels generated by the full quantum dynamical evolution of the density matrix [12, 13, 14],
$\frac{d\hat{\varrho}}{dt}=\mathcal{L}\hat{\varrho}$, (1)
where $\mathcal{L}$ denotes the Liouvillian generator of the effective (open or coarse-grained) dynamics.
The eigenmodes of $\mathcal{L}$,
$\mathcal{L}R_{n}=\lambda_{n}R_{n},$ (2)
define the fundamental decay channels of the interacting system. For stable dynamics, $\mathrm{Re}\,\lambda_{n}\leq 0$, and we define the collective relaxation rates as
$\lambda\equiv-\mathrm{Re}\,\lambda_{n}\geq 0.$ (3)
Infrared dynamics is therefore organized by the distribution of these decay scales rather than by Hamiltonian excitation energies.
Observable collective responses—charge, spin, pairing, current, and phase-coherence fluctuations—arise from projections of the full relaxation spectrum onto experimentally accessible operators. The resulting low-energy theory is thus controlled by the statistical structure of relaxation rates. We define the TDOS as the infrared distribution of Liouvillian relaxation eigenvalues,
$\rho(\lambda)=\sum_{n}\delta\!\left(\lambda+\mathrm{Re}\,\lambda_{n}\right),$ (4)
or its continuum limit.
The spectral structure introduced here should therefore be understood at the level of the full dynamical generator rather than solely in terms of Hamiltonian energy eigenmodes. Energy spectra characterize the conservative limit in which all Liouvillian eigenvalues lie on the imaginary axis, whereas generic interacting or coarse-grained systems develop complex spectra with nonzero decay rates. A detailed discussion of this relaxation-rate (Liouvillian) formulation and its connection to memory kernels is provided in Appendix A.
II.1 A. Dynamical stability spectrum and time-scale density of states
The infrared structure of correlated matter is encoded in the spectrum of collective relaxation rates. Rather than softening a single order parameter, correlated critical regimes reorganize the entire decay spectrum, producing a dense accumulation of near-marginal modes.
We characterize this structure by introducing the TDOS, which measures the density of collective decay channels at relaxation rate $\lambda$. This object plays a role analogous to the energy density of states, but in relaxation-rate space. It determines how many collective modes decay on a given time scale.
In weakly interacting systems, the number of slow modes remains finite, and the TDOS vanishes as $\lambda\to 0$, reflecting the presence of well-defined quasiparticles and isolated collective excitations.
By contrast, in strongly correlated critical regimes, the relaxation spectrum generically develops an extensive accumulation of near-marginal decay channels,
$\rho(\lambda\to 0)=\rho_{0}\neq 0,$ (5)
forming a slow-mode reservoir that dominates infrared collective dynamics.
This accumulation signifies a dynamical approach to marginal stability along a macroscopic manifold of collective directions, rather than along a single order-parameter coordinate.
II.2 B. Functional representation of the relaxation spectrum
To compute collective response functions, we employ a functional representation of the dissipative dynamics. The dissipative quantum dynamics of interacting electrons may be formulated in the real–time Keldysh functional integral framework [23, 24]. Upon projection onto slow collective operators and in the regime where decoherence suppresses quantum interference between forward and backward contours, the theory reduces to a classical MSRJD functional integral describing the infrared stochastic dynamics.
Let $\phi(\mathbf{r},t)$ denote a coarse–grained collective bosonic field representing charge, spin, pairing, or phase fluctuations. Projecting the full quantum dynamics onto this operator sector yields an exact decomposition in terms of relaxation eigenmodes.
Because the dissipative spectrum consists of decay rates $\lambda\geq 0$, the projected dynamics may be represented as a continuum of relaxational modes labeled by their decay scales,
$\dot{\phi}$ $=-(r+c\nabla^{2})\phi+\int_{0}^{\Lambda}d\lambda\,g(\lambda)X_{\lambda}+\eta,$ (6)
$\dot{X}_{\lambda}$ $=-\lambda X_{\lambda}+g(\lambda)\phi+\zeta_{\lambda},$ (7)
where $X_{\lambda}$ represents a decay mode with relaxation rate $\lambda$, and $g(\lambda)$ encodes its overlap with $\phi$.
Introducing response fields $\tilde{\phi}$ and $\tilde{X}_{\lambda}$, the MSRJD action reads
$S=$ $\int dt\,d^{d}r\,\tilde{\phi}\Big[\dot{\phi}+(r+c\nabla^{2})\phi-\int_{0}^{\Lambda}d\lambda\,g(\lambda)X_{\lambda}\Big]-D_{\phi}\tilde{\phi}^{2}$
$+\int dt\,d^{d}r\int_{0}^{\Lambda}d\lambda\,\tilde{X}_{\lambda}\big[\dot{X}_{\lambda}+\lambda X_{\lambda}-g(\lambda)\phi\big]$
$-\int dt\,d^{d}r\int_{0}^{\Lambda}d\lambda\,D_{\lambda}\tilde{X}_{\lambda}^{2}.$ (8)
Because the reservoir fields $X_{\lambda}$ enter quadratically, they can be integrated out exactly. Performing the Gaussian functional integration generates a nonlocal-in-time memory kernel in the effective equation of motion,
$\partial_{t}\phi(t)=-(r+c\nabla^{2})\phi(t)-\int_{0}^{t}dt^{\prime}\,K(t-t^{\prime})\,\phi(t^{\prime})+\eta_{\rm eff}(t),$ (9)
where the memory kernel is given exactly by
$K(t)=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\,e^{-\lambda t},$ (10)
with the time-scale density of states
$\rho(\lambda)\equiv g(\lambda)^{2}.$ (11)
Equation (10) makes explicit that the slow-mode reservoir is entirely controlled by the distribution of relaxation rates. The long-time structure of $K(t)$ therefore encodes the infrared organization of decay modes in the interacting system.
Fourier transforming the kernel yields the retarded self-energy,
$\Sigma^{R}(\omega)=\int_{0}^{\Lambda}d\lambda\frac{\rho(\lambda)}{\lambda-i\omega},$ (12)
so that the effective action becomes
$S_{\rm eff}=\int_{\omega,\mathbf{q}}\tilde{\phi}(-\omega,-\mathbf{q})\Big[-i\omega+r+cq^{2}+\Sigma^{R}(\omega)\Big]\phi(\omega,\mathbf{q})+\cdots.$ (13)
Separating real and imaginary parts,
$\mathrm{Re}\,\Sigma^{R}(\omega)$ $=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{\lambda}{\lambda^{2}+\omega^{2}},$ (14)
$\mathrm{Im}\,\Sigma^{R}(\omega)$ $=\omega\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{1}{\lambda^{2}+\omega^{2}}.$ (15)
The collective inverse susceptibility therefore assumes the exact spectral representation
$\chi_{R}^{-1}(\omega,\mathbf{q})$ $=r+cq^{2}+\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{\lambda}{\lambda^{2}+\omega^{2}}$ (16)
$\quad+i\omega\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{1}{\lambda^{2}+\omega^{2}}.$
Equation (16) provides an exact infrared response fully controlled by the relaxation-rate density. The slow-mode reservoir thus appears as a memory-dominated fixed point, whose long-time kernel structure governs the dynamical amplification mechanism discussed below.
The explicit Gaussian integration and derivation of the memory kernel are presented in Appendix B.
II.3 C. Infrared universality classes
[…]
$\rho(\lambda)\sim\lambda^{\alpha},\qquad\lambda\to 0,$ (17)
the collective susceptibility follows directly from the spectral representation in Eq. (16).
The dissipative part of the kernel is
$\mathrm{Im}\,\chi_{R}^{-1}(\omega)=\omega\int_{0}^{\Lambda}d\lambda\,\frac{\rho(\lambda)}{\lambda^{2}+\omega^{2}}\;\sim\;\mathrm{sgn}(\omega)\,|\omega|^{\alpha}.$ (18)
Because the response function is causal, the reactive component is not independent but is related to the dissipative part through the Kramers–Kronig relation. The real part therefore inherits the same infrared scaling structure and, depending on the TDOS exponent, produces logarithmic or subleading corrections to the static stiffness.
It is convenient to express the long–wavelength response in the scaling form
$\chi^{-1}(\omega,\mathbf{q})=r+cq^{2}+\gamma|\omega|^{s},$ (19)
where the TDOS scaling directly implies the identification $s=\alpha$.
While this form resembles the conventional Hertz–Millis representation, its physical interpretation is different. In the present framework the exponent $s$ does not originate from fermionic damping of a single order–parameter mode but instead reflects the infrared organization of the entire relaxation spectrum.
The appropriate renormalization viewpoint is therefore naturally formulated in terms of temporal coarse–graining. Integrating out fast relaxation processes with decay rates $\lambda>\Lambda/b$ corresponds to eliminating short time scales $t\lesssim b^{-1}\Lambda^{-1}$. Under this temporal renormalization step the remaining effective theory is characterized by a renormalized TDOS
$\rho_{b}(\lambda)=b^{\alpha}\rho(b\lambda),$ (20)
so that the infrared scaling exponent $\alpha$ defines the universality class of the relaxation spectrum. Infrared universality classes are therefore classified directly by the low–energy structure of the relaxation–rate density.
A particularly important case occurs when the TDOS remains finite at the origin,
$\rho(\lambda)\xrightarrow{\lambda\to 0}\rho_{0}.$ (21)
In this flat–TDOS regime the dissipative kernel simplifies to
$\mathrm{Im}\,\chi_{R}^{-1}(\omega)$ $=\omega\int_{0}^{\Lambda}d\lambda\,\frac{\rho_{0}}{\lambda^{2}+\omega^{2}}$
$\xrightarrow{|\omega|\ll\Lambda}\frac{\pi}{2}\rho_{0}\,\mathrm{sgn}(\omega),$ (22)
while the reactive component develops a logarithmic infrared singularity,
$\mathrm{Re}\,\chi_{R}^{-1}(\omega)=\int_{0}^{\Lambda}d\lambda\,\rho_{0}\frac{\lambda}{\lambda^{2}+\omega^{2}}\simeq\rho_{0}\ln\frac{\Lambda}{|\omega|}.$ (23)
Collective infrared dynamics is therefore governed by the marginal non–Markovian kernel
$\chi_{R}^{-1}(\omega,\mathbf{q})\simeq r+cq^{2}+\rho_{0}\ln\frac{\Lambda}{|\omega|}+i\frac{\pi}{2}\rho_{0}\,\mathrm{sgn}(\omega)+\cdots,$ (24)
which corresponds in the time domain to a universal long–memory response $K(t)\sim 1/t$, derived explicitly in Appendix C.
This flat–TDOS fixed point represents a dynamical regime in which the infrared behavior is governed by the accumulation of a continuum of slow decay modes. Rather than arising from overdamped fluctuations of a single soft order parameter, collective relaxation is controlled by the spectral organization of the full relaxation spectrum.
III III. Memory–Enhanced Pairing Beyond Eliashberg Theory
The dynamical framework developed in the previous section implies that strongly correlated quantum matter can self–organize into a regime dominated by a continuum of slow collective modes. Instead of a single overdamped bosonic mediator, the infrared dynamics is governed by the full relaxation spectrum characterized by the time–scale density of states $\rho(\lambda)$.
This reorganization of the collective spectrum qualitatively alters the structure of the Cooper channel. In conventional approaches superconductivity is typically attributed to the exchange of a specific bosonic “pairing glue”. In contrast, in the present framework superconductivity emerges from the dynamical amplification of pairing correlations generated by the slow–mode reservoir itself.
In this section we derive the infrared structure of the Cooper channel in this regime and contrast it with the conventional Eliashberg mechanism.
III.1 A. Marginal pairing from overdamped collective dynamics
Near metallic quantum critical points, electrons often couple to overdamped collective modes whose retarded propagator takes the form
$D_{R}^{-1}(\omega,\mathbf{q})=r+cq^{2}+\gamma|\omega|.$ (25)
Here $D_{R}$ denotes the effective propagator of a coarse-grained order-parameter (particle–hole) collective mode in the Hertz–Millis sense: the $|\omega|$ term is generated by the electron–hole polarization of the metallic continuum (Landau damping) [3, 4]. This object should be distinguished from the Cooper-channel polarization $\Pi(\omega)$ introduced below, which diagnoses the pairing instability itself.
The linear frequency dependence reflects Landau damping due to decay of collective excitations into the electronic continuum. After momentum integration, the pairing kernel entering the Cooper channel behaves approximately as
$D_{R}(\omega)\sim\frac{1}{|\omega|}.$ (26)
The linearized gap equation may be written schematically as
$\Delta(\omega)=\int_{T}^{\Lambda}\frac{d\omega^{\prime}}{|\omega^{\prime}|}\,g(\omega-\omega^{\prime})\,\Delta(\omega^{\prime}),$ (27)
where $\Lambda$ is the ultraviolet cutoff and $T$ provides the infrared cutoff.
To analyze the infrared flow we eliminate fast frequency modes with $\omega^{\prime}\in[\Lambda/b,\Lambda]$, corresponding to coarse–graining of short time scales. The ladder diagram gives the leading correction
$dg\propto g^{2}\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{\omega^{\prime}}.$ (28)
The shell integral yields
$\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{\omega^{\prime}}=\ln b.$ (29)
Introducing the RG scale $l=\ln b$ gives
$\frac{dg}{dl}=Cg^{2}.$ (30)
The Cooper interaction therefore grows only logarithmically under infrared coarse graining. In RG language the pairing interaction is marginal. Physically, overdamped collective modes dissipate into the electronic continuum and repeatedly lose phase coherence. As a result superconducting correlations accumulate only logarithmically as the energy scale is reduced.
III.2 B. Collective relaxation spectrum and TDOS representation
Strongly correlated quantum matter, however, can develop a much richer infrared structure. Instead of a single overdamped bosonic excitation, the system may host a dense continuum of slow collective modes.
The collective dynamics can then be represented by the TDOS $\rho(\lambda)$, which measures the density of relaxation rates $\lambda$. Projecting the dynamics onto the slow collective sector yields the spectral representation
$\Sigma(\omega)=\int_{0}^{\Lambda}d\lambda\frac{\rho(\lambda)}{\lambda-i\omega}.$ (31)
The polarization kernel entering the Cooper channel becomes
$\Pi(\omega)=\int_{0}^{\Lambda}d\lambda\frac{\rho(\lambda)}{\lambda^{2}+\omega^{2}}.$ (32)
Here $\Pi(\omega)$ represents the polarization of the Cooper channel generated by the collective relaxation spectrum. It should be distinguished from the conventional bosonic polarization that appears in propagator renormalization. In the present framework $\Pi(\omega)$ directly measures the dynamical pair susceptibility produced by the slow collective modes.
For a flat TDOS in the infrared,
$\rho(\lambda\to 0)=\rho_{0},$ (33)
the polarization acquires a universal singular form
$\Pi(\omega)\sim\frac{\rho_{0}}{|\omega|}.$ (34)
In the time domain this corresponds to a long–memory kernel
$K(t)\sim\frac{1}{t}.$ (35)
Importantly, this singular behavior does not originate from Landau damping into a fast electronic bath. Instead it arises from the spectral accumulation of slow relaxation modes themselves.
III.3 C. Infrared amplification of pairing correlations
The Cooper susceptibility takes the form
$\chi_{\Delta}^{-1}(\omega)=\frac{1}{g_{pair}}-\Pi(\omega).$ (36)
Because $\Pi(\omega)$ diverges as $|\omega|\to 0$, pairing correlations are strongly enhanced as the energy scale is lowered.
To understand the infrared flow more generally, consider a TDOS scaling
$\rho(\lambda)\sim\lambda^{\eta}.$ (37)
Substituting this form into the polarization integral yields
$\Pi(\omega)\sim\int d\lambda\frac{\lambda^{\eta}}{\lambda^{2}+\omega^{2}}\sim|\omega|^{-(1-\eta)}.$ (38)
The effective pairing kernel therefore scales as
$g_{\mathrm{eff}}(\omega)\sim|\omega|^{-(1-\eta)}.$ (39)
Under RG rescaling $\omega\rightarrow\omega/b$ the kernel transforms as
$g^{\prime}=b^{1-\eta}g.$ (40)
Introducing $l=\ln b$ gives the tree–level RG flow
$\frac{dg}{dl}=(1-\eta)g.$ (41)
For $\eta<1$ the Cooper interaction becomes relevant in the infrared. Including the usual ladder correction yields
$\frac{dg}{dl}=(1-\eta)g+Cg^{2}.$ (42)
The linear term dominates the early RG flow, leading to algebraic growth
$g(l)=g_{0}e^{(1-\eta)l}.$ (43)
Physical interpretation. The qualitative difference between the Eliashberg regime and the present memory–dominated regime originates from the structure of the collective spectrum, as illustrated schematically in Fig. 2.
In the Eliashberg case, each logarithmic shell contributes independently, leading to marginal enhancement of pairing correlations. In contrast, a flat TDOS corresponds to a dense continuum of near–marginal relaxation modes. Fluctuations across widely separated time scales remain dynamically correlated, allowing pairing correlations to accumulate coherently.
Superconductivity therefore emerges as a dynamical instability of the collective relaxation spectrum itself, rather than from the exchange of a specific bosonic mediator.
Figure 2: Infrared renormalization-group flow of the Cooper-channel interaction. Shown schematically is the flow of the dimensionless pairing coupling $g$ under infrared coarse graining. Ohmic Landau damping of collective modes produces marginal logarithmic growth, $dg/dl\sim g^{2}$, characteristic of conventional Eliashberg theory (blue). In contrast, memory–dominated dynamics generated by a flat relaxation–rate density of states yields a relevant infrared instability, $dg/dl=\alpha g$, leading to algebraic amplification of the pairing interaction (orange). The slow–mode reservoir therefore dynamically enhances intrinsic electronic pairing tendencies rather than merely renormalizing a fixed microscopic coupling.
IV IV. Pairing from Memory–Dominated Critical Dynamics
The dynamical framework developed in the preceding sections establishes that strongly correlated matter generically self–organizes into a regime dominated by an extensive reservoir of slow collective modes. Rather than producing conventional overdamped fluctuations described by Ohmic Eliashberg dynamics, the accumulation of near–marginal relaxation rates generates long–time memory kernels and nonanalytic infrared response.
Superconductivity in this setting emerges as a dynamical instability of the memory–dominated collective spectrum. The Cooper channel is reorganized by the slow–mode reservoir into a relevant infrared interaction, fundamentally distinct from marginal pairing mediated by fast bosonic modes. Consequently, even moderate intrinsic electronic pairing scales are dynamically amplified into an algebraic superconducting instability.
In this section we demonstrate explicitly how microscopic electronic correlations seed pairing and how the slow–mode TDOS reorganizes the Cooper channel into a memory–enhanced critical regime, ultimately yielding algebraic scaling of the transition temperature and robust phase coherence.
IV.1 A. Intrinsic electronic pairing and dynamical formulation
In strongly correlated materials proximate to Mott insulating phases, short–range Coulomb interactions naturally generate local singlet correlations through exchange processes. Starting from the Hubbard Hamiltonian,
$H=-t\sum_{\langle ij\rangle\sigma}c_{i\sigma}^{\dagger}c_{j\sigma}+U\sum_{i}n_{i\uparrow}n_{i\downarrow},$ (44)
projection to the strong–coupling sector yields the superexchange interaction
$H_{J}=J\sum_{\langle ij\rangle}\Big(\mathbf{S}_{i}\cdot\mathbf{S}_{j}-\frac{1}{4}n_{i}n_{j}\Big),\qquad J=\frac{4t^{2}}{U},$ (45)
which favors local singlet formation. This interaction provides a natural microscopic origin of electronic pairing in systems near Mott localization.
More generally, interacting electrons are described by a generic four–fermion interaction
$H_{\rm int}=\sum_{k,k^{\prime},q}V(q)\,c^{\dagger}_{k+q\sigma}c^{\dagger}_{k^{\prime}-q\sigma^{\prime}}c_{k^{\prime}\sigma^{\prime}}c_{k\sigma}.$ (46)
This interaction contains multiple collective channels. In conventional diagrammatic approaches these channels are analyzed through resummations such as particle–hole bubble diagrams or particle–particle ladder diagrams. Rather than focusing on a specific mediator, the present framework emphasizes the dynamical organization of the collective spectrum itself.
While the four–fermion interaction generically allows multiple ordering channels, their ordering tendencies are not equivalent. Particle–hole instabilities such as charge or spin density waves typically require specific momentum structure, most notably Fermi surface nesting or commensurate wave vectors determined by the lattice. Such kinematic constraints restrict the available phase space for density–wave ordering.
Superconductivity, by contrast, corresponds to condensation in the particle–particle channel at zero total momentum,
$\Delta=\langle c_{k\uparrow}c_{-k\downarrow}\rangle,$ (47)
and therefore does not rely on special nesting conditions of the Fermi surface. Consequently the Cooper channel can efficiently accumulate infrared spectral weight generated by collective fluctuations.
Within the dynamical framework developed in Sec. II, the pairing susceptibility assumes the universal tt–matrix form
$\chi_{\Delta}^{-1}(\omega)=\frac{1}{g_{\rm pair}}-\Pi(\omega),$ (48)
where $g_{\rm pair}$ characterizes the intrinsic electronic pairing scale $E_{\rm pair}$ generated by short–range correlations. Diagrammatically $\Pi(\omega)$ corresponds to the Cooper bubble, which resums particle–particle ladder diagrams.
In weakly interacting systems the pairing instability is exponentially suppressed as in conventional BCS theory. Near correlated criticality, however, the situation changes qualitatively. Because the relaxation spectrum becomes dense near $\lambda\to 0$, an extensive reservoir of slow collective modes emerges. This slow–mode reservoir reorganizes the infrared dynamics of the electronic system and strongly enhances the Cooper channel.
Within the present framework critical dynamics does not create attraction. Instead it amplifies a preexisting electronic pairing tendency through infrared spectral reorganization. The intrinsic pairing scale $E_{\rm pair}$ therefore acts as the microscopic seed, while the slow collective spectrum dynamically amplifies the Cooper instability.
IV.2 B. Memory–dominated Cooper kernel from flat TDOS
Because single–particle propagation inherits the relaxational spectral structure generated by the slow–mode reservoir, the Cooper bubble reduces to a continuum integral over relaxation rates,
$\Pi(\omega)\sim\int_{0}^{\Lambda}d\lambda\,\rho_{\rm eff}(\lambda)\frac{1}{\lambda^{2}+\omega^{2}}.$ (49)
When the slow–mode reservoir develops,
$\rho_{\rm eff}(\lambda\to 0)=\rho_{0}\neq 0,$ (50)
the infrared behavior becomes
$\Pi(\omega)\xrightarrow{|\omega|\ll\Lambda}\frac{\pi\rho_{0}}{2}\,\frac{1}{|\omega|}.$ (51)
This nonanalytic enhancement reflects the accumulation of collective memory across widely separated time scales. Physically, the flat TDOS implies that an extensive set of slow collective modes contributes coherently to the Cooper channel. Unlike conventional Eliashberg kernels, which remain marginal in the infrared, the present Cooper kernel becomes strongly enhanced by the infrared spectral weight of the slow–mode reservoir. The pairing interaction is therefore dynamically amplified by the collective memory of the system.
IV.3 C. Algebraic Instability and Transition Scale
The superconducting transition follows from the Thouless criterion
$\chi_{\Delta}^{-1}(\omega\sim T_{c})=0,$ (52)
which yields the algebraic transition scale
$T_{c}\sim\rho_{0}\,E_{\rm pair},$ (53)
up to numerical factors of order unity. The finite-temperature regularization of the infrared divergence, leading to $\Pi(T)\sim 1/T$, is derived explicitly in Appendix D.
The exponential suppression characteristic of BCS theory is thus replaced by algebraic scaling governed by the infrared spectral weight of the slow–mode reservoir. High transition temperatures arise generically from dynamical amplification of moderate intrinsic pairing scales.
It is important to emphasize that superconductivity requires not only pair formation but also global phase coherence. The transition temperature is therefore ultimately bounded by the smaller of the pairing amplitude and the superfluid phase stiffness,
$k_{B}T_{c}\sim\min\left(\Delta,\,\rho_{s}\right),$ (54)
with $\rho_{s}\sim n_{s}/m^{*}$. The detailed scaling of the phase stiffness within the slow–mode framework is discussed in Sec. V.
The key result of the present section is that the pairing scale itself is controlled by a single infrared parameter, the zero–frequency TDOS $\rho_{0}$. The slow–mode reservoir therefore acts as a universal dynamical amplifier: it converts intrinsic electronic correlation energy into an algebraically enhanced instability of the Cooper channel, independent of any specific bosonic mediator.
In this sense, the TDOS–controlled fixed point defines a dynamical stability class distinct from both weak–coupling BCS superconductors and conventional Eliashberg scenarios. The emergence of superconductivity is governed not by the exchange of a particular glue, but by infrared spectral organization of collective relaxation modes.
Figure 3: Experimental consequences of memory–dominated critical dynamics. (a) Superconducting dome emerging from truncation of the slow–mode reservoir away from criticality. The transition temperature follows $T_{c}(p)\sim E_{\rm pair}\rho_{0}(p)$ and is maximal near the correlated critical point $p_{c}$, where infrared spectral weight of collective relaxation modes is largest. Suppression of $T_{c}$ on both sides arises solely from the finite infrared extent of the slow–mode reservoir, without invoking competing orders or fine tuning. (b) Uemura scaling collapse between superconducting transition temperature and superfluid stiffness, $T_{c}\propto\rho_{s}$. Both pairing strength and phase coherence are controlled by the same infrared spectral weight $\rho_{0}$ of slow collective modes, yielding universal linear scaling across material families.
V V. Experimental consequences of memory–dominated critical dynamics
[…]
V.1 A. Infrared spectral control and automatic emergence of the dome
The linearized transition scale derived in Eq. (53) shows that superconductivity is governed by the infrared spectral weight $\rho_{0}$ of the slow–mode reservoir. We interpret this regime not as a finely tuned quantum critical point, but as a self–organized dynamical fixed point of the relaxation spectrum.
In this phase, feedback between electronic correlations and collective relaxation continuously regenerates an extensive near–marginal sector, leading to a scale–free infrared TDOS,
$\rho(\lambda\to 0)=\rho_{*}>0,$ (55)
without the need for fine tuning of microscopic parameters. The flat infrared spectrum therefore characterizes a dynamical phase rather than a measure–zero critical point.
External control parameters such as doping or pressure primarily modulate the effective infrared cutoff scale of this reservoir. Let $p$ denote a tuning parameter and $p_{\star}$ the point at which the infrared cutoff is minimal. The collective mass term behaves as
$r(p)\sim|p-p_{\star}|,$ (56)
so that the associated cutoff obeys
$\Lambda(p)\sim r^{z\nu}\sim|p-p_{\star}|^{z\nu}.$ (57)
Away from $p_{\star}$, the slow–mode reservoir remains present but is truncated in the deep infrared, yielding
$\rho_{0}(p)\sim\rho_{*}\,f\!\left(\frac{|p-p_{\star}|}{p_{0}}\right),$ (58)
with $f$ a decreasing crossover function.
Substituting into Eq. (53) gives
$T_{c}(p)\sim E_{\rm pair}\,\rho_{*}\,f\!\left(\frac{|p-p_{\star}|}{p_{0}}\right),$ (59)
which naturally produces a superconducting dome. Importantly, the suppression of $T_{c}$ away from $p_{\star}$ does not arise from the disappearance of the dynamical fixed point, but from the progressive truncation of its infrared extent.
In this interpretation, optimal doping does not correspond to a conventional symmetry–breaking quantum critical point. Rather, it marks the regime in which the self–organized infrared spectral phase is most fully exposed.
V.2 B. Phase stiffness and Uemura scaling from the slow–mode reservoir
While Sec. IV established that the pairing scale is algebraically enhanced by the infrared TDOS, the observable transition temperature is ultimately limited by global phase coherence. In systems with reduced superfluid density, the ordering temperature may be controlled by phase stiffness rather than by gap formation.
The long-wavelength phase dynamics of the Cooper field is described by the effective action
$S_{\theta}=\frac{\rho_{s}}{2}\int d\tau d^{d}r\left[(\partial_{\tau}\theta)^{2}+v_{s}^{2}(\nabla\theta)^{2}\right],$ (60)
where $\rho_{s}$ denotes the superfluid stiffness. The corresponding phase-ordering temperature scales as
$T_{\theta}\sim\rho_{s}.$ (61)
Within the slow–mode reservoir framework, the same infrared spectral weight that enhances pairing also governs coherent charge transport. The superfluid stiffness measures the energy cost of imposing a spatial phase twist, and is therefore proportional to the density of coherent paired carriers that remain dynamically correlated over long time scales.
In the memory–dominated regime, long–time coherence is controlled by the infrared TDOS weight $\rho_{0}$. Because the pairing amplitude itself scales as $\Delta\sim E_{\rm pair}\,\rho_{0}$, the density of phase-coherent pairs is likewise proportional to the same infrared spectral weight. Integrating out amplitude fluctuations in the presence of the slow–mode reservoir therefore yields
$\rho_{s}\sim E_{\rm pair}\,\rho_{0},$ (62)
up to nonuniversal numerical factors.
Both pairing amplitude and phase rigidity are thus controlled by a single infrared parameter that quantifies the extent of the slow–mode reservoir.
Combining Eqs. (61) and (62) gives
$T_{c}\sim\rho_{s}\sim E_{\rm pair}\,\rho_{0},$ (63)
establishing a direct proportionality between the superconducting transition temperature and superfluid density.
This relation reproduces the empirical Uemura scaling observed across cuprates and other unconventional superconductors, where $T_{c}$ tracks the zero-temperature superfluid density. In contrast to weak-coupling BCS superconductors—where phase stiffness is large and essentially decoupled from gap formation—correlated systems operate in a regime in which infrared collective dynamics simultaneously controls pairing and coherence.
Within the present framework, this proportionality is not imposed phenomenologically. It follows directly from the TDOS-controlled infrared fixed point, where a single spectral parameter $\rho_{0}$ governs both the Cooper instability and phase stiffness. The observed scaling of $T_{c}$ with superfluid density thus emerges as a natural consequence of memory-dominated critical dynamics.
V.3 C. Dynamical signatures: long–time correlations and anomalous dissipation
A distinctive prediction of the present theory concerns dynamical response in the normal state. As shown in Sec. II, a flat TDOS produces a marginal nonanalytic retarded kernel with
$\mathrm{Im}\,\chi_{R}^{-1}(\omega)\propto\mathrm{sgn}(\omega),$ (64)
corresponding in the time domain to a universal long–memory kernel $K(t)\sim 1/t$. This implies scale–free temporal correlations extending over broad time windows.
Such long–time tails can produce 1/f1/f–type noise spectra and anomalously slow relaxation, phenomena widely reported in strange metals and cuprate superconductors. Unlike conventional overdamped criticality, where correlations decay exponentially and dynamics is effectively Markovian, the memory–dominated regime predicts non-Markovian response with persistent temporal structure. Because the same slow–mode reservoir controls both pairing enhancement and dynamical response, superconducting and normal–state anomalies become two manifestations of a single infrared spectral reorganization.
V.4 D. Unified infrared scaling picture
We emphasize that all experimental signatures discussed above—the dome, Uemura scaling, and long–time dynamical correlations—originate from the same infrared parameter $\rho_{0}$, the low–energy spectral weight of the slow–mode reservoir.
The density of relaxation rates controls: (i) the infrared extent of the slow–mode reservoir, (ii) the dynamical amplification of intrinsic pairing tendencies, (iii) the superfluid stiffness and phase coherence, and (iv) the emergence of non-Markovian long–time response. Quantum criticality thus acts not merely as a source of fluctuations but as a dynamical amplifier governed by infrared spectral reorganization that simultaneously governs thermodynamic and dynamical phenomena in strongly correlated superconductors.
This unified picture yields a concrete strategy for experimental tests: systems that exhibit enhanced low–frequency TDOS weight (or equivalently, strong long–time correlations) are expected to display increased pairing susceptibility, elevated $T_{c}$, and reduced phase stiffness consistent with Uemura scaling. The resulting real-frequency spectral function generated by the memory-dominated self-energy, including its universal scaling form and infrared structure, is derived in Appendix E.
Figure 4: Universality map relating the relaxation-rate spectrum to infrared collective dynamics and superconducting pairing. The TDOS exponent $\alpha$ determines the scaling of the polarization kernel $\Pi(\omega)\sim|\omega|^{\alpha-1}$. The Ohmic Hertz–Millis theory corresponds to the marginal point $\alpha=1$, while a flat TDOS ($\alpha=0$) produces memory-dominated critical dynamics with $\Pi(\omega)\sim 1/|\omega|$ and enhanced pairing.
VI VI. Discussion
The central result of this work is that infrared collective dynamics in strongly correlated quantum matter is organized by the structure of its relaxation-rate spectrum rather than solely by the softening of a single order parameter. The time–scale density of states provides a quantitative measure of this organization. When the TDOS remains finite as $\lambda\to 0$, the system enters a memory–dominated regime characterized by long–time kernels $K(t)\sim 1/t$ and nonanalytic infrared response functions.
This viewpoint clarifies the limits of conventional coarse-grained field theories. Away from criticality, the relaxation spectrum is gapped in the infrared, leading to finite memory times and justifying reduction to local-in-time damping and a small number of collective fields. Near correlated critical regimes, however, relaxation rates condense into a continuous spectrum extending toward zero. In this regime internal modes cannot be integrated out: the full decay spectrum remains dynamically active across time scales. The breakdown of Markovian coarse-graining is therefore not a model-dependent artifact but a universal consequence of dynamical criticality.
This observation highlights a structural difference from Hertz–Millis–type approaches. Conventional quantum critical theories assume overdamped local dynamics of a small set of bosonic modes, with dissipation controlled by Landau damping. In contrast, the present framework allows for spectral condensation of near-marginal relaxation channels. Criticality is reinterpreted not as single-mode softening but as a reorganization of the entire decay spectrum. The TDOS therefore replaces phenomenological damping coefficients with a microscopic spectral quantity governing infrared universality.
This structure is summarized schematically in Fig. 4. The figure illustrates how the infrared universality class of the collective dynamics is determined by the exponent of the relaxation-rate density of states. For a TDOS scaling $\rho(\lambda)\sim\lambda^{\alpha}$, the polarization kernel assumes the universal infrared form
$\Pi(\omega)\sim|\omega|^{\alpha-1}.$ (65)
Different dynamical regimes therefore emerge as special cases of the same spectral organization. For $\alpha>1$ the density of slow modes vanishes in the infrared, so that $\Pi(\omega)\rightarrow 0$ and the collective dynamics reduces to weakly damped propagating modes with $\chi^{-1}\simeq r+cq^{2}$. The marginal case $\alpha=1$ corresponds to Ohmic dissipation, reproducing the Hertz–Millis form $\chi^{-1}\simeq r+cq^{2}+\gamma|\omega|$. Most importantly, when the TDOS approaches a finite constant ($\alpha=0$), a dense continuum of near-marginal relaxation modes emerges. In this regime the polarization acquires the singular form $\Pi(\omega)\sim 1/|\omega|$, producing long-time memory kernels $K(t)\sim 1/t$ and the non-Markovian infrared dynamics discussed throughout this work.
The lower panel of Fig. 4 illustrates the direct consequence of this spectral organization for superconductivity. Within the Thouless criterion $\chi_{\Delta}^{-1}(0,T)=g_{\rm pair}^{-1}-\Pi(T)$, Ohmic dynamics leads to the familiar logarithmic kernel $\Pi(T)\sim\ln(\Omega/T)$ and the exponentially small BCS transition temperature. In contrast, memory-dominated dynamics with a flat TDOS generates the stronger scaling $\Pi(T)\sim C/T$, converting moderate microscopic pairing into an algebraically enhanced instability $T_{c}\sim\rho_{0}E_{\rm pair}$. The figure therefore provides a compact universality map linking the relaxation-rate spectrum, infrared collective dynamics, and the resulting pairing scale.
A direct physical consequence of this spectral reorganization is the enhancement of superconducting instabilities. Superconductivity need not rely on finely tuned bosonic mediators. Instead, intrinsic short-range electronic pairing tendencies are dynamically amplified by the slow-mode reservoir. The exponential suppression characteristic of weak-coupling BCS theory is replaced by algebraic scaling controlled by infrared spectral weight. In this sense, superconducting order corresponds to condensation within an extended slow-mode manifold rather than binding via a sharp bosonic excitation.
A single infrared parameter, $\rho_{0}$, governs both thermodynamic and dynamical phenomena. The superconducting dome reflects truncation of long-time memory away from the optimal spectral reorganization point; Uemura scaling follows from the shared spectral origin of pairing and phase coherence; and anomalous normal-state dynamics arises from the same non-Markovian kernel. Quantum criticality therefore acts as a dynamical amplifier linking transport, coherence, and pairing within a unified spectral framework.
Liouvillian description of infrared collective dynamics.
The relaxation-rate organization emphasized here is not merely a phenomenological assumption. Both Liouvillian evolution in open quantum systems and generalized memory-kernel dynamics naturally decompose into continua of decay modes characterized by complex relaxation eigenvalues. Whenever long-time correlations are present, Laplace-domain representations generically produce continuous relaxation spectra. A dense TDOS therefore reflects a universal mathematical property of dynamical systems with extended temporal correlations rather than a model-specific constructione, is derived in Appendix F.
This viewpoint is closely related to the distinction between closed and effective open descriptions of many-body dynamics. While the complete microscopic system evolves unitarily under a Hamiltonian, the experimentally accessible collective sector behaves as a coarse-grained subsystem continuously exchanging coherence and information with the remaining degrees of freedom. From the perspective of such a subsystem, the relevant generator of dynamics is not the Hamiltonian itself but the effective Liouvillian governing relaxation and memory processes. The infrared behavior of strongly correlated matter is therefore naturally organized by the spectrum of relaxation eigenvalues rather than solely by the spectrum of excitation energies. In this sense the TDOS plays a role analogous to a Liouvillian spectral density, providing a dynamical counterpart to the Lehmann representation familiar from Hamiltonian many-body theory.
From this perspective Hamiltonian energy eigenmodes describe conservative reversible dynamics, whereas relaxation spectra encode the irreversible organization governing collective behavior near criticality. Strongly correlated quantum matter near dynamical critical points may thus be more naturally classified by its relaxation-rate spectrum than by its excitation spectrum alone.
Temporal scaling and dynamical criticality.
An important conceptual aspect of the present framework concerns the role of temporal scaling in dynamical criticality. In conventional renormalization-group treatments of equilibrium critical phenomena, spatial scale invariance provides the primary structure, while temporal scaling emerges indirectly through the dynamic exponent $z$, relating relaxation time and correlation length via $\tau\sim\xi^{z}$. The Hertz–Millis theory of quantum critical metals follows a similar philosophy, in which spatial critical fluctuations generate an effective dynamical scaling once fermionic degrees of freedom are integrated out.
In contrast, the relaxation-spectrum formulation adopted here emphasizes the infrared organization of time scales themselves. When the density of relaxation rates develops a finite limit $\rho(\lambda\to 0)=\rho_{0}$, the system possesses a continuum of slow modes extending to arbitrarily long time scales. The resulting dynamics naturally exhibits scale-free temporal behavior, including long-time memory kernels $K(t)\sim 1/t$ and logarithmic infrared response functions. Scale-free temporal dynamics therefore arises directly from the structure of the relaxation spectrum rather than secondarily from spatial criticality. This perspective is closely related to the scale-free lifetime statistics observed in self-organized critical systems and suggests that infrared relaxation spectra provide a natural language for describing dynamical critical phenomena.
Memory–dominated dynamics as a guiding principle for room-temperature superconductivity. Most contemporary efforts toward room-temperature superconductivity focus on strengthening microscopic pairing interactions through pressure, lattice engineering, or exotic chemical composition. Within the present framework, however, enhancing the bare pairing scale alone is insufficient. The decisive control parameter for high transition temperatures is the infrared dynamical organization of collective modes encoded in the relaxation-rate density of states.
A dense reservoir of near-marginal slow modes reorganizes the Cooper channel, converting moderate intrinsic electronic pairing into an algebraically amplified instability with $T_{c}\sim\rho_{0}E_{\rm pair}$. From this viewpoint certain high-pressure hydrides or correlated materials may fail to reach extreme transition temperatures not because pairing is weak, but because their collective spectra do not self-organize into a flat-TDOS regime.
The present theory therefore suggests a complementary design principle for superconductivity beyond material-specific glue engineering: stabilization of dynamical phases that support extensive slow-mode continua [27]. Control of electronic frustration, dimensionality, feedback mechanisms, or proximity to dynamical criticality may be as crucial as increasing microscopic attraction itself. In this perspective high-temperature superconductivity corresponds to a dynamical phase characterized by maximal infrared memory rather than simply by extreme coupling strength.
VII VII. Conclusion
We have identified dynamical spectral organization as a fundamental principle governing infrared behavior in strongly correlated quantum matter. By formulating collective dynamics in terms of the time–scale density of states, we show that criticality is generically controlled not by local Markovian dissipation but by the accumulation of slow collective relaxation modes.
A finite infrared TDOS defines a memory–dominated universality class characterized by long–time kernels $K(t)\sim 1/t$ and nonanalytic dynamical response. Within this regime, intrinsic electronic pairing tendencies are dynamically amplified, converting marginal logarithmic instabilities into algebraic superconducting transitions governed by infrared spectral weight.
This mechanism provides a unified account of key experimental phenomena in unconventional superconductors, including superconducting domes, Uemura scaling of transition temperatures with superfluid stiffness, and anomalous long–time dynamical correlations in the normal state. These signatures arise naturally from infrared spectral reorganization, without requiring specific bosonic mediators or material-dependent fine tuning.
More broadly, the results recast quantum criticality as a dynamical process that enhances collective quantum phenomena through extended temporal correlations. Memory–dominated criticality therefore represents a universality class distinct from traditional dissipative field theories, with implications for non-Fermi liquids, glassy dynamics, and engineered quantum platforms.
In this perspective, high–temperature superconductivity appears as a generic consequence of infrared spectral organization in strongly correlated matter. Systematic exploration of TDOS-controlled dynamics may provide a pathway toward stabilizing and enhancing collective quantum states in future materials and devices.
Acknowledgements—This work was partially supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (IITP-RS-2025-02214780).
The author acknowledges the support of ChatGPT (GPT-5, OpenAI) for assistance in literature review and conceptual structuring during development.
Supplementary Materials
Appendix A Appendix A: Relaxation-rate spectra as the fundamental dynamical basis of collective dynamics
In conventional quantum many-body theory, collective behavior is often organized in terms of energy eigenmodes of an underlying Hamiltonian,
$H|n\rangle=E_{n}|n\rangle,$ (S1)
with dynamical response expressed through oscillatory Fourier components at frequencies set by energy differences. This viewpoint is natural for closed systems whose dynamics is unitary and time-reversal invariant, and it emphasizes stationary states and conservative motion.
However, energy eigenmodes provide only a special organizing basis. In generic interacting settings—especially in open, driven, or internally coarse-grained systems—the physically relevant long-time organization is not set by Hamiltonian energies but by the stability spectrum of the effective dynamical generator. In this work, the central spectral object is therefore the distribution of relaxation rates governing irreversible decay and memory.
(I) Dynamical linearization and relaxation modes: For a general high-dimensional (possibly coarse-grained) dynamical system,
$\dot{x}(t)=F(x(t)),$ (S2)
linearization about a metastable configuration $x^{*}$ yields
$\delta\dot{x}(t)=-A\,\delta x(t),\qquad A=-\left.\frac{\partial F}{\partial x}\right|_{x^{*}}.$ (S3)
The eigenvalue problem
$Av_{i}=\lambda_{i}v_{i}$ (S4)
defines a spectrum of collective relaxation modes with temporal evolution
$\delta x_{i}(t)\sim e^{-\lambda_{i}t}.$ (S5)
The real parts of $\lambda_{i}$ set decay time scales, while imaginary parts encode oscillatory components. Purely oscillatory energy-mode dynamics corresponds to the singular limit $\mathrm{Re}\,\lambda_{i}=0$ for all modes, which is realized only in strictly Hamiltonian, non-dissipative evolution.
Thus the relaxation-rate spectrum $\{\lambda_{i}\}$ constitutes the natural dynamical basis for generic collective phenomena, including metastability, critical slowing down, and non-Markovian memory.
(II) Liouvillian formulation (closed, open, and non-Markovian dynamics): A conceptually complete formulation of nonequilibrium quantum dynamics is given in terms of the density matrix $\hat{\varrho}(t)$ and its dynamical generator. For a closed system, $\hat{\varrho}(t)$ evolves by the von Neumann equation,
$\frac{d\hat{\varrho}}{dt}=-i[H,\hat{\varrho}]\equiv\mathcal{L}_{H}\hat{\varrho}$, (S6)
where $\mathcal{L}_{H}$ is the Hamiltonian Liouvillian superoperator. If $H|n\rangle=E_{n}|n\rangle$, then the operator basis $|n\rangle\langle m|$ diagonalizes $\mathcal{L}_{H}$:
$\mathcal{L}_{H}\big(|n\rangle\langle m|\big)=-i(E_{n}-E_{m})\,|n\rangle\langle m|,$ (S7)
so the Liouvillian eigenvalues are
$\lambda_{nm}=-i(E_{n}-E_{m}),$ (S8)
lying purely on the imaginary axis and reproducing undamped oscillations.
For open quantum systems, the reduced density matrix typically obeys a master equation with a dissipative contribution. In the Markovian limit, the evolution takes Lindblad form,
[…]
where the jump operators $L_{\mu}$ encode decoherence and dissipation. The resulting Liouvillian $\mathcal{L}$ is generically non-Hermitian as a superoperator and therefore possesses complex eigenvalues
$\mathcal{L}R_{\alpha}=\lambda_{\alpha}R_{\alpha},\qquad\mathrm{Re}\,\lambda_{\alpha}\leq 0,$ (S10)
with decay rates set by $\mathrm{Re}\,\lambda_{\alpha}$. The density matrix admits the modal expansion
$\hat{\varrho}(t)=\hat{\varrho}_{\rm ss}+\sum_{\alpha\neq 0}c_{\alpha}\,e^{\lambda_{\alpha}t}\,R_{\alpha},$ (S11)
where $\hat{\varrho}_{\rm ss}$ is the stationary state ($\lambda_{0}=0$) and the long-time approach is controlled by the Liouvillian spectral gap and any near-marginal continuum.
More generally, when the environment or eliminated internal coordinates retain memory, the reduced dynamics becomes non-Markovian and takes a memory-kernel form (Nakajima–Zwanzig structure),
$\frac{d\hat{\varrho}}{dt}=\int_{0}^{t}ds\;\mathcal{K}(t-s)\,\hat{\varrho}(s),$ (S12)
where $\mathcal{K}(t)$ is a causal superoperator kernel. In Laplace space,
$\tilde{\hat{\varrho}}(s)=\big[s-\tilde{\mathcal{K}}(s)\big]^{-1}\hat{\varrho}(0),$ (S13)
so the poles and branch cuts of $s-\tilde{\mathcal{K}}(s)$ define the intrinsic relaxation spectrum. In this sense, even for quantum systems the relevant organizing structure for long-time dynamics is generically a relaxation-rate spectrum (possibly continuous), not an energy spectrum.
(III) Memory kernels and continuous relaxation spectra: The same spectral logic appears already at the level of linear causal dynamics for a collective coordinate $\phi(t)$:
$\dot{\phi}(t)+\int_{0}^{t}ds\,K(t-s)\phi(s)=\eta(t).$ (S14)
Taking the Laplace transform gives
$\tilde{\phi}(s)=\frac{\tilde{\eta}(s)}{s+\tilde{K}(s)}.$ (S15)
The analytic structure of $s+\tilde{K}(s)$ determines the intrinsic relaxation content: isolated poles correspond to discrete decay modes, while branch cuts encode continua of decay rates. Whenever long-time memory is present, the causal kernel admits a spectral representation of the form
$K(t)=\int_{0}^{\infty}d\lambda\,\rho(\lambda)e^{-\lambda t},$ (S16)
which defines the time-scale spectrum. A flat infrared TDOS produces the universal tail $K(t)\sim 1/t$.
Thus continuous relaxation spectra arise generically from causality and memory, independent of microscopic Hamiltonian structure.
(IV) Energy modes as a special limit: The familiar Fourier-mode picture of conservative dynamics corresponds to the singular limit in which the relevant spectrum collapses onto the imaginary axis. In the closed-system Liouvillian case this appears as $\lambda_{nm}=-i(E_{n}-E_{m})$, while in the memory-kernel representation it corresponds to eliminating all decay rates (no weight at $\mathrm{Re}\,\lambda>0$), recovering purely oscillatory behavior. Away from this limit, generic interacting and/or open systems naturally develop broad spectra of relaxation rates controlling infrared dynamics.
(V) Physical implication: The TDOS—the infrared distribution of collective relaxation rates— is therefore the fundamental organizing variable of collective dynamics in correlated systems. Energy eigenmodes describe an integrable, conservative limit, whereas relaxation-rate spectra capture the universal structure of irreversible, critical, and strongly interacting matter.
Memory-dominated criticality corresponds precisely to the formation of an extensive continuum of near-marginal relaxation modes. The resulting TDOS controls dynamical response, pairing enhancement, and infrared universality, providing a deeper and more generally applicable foundation than energy-based descriptions.
Appendix B Appendix B: Gaussian MSRJD integration of a relaxation-mode reservoir and emergence of the self-energy
In this appendix we show how a continuum of relaxational (Ornstein–Uhlenbeck) modes generates a nonlocal memory kernel and the associated frequency-dependent self-energy $\Sigma(\omega)$ through Gaussian functional integration within the Martin–Siggia–Rose–Janssen–De Dominicis (MSRJD) formalism. The derivation follows the standard structure of Gaussian path integration but is formulated in a way adapted to the slow-mode reservoir perspective developed in this work.
The construction should be understood as an effective infrared description of collective dynamics rather than a microscopic derivation from an explicit fermionic Hamiltonian. At the microscopic level one may introduce collective Hubbard–Stratonovich fields and integrate out fast fermionic degrees of freedom, in the spirit of Hertz-type constructions, which generally produces a nonlocal dynamical kernel for the collective sector. Instead of evaluating this kernel microscopically, we represent its infrared structure through a continuum of relaxational modes.
Within this coarse-grained description the relaxational reservoir is not an additional assumption but a spectral representation of a causal memory kernel. Integrating out the reservoir therefore yields a retarded dynamical correction and a frequency-dependent self-energy for the collective field.
Such a formulation is particularly natural in strongly correlated systems where the infrared dynamics is governed not by a single soft mode but by a broad spectrum of slow relaxation processes. In this regime the collective response is controlled by the spectral weight distribution of these modes, leading naturally to the TDOS-based description employed throughout this work.
(I) Coupled stochastic dynamics: We consider a collective degree of freedom $\phi(t)$ coupled linearly to a continuum of reservoir modes $X_{\lambda}(t)$ labeled by their relaxation rates $\lambda>0$. The coupled stochastic dynamics is
$\dot{\phi}(t)$ $=-r\,\phi(t)+\int d\lambda\;g(\lambda)\,X_{\lambda}(t)+\eta(t),$ (S17)
$\dot{X}_{\lambda}(t)$ $=-\lambda\,X_{\lambda}(t)+g(\lambda)\,\phi(t)+\xi_{\lambda}(t).$ (S18)
The noises are assumed Gaussian, stationary, and white:
$\langle\eta(t)\eta(t^{\prime})\rangle$ $=2D_{\phi}\,\delta(t-t^{\prime}),$ (S19)
$\langle\xi_{\lambda}(t)\,\xi_{\lambda^{\prime}}(t^{\prime})\rangle$ $=2D_{\lambda}\,\delta(\lambda-\lambda^{\prime})\,\delta(t-t^{\prime}).$ (S20)
[…]
(II) MSRJD action for the coupled system: For a generic additive-noise Langevin equation $\dot{y}=f(y)+\zeta$ with $\langle\zeta(t)\zeta(t^{\prime})\rangle=2D\,\delta(t-t^{\prime})$, the MSRJD functional integral may be written (Ito convention) as
$Z$ $=\int\mathcal{D}y\,\mathcal{D}\tilde{y}\;$ (S21)
$\exp\!\left[-\int dt\;\tilde{y}(t)\big(\dot{y}(t)-f(y(t))\big)+\int dt\;D\,\tilde{y}(t)^{2}\right],$
up to an overall normalization independent of $y$. Applying (S21) to (S17)–(S18) gives
$Z=\int\mathcal{D}\phi\,\mathcal{D}\tilde{\phi}\;\prod_{\lambda}\mathcal{D}X_{\lambda}\,\mathcal{D}\tilde{X}_{\lambda}\;e^{-S[\phi,\tilde{\phi},\{X_{\lambda},\tilde{X}_{\lambda}\}]},$ (S22)
with total action
$S=S_{\phi}+\int d\lambda\;S_{\lambda},$ (S23)
where
$S_{\phi}$ $=\int dt\;\tilde{\phi}(t)\!\left[\dot{\phi}(t)+r\,\phi(t)-\int d\lambda\;g(\lambda)\,X_{\lambda}(t)\right]$
$\quad-\int dt\;D_{\phi}\,\tilde{\phi}(t)^{2},$ (S24)
$S_{\lambda}$ $=\int dt\;\tilde{X}_{\lambda}(t)\!\left[\dot{X}_{\lambda}(t)+\lambda\,X_{\lambda}(t)-g(\lambda)\,\phi(t)\right]$
$\quad-\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}.$ (S25)
(III) Isolating a single $\lambda$-mode contribution: Fix $\lambda$ and collect all terms in $S$ that contain $X_{\lambda}$ or $\tilde{X}_{\lambda}$. From (S24) we extract
$S_{\phi}\supset-\int dt\;\tilde{\phi}(t)\,g(\lambda)\,X_{\lambda}(t).$ (S26)
Combining (S26) with (S25) yields the $\lambda$-sector action
$S^{(\lambda)}[X_{\lambda},\tilde{X}_{\lambda};\phi,\tilde{\phi}]=\int dt\;\tilde{X}_{\lambda}(t)\big(\dot{X}_{\lambda}(t)+\lambda X_{\lambda}(t)\big)$ (S27)
$\quad-\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,\phi(t)-\int dt\;g(\lambda)\,\tilde{\phi}(t)\,X_{\lambda}(t)$
$\quad-\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}.$
In what follows, $\phi,\tilde{\phi}$ are treated as external sources.
(IV) Functional integration over $X_{\lambda}$: Define the linear operator
$\mathcal{L}_{\lambda}\equiv\partial_{t}+\lambda.$ (S28)
Then the first term in (S27) is
$\int dt\;\tilde{X}_{\lambda}(t)\,\mathcal{L}_{\lambda}X_{\lambda}(t)=\int dt\;\tilde{X}_{\lambda}(t)\big(\partial_{t}X_{\lambda}(t)+\lambda X_{\lambda}(t)\big).$ (S29)
We now rewrite this term so that $X_{\lambda}$ appears without derivatives. Using integration by parts,
$\int dt\;\tilde{X}_{\lambda}(t)\,\partial_{t}X_{\lambda}(t)$ (S30)
$\quad=\big[\tilde{X}_{\lambda}(t)\,X_{\lambda}(t)\big]_{t_{i}}^{t_{f}}-\int dt\;(\partial_{t}\tilde{X}_{\lambda}(t))\,X_{\lambda}(t).$
Assuming boundary terms vanish (e.g. $t_{i}\to$-$\infty$, $t_{f}\to+\infty$ with sufficient decay, or appropriate causal boundary conditions), we drop the bracketed term and obtain
$\int dt\;\tilde{X}_{\lambda}(t)\,\partial_{t}X_{\lambda}(t)=-\int dt\;(\partial_{t}\tilde{X}_{\lambda}(t))\,X_{\lambda}(t).$ (S31)
Therefore,
$\int dt\;\tilde{X}_{\lambda}(t)\,\mathcal{L}_{\lambda}X_{\lambda}(t)=-\int dt\;(\partial_{t}\tilde{X}_{\lambda}(t))\,X_{\lambda}(t)$ (S32)
$+\int dt\;\lambda\,\tilde{X}_{\lambda}(t)\,X_{\lambda}(t)=\int dt\;X_{\lambda}(t)\big(-\partial_{t}+\lambda\big)\tilde{X}_{\lambda}(t).$
Define the adjoint operator
$\mathcal{L}_{\lambda}^{\dagger}\equiv-\partial_{t}+\lambda.$ (S33)
Then (S27) becomes
$S^{(\lambda)}$ $=\int dt\;X_{\lambda}(t)\Big(\mathcal{L}_{\lambda}^{\dagger}\tilde{X}_{\lambda}(t)-g(\lambda)\tilde{\phi}(t)\Big)$ (S34)
$\quad-\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,\phi(t)-\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}.$
Now $X_{\lambda}$ enters only linearly. Hence its functional integral produces a functional delta constraint:
$\int\mathcal{D}X_{\lambda}\;\exp\!\left\{-\int dt\;X_{\lambda}(t)\,A(t)\right\}$ $\propto\delta[A(t)],$ (S35)
where here
$A(t)\equiv\mathcal{L}_{\lambda}^{\dagger}\tilde{X}_{\lambda}(t)-g(\lambda)\tilde{\phi}(t).$ (S36)
Therefore,
[…]
$\quad\exp\!\left[+\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,\phi(t)-\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}\right].$
The Jacobian factor associated with the functional constraint, $|\det\mathcal{L}_{\lambda}^{\dagger}|^{-1}$, is independent of $\phi,\tilde{\phi}$ and can therefore be absorbed into the overall normalization of $Z$.
(V) Solving the delta constraint for $\tilde{X}_{\lambda}$ via Green functions: The constraint
$\mathcal{L}_{\lambda}^{\dagger}\tilde{X}_{\lambda}(t)=g(\lambda)\tilde{\phi}(t),\qquad\mathcal{L}_{\lambda}^{\dagger}=-\partial_{t}+\lambda,$ (S38)
is a linear inhomogeneous ODE for $\tilde{X}_{\lambda}(t)$. Introduce the Green function $G_{\lambda}^{A}(t-t^{\prime})$ solving
$(-\partial_{t}+\lambda)\,G_{\lambda}^{A}(t-t^{\prime})=\delta(t-t^{\prime}).$ (S39)
Because the constraint is written with the adjoint operator $\mathcal{L}_{\lambda}^{\dagger}=-\partial_{t}+\lambda$, its Green function carries advanced support (equivalently, the retarded kernel will reappear after relabelling, ensuring causality of the $\phi$-sector response).
The solution with advanced support is
$G_{\lambda}^{A}(t)=\Theta(-t)\,e^{\lambda t}.$ (S40)
(Indeed, for $t\neq 0$, $(-\partial_{t}+\lambda)G_{\lambda}^{A}=0$ and the jump at $t=0$ produces the delta distribution.) Multiplying (S38) by $G_{\lambda}^{A}$ and integrating, we obtain
$\tilde{X}_{\lambda}(t)$ $=g(\lambda)\int dt^{\prime}\;G_{\lambda}^{A}(t-t^{\prime})\,\tilde{\phi}(t^{\prime}).$ (S41)
[…]
(VI) Effective action: memory kernel (self-energy) and induced colored noise: We now evaluate the two remaining terms in the exponent of (S37) using (S41).
1. The $\tilde{X}_{\lambda}\phi$ term (deterministic memory kernel).
Start from
$\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,\phi(t).$ (S42)
Insert (S41):
$\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,\phi(t)=\int dt\,dt^{\prime}\;g(\lambda)^{2}\,G_{\lambda}^{A}(t-t^{\prime})\,\tilde{\phi}(t^{\prime})\,\phi(t).$ (S43)
Relabel integration variables $(t,t^{\prime})\to(t^{\prime},t)$:
$\int dt\,dt^{\prime}\;g(\lambda)^{2}\,G_{\lambda}^{A}(t^{\prime}-t)\,\tilde{\phi}(t)\,\phi(t^{\prime}).$ (S44)
Using the identity $G_{\lambda}^{A}(t^{\prime}-t)=G_{\lambda}^{R}(t-t^{\prime})$ with the retarded Green function
$G_{\lambda}^{R}(t)\equiv\Theta(t)\,e^{-\lambda t},\qquad(\partial_{t}+\lambda)G_{\lambda}^{R}(t)=\delta(t),$ (S45)
we obtain
$\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,\phi(t)$ $=\int dt\,dt^{\prime}\;\tilde{\phi}(t)\,\underbrace{g(\lambda)^{2}\,G_{\lambda}^{R}(t-t^{\prime})}_{\equiv K_{\lambda}(t-t^{\prime})}\,\phi(t^{\prime}).$ (S46)
Thus a single OU mode contributes the causal memory kernel
$K_{\lambda}(t-t^{\prime})=g(\lambda)^{2}\,\Theta(t-t^{\prime})\,e^{-\lambda(t-t^{\prime})}.$ (S47)
2. The $\tilde{X}_{\lambda}^{2}$ term (induced colored noise kernel).
Next consider
$\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}.$ (S48)
Insert (S41):
$\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}$
$=\int dt\,dt^{\prime}\,dt^{\prime\prime}\;D_{\lambda}\,g(\lambda)^{2}\,G_{\lambda}^{A}(t-t^{\prime})\,G_{\lambda}^{A}(t-t^{\prime\prime})\,\tilde{\phi}(t^{\prime})\,\tilde{\phi}(t^{\prime\prime}).$ (S49)
Define the induced (generally nonlocal) noise kernel
$\mathcal{N}_{\lambda}(t^{\prime},t^{\prime\prime})\equiv D_{\lambda}\,g(\lambda)^{2}\int dt\;G_{\lambda}^{A}(t-t^{\prime})\,G_{\lambda}^{A}(t-t^{\prime\prime}),$ (S50)
so that
$\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}=\int dt^{\prime}\,dt^{\prime\prime}\;\tilde{\phi}(t^{\prime})\,\mathcal{N}_{\lambda}(t^{\prime},t^{\prime\prime})\,\tilde{\phi}(t^{\prime\prime}).$ (S51)
In equilibrium one may impose a fluctuation–dissipation relation linking $\mathcal{N}_{\lambda}$ and $K_{\lambda}$, but this is not required for the deterministic self-energy structure derived below.
3. Summing the reservoir and assembling $S_{\mathrm{eff}}$.
Combining (S46)–(S51) for all $\lambda$ yields the exact effective action
$S_{\mathrm{eff}}[\phi,\tilde{\phi}]=\int dt\;\tilde{\phi}(t)\big(\dot{\phi}(t)+r\phi(t)\big)-\int dt\;D_{\phi}\,\tilde{\phi}(t)^{2}$
$-\int dt\,dt^{\prime}\;\tilde{\phi}(t)\,K(t-t^{\prime})\,\phi(t^{\prime})-\int dt\,dt^{\prime}\;\tilde{\phi}(t)\,\mathcal{N}(t,t^{\prime})\,\tilde{\phi}(t^{\prime}),$ (S52)
where the total memory kernel is the reservoir sum
$K(t-t^{\prime})=\int d\lambda\;K_{\lambda}(t-t^{\prime})=\int d\lambda\;g(\lambda)^{2}\,\Theta(t-t^{\prime})\,e^{-\lambda(t-t^{\prime})},$ (S53)
and likewise $\mathcal{N}=\int d\lambda\,\mathcal{N}_{\lambda}$.
(VII) Frequency-domain self-energy $\Sigma(\omega)$ from the memory kernel: To connect (S52) to the standard response-function form, we Fourier transform using
$f(\omega)=\int_{-\infty}^{\infty}dt\;e^{i\omega t}\,f(t),\qquad f(t)=\int\frac{d\omega}{2\pi}\;e^{-i\omega t}\,f(\omega).$ (S54)
The bilinear term involving $K$ becomes
$\int dt\,dt^{\prime}\;\tilde{\phi}(t)\,K(t-t^{\prime})\,\phi(t^{\prime})$ $=\int\frac{d\omega}{2\pi}\;\tilde{\phi}(-\omega)\,K(\omega)\,\phi(\omega),$ (S55)
where $K(\omega)$ is the Fourier transform of the causal kernel (S53):
$K(\omega)$ $=\int_{-\infty}^{\infty}dt\;e^{i\omega t}\,K(t)=\int d\lambda\;g(\lambda)^{2}\int_{0}^{\infty}dt\;e^{i\omega t}e^{-\lambda t}.$ (S56)
The time integral is elementary:
$\int_{0}^{\infty}dt\;e^{-(\lambda-i\omega)t}$ $=\left[\frac{-1}{\lambda-i\omega}e^{-(\lambda-i\omega)t}\right]_{0}^{\infty}=\frac{1}{\lambda-i\omega},$
$\qquad(\mathrm{Re}\,\lambda>0).$ (S57)
Substituting (S57) into (S56) gives the exact spectral representation
$K(\omega)=\int d\lambda\;\frac{g(\lambda)^{2}}{\lambda-i\omega}.$ (S58)
In the MSRJD response sector, the retarded inverse propagator for $\phi$ is read off from the $\tilde{\phi}(-\omega)\phi(\omega)$ coefficient in $S_{\mathrm{eff}}$:
$G_{R}^{-1}(\omega)=-i\omega+r+K(\omega).$ (S59)
It is therefore natural to define the (retarded) self-energy as
$\Sigma_{R}(\omega)\equiv K(\omega)=\int d\lambda\;\frac{g(\lambda)^{2}}{\lambda-i\omega}.$ (S60)
Our convention is $\Sigma_{R}(\omega)\equiv K(\omega)$ so that $G_{R}^{-1}(\omega)=-i\omega+r+\Sigma_{R}(\omega)$; other sign conventions for the self-energy appear in the literature.
Equations (S58)–(S60) demonstrate that integrating out a continuum of OU relaxation modes produces a frequency- dependent self-energy whose analytic structure is determined by the relaxation-rate spectrum and couplings.
(VIII) Connection to TDOS (relaxation-rate spectral weight): The spectral representation derived above shows that each relaxational mode contributes to the retarded self-energy with pole $(\lambda_{\alpha}-i\omega)^{-1}$ weighted by the square of its overlap with the collective field. Starting from the discrete mode expansion,
$\Sigma_{R}(\omega)=\sum_{\alpha}\frac{g_{\alpha}^{2}}{\lambda_{\alpha}-i\omega},$ (S61)
it is therefore natural to define the TDOS directly as the coupling-weighted spectral distribution of relaxation rates,
$\rho(\lambda)\equiv\sum_{\alpha}g_{\alpha}^{2}\,\delta(\lambda-\lambda_{\alpha}).$ (S62)
With this definition, the continuum form of the self-energy becomes
$\Sigma_{R}(\omega)=\int d\lambda\;\frac{\rho(\lambda)}{\lambda-i\omega}.$ (S63)
In this sense, $\rho(\lambda)$ plays a role directly analogous to the Lehmann spectral function in Hamiltonian many-body theory. For an operator $O$ with energy eigenstates $|n\rangle$, the usual spectral representation involves weights $|\langle 0|O|n\rangle|^{2}$ multiplying poles at energies $E_{n}$. Here, by contrast, the relevant poles are located at the relaxation rates $\lambda_{\alpha}$, and the corresponding spectral weights are the squared overlap amplitudes $g_{\alpha}^{2}$ between the collective field and the relaxation eigenmodes.
Thus the TDOS should be understood not merely as a bare counting density of modes, but as the physically observable spectral weight distribution of the relaxation spectrum for the collective field under consideration. Equation (S63) makes explicit, at the level of a controlled Gaussian MSRJD integration, why the relaxation-rate spectrum provides the natural basis for the effective dynamics and response of the collective field.
(IX) Coarse-grained field extension: We promote the collective coordinate to a coarse-grained field $\phi(\mathbf{r},t)$. At the Gaussian level the MSRJD action reads
$S_{0}=\int dt\,d^{d}r\;\tilde{\phi}\big(\partial_{t}+\Gamma(r-c\nabla^{2})\big)\phi-\int dt\,d^{d}r\,D\,\tilde{\phi}^{2}.$ (S64)
Coupling $\phi$ locally to an OU reservoir $X_{\lambda}(\mathbf{r},t)$ and integrating out $\{X_{\lambda},\tilde{X}_{\lambda}\}$ yields the nonlocal-in-time term
$S_{\rm eff}\supset-\int dt\,dt^{\prime}\,d^{d}r\;\tilde{\phi}(\mathbf{r},t)\,K(t-t^{\prime})\,\phi(\mathbf{r},t^{\prime}),$ (S65)
with $K(\omega)=\int d\lambda\,g(\lambda)^{2}/(\lambda-i\omega)\equiv\Sigma_{R}(\omega)$. Hence,
$\chi^{-1}(\omega,\mathbf{q})=r+cq^{2}+\Sigma(\omega),$ (S66)
up to the conventional factor $\Gamma$ and the $i\omega$ term in the retarded response.
Summary of the exact Gaussian result. Starting from the coupled Langevin system (S17)–(S18), the MSRJD action (S24)–(S25) is Gaussian in the reservoir variables. Performing the functional integration over $\{X_{\lambda},\tilde{X}_{\lambda}\}$ yields the exact nonlocal effective action (S52) with a causal memory kernel (S53). In frequency space the kernel takes the spectral form (S58), which is naturally identified with the retarded self-energy (S60). This derivation is purely algebraic and involves no approximation beyond standard boundary-condition assumptions required for integration by parts.
(X) Relation to the Keldysh functional integral: We briefly clarify the formal relation between the MSRJD formulation employed in the present work and the real–time Keldysh functional integral for driven–dissipative quantum systems.
The MSRJD action for a stochastic dynamical variable $\phi(t)$ with memory kernel and Gaussian noise may be written as
$S_{\mathrm{MSRJD}}$ $=\int dt\,dt^{\prime}\,\tilde{\phi}(t)\left[\delta(t-t^{\prime})(\partial_{t}+r)-K_{R}(t-t^{\prime})\right]$ (S67)
$\quad\phi(t^{\prime})-\int dt\,dt^{\prime}\,\tilde{\phi}(t)\,\mathcal{N}(t,t^{\prime})\,\tilde{\phi}(t^{\prime}),$
where $\tilde{\phi}$ denotes the response field, $K_{R}$ is the retarded dynamical kernel (memory kernel), and $\mathcal{N}$ is the noise correlator introduced above.
The real–time Keldysh functional integral for an open quantum system is naturally formulated in terms of classical and quantum fields, $\phi_{\mathrm{cl}}$ and $\phi_{q}$, defined by the Keldysh rotation
$\phi_{\mathrm{cl}}=\frac{\phi_{+}+\phi_{-}}{2},\qquad\phi_{q}=\phi_{+}-\phi_{-},$ (S68)
where $\phi_{\pm}$ denote fields on the forward and backward branches of the time contour.
To quadratic order in the quantum field, the Keldysh action assumes the general structure
$S_{\mathrm{K}}=\int_{\omega}\begin{pmatrix}\phi_{\mathrm{cl}}&\phi_{q}\end{pmatrix}_{-\omega}\begin{pmatrix}0&K_{A}(\omega)\\ K_{R}(\omega)&iK_{K}(\omega)\end{pmatrix}\begin{pmatrix}\phi_{\mathrm{cl}}\\ \phi_{q}\end{pmatrix}_{\omega},$ (S69)
where $K_{R,A}$ are retarded/advanced kernels and $K_{K}$ is the Keldysh (noise) component. The retarded kernel encodes dissipative response, while the Keldysh component determines fluctuation amplitudes.
In the regime of strong decoherence or large occupation, where quantum interference effects are suppressed, the quantum field $\phi_{q}$ acts as a response field. Identifying
$\phi_{q}\;\leftrightarrow\;i\tilde{\phi}$, (S70)
and integrating out $\phi_{q}$ via a Hubbard–Stratonovich transformation, one recovers a stochastic Langevin equation with noise correlator set by $K_{K}$. The corresponding MSRJD action then takes the form (S67), with the correspondence
$K_{R}(\omega)\;\leftrightarrow\;K_{R}(\omega),\qquad\mathcal{N}(\omega)\;\leftrightarrow\;\tfrac{1}{2}K_{K}(\omega).$ (S71)
Thus the MSRJD formulation used in the present work can be viewed as the classical (or strongly dissipative) limit of the Keldysh functional integral, in which the relevant dynamical information is encoded in the retarded kernel and its associated spectral properties.
Importantly, our theory does not rely on coherent quantum interference or operator-valued field structure. Instead, the central object is the spectrum of the dynamical generator (equivalently, the retarded kernel), whose low-frequency density of states controls the infrared behavior and pairing amplification mechanism discussed in the main text.
Appendix C Appendix C: Exact emergence of a memory kernel from integrating out an OU reservoir
In this Appendix we provide a self-contained derivation showing that integrating out a continuum of relaxational Ornstein–Uhlenbeck (OU) modes generates a nonlocal-in-time (memory) kernel for the remaining collective coordinate. The result makes explicit that long-memory response is the generic outcome of eliminating hidden internal degrees of freedom, and that a flat time-scale density of states yields the universal tail $K(t)\sim 1/t$. This time-domain elimination is equivalent in content to the Gaussian MSRJD integration, but is presented here directly at the level of the coupled equations of motion.
(I) Coupled Markovian dynamics (collective field plus relaxational reservoir): For clarity we first suppress spatial dependence and focus on temporal structure; the extension to $\mathbf{r}$ (including $cq^{2}$ terms) is immediate by Fourier transforming in space. We consider a slow collective variable $\phi(t)$ linearly coupled to a continuum of relaxational OU modes $X_{\lambda}(t)$ labeled by decay rate $\lambda\geq 0$:
$\dot{\phi}(t)$ $=-r\,\phi(t)+\int_{0}^{\Lambda}d\lambda\,g(\lambda)\,X_{\lambda}(t)+\eta(t),$ (S72)
$\dot{X}_{\lambda}(t)$ $=-\lambda\,X_{\lambda}(t)+g(\lambda)\,\phi(t)+\zeta_{\lambda}(t).$ (S73)
Here $r>0$ is the bare restoring (mass) parameter, $\Lambda$ is an ultraviolet cutoff on relaxation rates, and $\eta$ and $\zeta_{\lambda}$ are (possibly independent) noise sources. For $\lambda>0$ each reservoir coordinate is exponentially stable; the sector $\lambda\to 0^{+}$ should be understood as the infrared limit of a continuous spectrum of increasingly slow relaxational modes rather than an exactly conserved degree of freedom. Equations (S72)–(S73) define a Markovian (first-order) dynamical system in the enlarged state space $\{\phi,X_{\lambda}\}$.
(II) Exact solution for the OU reservoir modes: Equation (S73) is linear and can be solved exactly by the integrating-factor method. Multiply (S73) by $e^{\lambda t}$:
$e^{\lambda t}\dot{X}_{\lambda}(t)+\lambda e^{\lambda t}X_{\lambda}(t)=e^{\lambda t}\big[g(\lambda)\phi(t)+\zeta_{\lambda}(t)\big].$ (S74)
The left-hand side is the total derivative
$\frac{d}{dt}\Big(e^{\lambda t}X_{\lambda}(t)\Big)=e^{\lambda t}\big[g(\lambda)\phi(t)+\zeta_{\lambda}(t)\big].$ (S75)
Integrate from an initial time $t_{0}$ to $t$:
$e^{\lambda t}X_{\lambda}(t)-e^{\lambda t_{0}}X_{\lambda}(t_{0})=\int_{t_{0}}^{t}ds\,e^{\lambda s}\big[g(\lambda)\phi(s)+\zeta_{\lambda}(s)\big].$ (S76)
Solving for $X_{\lambda}(t)$ yields the exact causal representation
$X_{\lambda}(t)$ $=e^{-\lambda(t-t_{0})}X_{\lambda}(t_{0})$ (S77)
$\quad+\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}\Big[g(\lambda)\phi(s)+\zeta_{\lambda}(s)\Big].$
This expression already shows that the reservoir coordinate at time $t$ depends on the entire history $\{\phi(s)\}_{s<t}$.
(III) Substitution into the collective equation and emergence of a memory kernel: Substituting (S77) into the collective equation (S72) gives
$\dot{\phi}(t)$ $=-r\,\phi(t)+\int_{0}^{\Lambda}d\lambda\,g(\lambda)\,$
$\Bigg[e^{-\lambda(t-t_{0})}X_{\lambda}(t_{0})+\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}\Big(g(\lambda)\phi(s)+\zeta_{\lambda}(s)\Big)\Bigg]$
$+\eta(t)$
This motivates the definition of the causal memory kernel
$K(t-s)\equiv\Theta(t-s)\int_{0}^{\Lambda}d\lambda\,g(\lambda)^{2}\,e^{-\lambda(t-s)},$ (S79)
so that (LABEL:eq:kernel_identification_step_app) becomes
$\int_{0}^{\Lambda}d\lambda\,g(\lambda)\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}g(\lambda)\phi(s)=\int_{t_{0}}^{t}ds\,K(t-s)\,\phi(s).$ (S80)
With this definition, Eq. (LABEL:eq:kernel_identification_step_app) can be written as
$\dot{\phi}(t)=-r\,\phi(t)+\int_{t_{0}}^{t}ds\,K(t-s)\phi(s)+\eta(t)+\xi_{\rm eff}(t)+J_{\rm ic}(t),$ (S81)
where we have collected the remaining contributions into an effective noise term (generated by the reservoir noise) and an initial-condition transient:
$\xi_{\rm eff}(t)$ $\equiv\int_{0}^{\Lambda}d\lambda\,g(\lambda)\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}\zeta_{\lambda}(s),$ (S82)
$J_{\rm ic}(t)$ $\equiv\int_{0}^{\Lambda}d\lambda\,g(\lambda)\,e^{-\lambda(t-t_{0})}X_{\lambda}(t_{0}).$ (S83)
Equation (S81) is an exact generalized Langevin equation for $\phi(t)$. Although the enlarged system (S72)–(S73) is Markovian, eliminating the hidden reservoir coordinates produces a non-Markovian (memory) term and, generally, colored effective noise.
Long-time limit.
For $t-t_{0}\gg\Lambda^{-1}$, the transient $J_{\rm ic}(t)$ decays. More precisely, decay holds provided the effective spectral weight $g(\lambda)X_{\lambda}(t_{0})$ is integrable near $\lambda=0$ so that the $\lambda\to 0^{+}$ sector does not produce a non-decaying contribution.
(IV) TDOS form and Laplace-transform structure:
As derived in Appendix B, integrating out the continuum of relaxation modes produces a retarded self-energy of the form (S63), where the time-scale density of states is defined as the spectral weight distribution of relaxation eigenmodes,
$\rho(\lambda)=\sum_{\alpha}g_{\alpha}^{2}\,\delta(\lambda-\lambda_{\alpha}).$ (S84)
Here $g_{\alpha}$ represents the overlap amplitude between the collective field and the relaxation eigenmode $\alpha$, so that $g_{\alpha}^{2}$ plays the role of the spectral weight of that mode. In this sense the TDOS is directly analogous to the Lehmann spectral function of Hamiltonian many-body theory, but defined for the relaxation spectrum of the Liouvillian dynamics.
With this definition the memory kernel introduced above can be written in the compact form
$K(t)=\int_{0}^{\Lambda}d\lambda\;\rho(\lambda)\,e^{-\lambda t},\qquad t\geq 0,$ (S85)
showing that the memory kernel is the Laplace transform of the TDOS. Consequently the long-time dynamics of the collective field is fully controlled by the infrared structure of $\rho(\lambda)$.
(V) Flat TDOS implies universal long-memory tail $K(t)\sim 1/t$: If the TDOS is finite at the origin (“flat” TDOS),
$\rho(\lambda)\xrightarrow{\lambda\to 0}\rho_{0},$ (S86)
then for times $t\gg\Lambda^{-1}$ the infrared part of (S85) dominates and we may approximate $\rho(\lambda)\simeq\rho_{0}$ over the relevant range. In that case the kernel can be evaluated exactly:
$K(t)=\int_{0}^{\Lambda}d\lambda\,\rho_{0}\,e^{-\lambda t}.$ (S87)
Perform the change of variables $u=\lambda t$ (so $d\lambda=du/t$):
$K(t)$ $=\rho_{0}\int_{0}^{\Lambda t}\frac{du}{t}\,e^{-u}=\frac{\rho_{0}}{t}\Big(1-e^{-\Lambda t}\Big).$ (S88)
Therefore, at long times $t\gg\Lambda^{-1}$,
$K(t)\simeq\frac{\rho_{0}}{t},\qquad(t\gg\Lambda^{-1}),$ (S89)
which is the universal long-memory tail discussed in the main text.
(VI) Frequency-space form and connection to the retarded self-energy: For completeness we relate the time-domain memory kernel to the standard frequency-space self-energy. Taking the Fourier transform of the reservoir solution (S77) in the stationary (long-time) limit yields
$X_{\lambda}(\omega)=\frac{g(\lambda)}{\lambda-i\omega}\,\phi(\omega)+\frac{1}{\lambda-i\omega}\,\zeta_{\lambda}(\omega),$ (S90)
and substituting into (S72) gives
$\Big[-i\omega+r+\Sigma^{R}(\omega)\Big]\phi(\omega)=\eta(\omega)+\int_{0}^{\Lambda}d\lambda\,\frac{g(\lambda)}{\lambda-i\omega}\zeta_{\lambda}(\omega),$ (S91)
with the retarded self-energy
$\Sigma^{R}(\omega)=\int_{0}^{\Lambda}d\lambda\,\frac{g(\lambda)^{2}}{\lambda-i\omega}=\int_{0}^{\Lambda}d\lambda\,\frac{\rho(\lambda)}{\lambda-i\omega}.$ (S92)
We use the convention $G_{R}^{-1}(\omega)=-i\omega+r+\Sigma^{R}(\omega)$; other sign conventions for the self-energy appear in the literature.
Separating real and imaginary parts using
$\frac{1}{\lambda-i\omega}=\frac{\lambda}{\lambda^{2}+\omega^{2}}+i\,\frac{\omega}{\lambda^{2}+\omega^{2}},$ (S93)
one obtains
$\mathrm{Re}\,\Sigma^{R}(\omega)$ $=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{\lambda}{\lambda^{2}+\omega^{2}},$ (S94)
$\mathrm{Im}\,\Sigma^{R}(\omega)$ $=\omega\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\frac{1}{\lambda^{2}+\omega^{2}}.$ (S95)
For a flat TDOS $\rho(\lambda)\simeq\rho_{0}$ and $|\omega|\ll\Lambda$,
$\mathrm{Im}\,\Sigma^{R}(\omega)$ $\simeq\omega\rho_{0}\int_{0}^{\Lambda}\frac{d\lambda}{\lambda^{2}+\omega^{2}}=\omega\rho_{0}\left[\frac{1}{|\omega|}\arctan\!\Big(\frac{\Lambda}{|\omega|}\Big)\right]$
$\xrightarrow{|\omega|\ll\Lambda}\frac{\pi}{2}\rho_{0}\,\mathrm{sgn}(\omega),$ (S96)
and
$\mathrm{Re}\,\Sigma^{R}(\omega)$ $\simeq\rho_{0}\int_{0}^{\Lambda}d\lambda\,\frac{\lambda}{\lambda^{2}+\omega^{2}}=\frac{\rho_{0}}{2}\int_{0}^{\Lambda}d\lambda\,\frac{2\lambda}{\lambda^{2}+\omega^{2}}$
$=\frac{\rho_{0}}{2}\left[\ln(\lambda^{2}+\omega^{2})\right]_{\lambda=0}^{\lambda=\Lambda}$
$=\frac{\rho_{0}}{2}\ln\!\left(\frac{\Lambda^{2}+\omega^{2}}{\omega^{2}}\right)\simeq\rho_{0}\ln\!\left(\frac{\Lambda}{|\omega|}\right),$
$\qquad(|\omega|\ll\Lambda),$ (S97)
which reproduces the logarithmic reactive singularity discussed in the main text.
Summary. Equations (S81)–(S85) show that integrating out a continuum of OU relaxation modes yields an exact generalized Langevin equation for the collective coordinate with a causal memory kernel $K(t)$. The kernel is the Laplace transform of the TDOS. A finite TDOS at the origin produces the universal long-memory tail $K(t)\sim 1/t$, and the corresponding retarded self-energy exhibits the marginal nonanalytic structure $\mathrm{Im}\,\Sigma^{R}(\omega)\propto\mathrm{sgn}(\omega)$ together with a logarithmic real part. This makes explicit that memory-dominated infrared dynamics arises from the self-organized accumulation of slow internal relaxation modes, rather than from coupling to a fast Markovian bath.
(VII) Conceptual structure (from hidden modes to infrared memory): The derivation above establishes a general structural result. A Markovian dynamical system defined in an extended state space $\{\phi,X_{\lambda}\}$ becomes non-Markovian when the internal degrees of freedom $X_{\lambda}$ are eliminated. The resulting equation for $\phi$ necessarily acquires a causal memory kernel given exactly by the Laplace transform of the time-scale density of states,
$K(t)=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)e^{-\lambda t}.$ (S98)
This relation is purely kinematic and does not rely on perturbation theory, weak coupling, or specific microscopic assumptions. The infrared dynamical structure of the collective field is therefore completely determined by the low-$\lambda$ scaling of $\rho(\lambda)$.
If the TDOS vanishes at the origin, $\rho(\lambda)\to 0$ as $\lambda\to 0$, the kernel decays exponentially and the dynamics reduces to effectively Markovian overdamping. By contrast, when the TDOS remains finite at zero relaxation rate,
$\rho(\lambda\to 0)=\rho_{0}\neq 0,$ (S99)
the Laplace transform produces the universal long-time tail
$K(t)\simeq\frac{\rho_{0}}{t}.$ (S100)
Thus long-memory infrared dynamics is not an additional assumption but an unavoidable consequence of a flat slow-mode spectrum. The TDOS provides the fundamental organizing principle linking microscopic stability structure to macroscopic temporal response.
In this sense, the emergence of memory-dominated criticality reflects a dynamical phase characterized by an extensive accumulation of near-marginal relaxation modes rather than by coupling to a fast bath.
Appendix D Appendix D: Finite-temperature regulation of the infrared divergence
In this Appendix we explicitly demonstrate how the zero–temperature infrared divergence of the polarization kernel, $\Pi(\omega)\sim 1/|\omega|$, is regulated at finite temperature, yielding the universal scaling $\Pi(T)\sim 1/T$. This establishes rigorously that the thermal scale acts as the infrared cutoff in the memory–dominated critical regime.
(I) Spectral integral representation: In the slow–mode reservoir regime the Cooper-channel polarization admits the spectral form
$\Pi(\omega)=\int_{0}^{\Lambda}d\lambda\;\frac{\rho_{\rm eff}(\lambda)}{\lambda^{2}+\omega^{2}}.$ (S101)
[…]
$\Pi(\omega)\simeq\rho_{0}\int_{0}^{\Lambda}d\lambda\;\frac{1}{\lambda^{2}+\omega^{2}}.$ (S102)
The elementary integral yields
$\int_{0}^{\Lambda}\frac{d\lambda}{\lambda^{2}+\omega^{2}}=\frac{1}{|\omega|}\tan^{-1}\!\left(\frac{\Lambda}{|\omega|}\right).$ (S103)
In the low–frequency limit $|\omega|\ll\Lambda$, one obtains
$\Pi(\omega)\simeq\frac{\pi\rho_{0}}{2}\,\frac{1}{|\omega|},$ (S104)
which expresses the nonanalytic infrared enhancement at zero temperature.
(II) Thermal infrared cutoff: At finite temperature the imaginary–time direction is compactified to the interval $\tau\in[0,\beta]$ with $\beta=1/T$, and dynamical correlations are expressed in terms of discrete Matsubara frequencies. As a consequence, continuous access to arbitrarily small frequencies is lost, and the thermal scale provides an effective infrared regulator.
Operationally, the zero–temperature denominator $\omega^{2}$ in Eq. (S102) is replaced by the minimal dynamical scale of order $T^{2}$, yielding
$\Pi(T)\simeq\rho_{0}\int_{0}^{\Lambda}d\lambda\;\frac{1}{\lambda^{2}+(cT)^{2}},$ (S105)
where $c$ is a numerical constant of order unity determined by the Matsubara spectrum.
Evaluating the integral gives
$\Pi(T)=\rho_{0}\,\frac{1}{cT}\tan^{-1}\!\left(\frac{\Lambda}{cT}\right)\simeq\frac{\pi\rho_{0}}{2c}\,\frac{1}{T},\qquad T\ll\Lambda.$ (S106)
Thus the infrared divergence is universally cut off by the thermal scale,
$\Pi(T)\propto\frac{1}{T}.$ (S107)
(III) Equivalent Matsubara summation: The same result follows directly from the finite–temperature Matsubara representation of the polarization bubble,
$\Pi(0,T)\sim T\sum_{\omega_{m}}\int_{0}^{\Lambda}d\lambda\;\frac{\rho_{0}}{\lambda^{2}+\omega_{m}^{2}},\qquad\omega_{m}=(2m+1)\pi T.$ (S108)
Using the standard identity
$T\sum_{\omega_{m}}\frac{1}{\lambda^{2}+\omega_{m}^{2}}=\frac{1}{2\lambda}\tanh\!\left(\frac{\lambda}{2T}\right),$ (S109)
and expanding for $\lambda\ll T$,
$\tanh\!\left(\frac{\lambda}{2T}\right)\simeq\frac{\lambda}{2T},$ (S110)
one obtains
$T\sum_{\omega_{m}}\frac{1}{\lambda^{2}+\omega_{m}^{2}}\simeq\frac{1}{2T}.$ (S111)
Substitution into Eq. (S108) yields
$\Pi(0,T)\sim\int_{0}^{\Lambda}d\lambda\;\rho_{0}\frac{1}{2T}\propto\frac{\rho_{0}}{T},$ (S112)
confirming that finite temperature universally regulates the infrared divergence.
(IV) Implication for the superconducting transition: The Thouless criterion,
$\chi_{\Delta}^{-1}(0,T_{c})=0\qquad\Longleftrightarrow\qquad\frac{1}{g_{\rm pair}}=\Pi(T_{c}),$ (S113)
therefore gives
$T_{c}\sim\rho_{0}\,g_{\rm pair},$ (S114)
or equivalently $T_{c}\sim\rho_{0}E_{\rm pair}$ in terms of the bare electronic pairing scale. The exponential suppression characteristic of BCS theory is thus replaced by algebraic scaling controlled by the infrared spectral weight of the slow–mode reservoir.
Figure S1: Experimental diagnostics distinguishing the memory-dominated TDOS regime from Hertz–Millis (HM) dynamics. (a) Absolute spectral line shape at $q=0$: the TDOS theory yields a broad continuum, whereas HM produces a sharper peak. (b) Momentum dependence of the TDOS spectral function: varying $q$ shifts the hump through $\Delta_{q}=r+cq^{2}$. (c) TDOS prediction (collapse): when plotted versus the scaling variable $u=\Delta_{q}/\rho_{0}+\ln(\Lambda/|\omega|)$, the TDOS spectra collapse onto a universal curve, $A(\omega,q)=\frac{\pi}{\rho_{0}}\,[u^{2}+(\pi/2)^{2}]^{-1}$. (d) Control test: HM spectra do not exhibit the same log-shift collapse. (e) Finite-temperature regulation implemented by $|\omega|\rightarrow\max\{|\omega|,cT\}$ yields infrared saturation. (f) Corresponding $T$-diagnostic at $q=0$: $A_{T}(\omega\!\to\!0)$ varies systematically with $\ln(\Lambda/T)$, providing a practical fitting protocol. All curves are shown in arbitrary units with consistent prefactor conventions.
Appendix E Appendix E: Spectral function generated by memory–dominated self–energy
In this Appendix we derive the real–frequency spectral function associated with the collective retarded susceptibility in the memory–dominated critical regime. We show that a flat slow–mode TDOS produces a broad non–Lorentzian continuum with logarithmic infrared scaling and a universal collapse form.
(I) Retarded susceptibility in the memory–dominated regime: In the infrared regime controlled by a flat TDOS, the retarded inverse susceptibility takes the universal form
$\chi_{R}^{-1}(\omega,\mathbf{q})=\Delta_{\mathbf{q}}+\rho_{0}\ln\!\left(\frac{\Lambda}{|\omega|}\right)+i\Gamma\,\mathrm{sgn}(\omega),$ (S115)
where
$\Delta_{\mathbf{q}}\equiv r+cq^{2},\qquad\Gamma=\frac{\pi}{2}\rho_{0},$ (S116)
and $\Lambda$ is the ultraviolet cutoff of the slow–mode reservoir. Accordingly,
$\chi_{R}(\omega,\mathbf{q})=\frac{1}{\Delta_{\mathbf{q}}+\rho_{0}\ln(\Lambda/|\omega|)+i\Gamma\,\mathrm{sgn}(\omega)}.$ (S117)
Throughout we use $\omega$ for real frequency; Matsubara frequencies are denoted separately in Appendix D.
(II) Definition of the spectral function: The spectral function of collective fluctuations is defined by
$A(\omega,\mathbf{q})\equiv-2\,\mathrm{Im}\,\chi_{R}(\omega,\mathbf{q}).$ (S118)
Writing $\chi_{R}^{-1}=X(\omega,\mathbf{q})+iY(\omega)$ with
$X(\omega,\mathbf{q})$ $=\Delta_{\mathbf{q}}+\rho_{0}\ln\!\left(\frac{\Lambda}{|\omega|}\right),$ (S119)
$Y(\omega)$ $=\Gamma\,\mathrm{sgn}(\omega),$ (S120)
one has $\chi_{R}=(X-iY)/(X^{2}+Y^{2})$, hence
$A(\omega,\mathbf{q})=\frac{2\Gamma}{\left[\Delta_{\mathbf{q}}+\rho_{0}\ln(\Lambda/|\omega|)\right]^{2}+\Gamma^{2}}.$ (S121)
(III) Infrared scaling at criticality: Along the critical manifold $\Delta_{\mathbf{q}}=0$,
$A(\omega,\mathbf{q}=0)=\frac{2\Gamma}{\rho_{0}^{2}\ln^{2}(\Lambda/|\omega|)+\Gamma^{2}}.$ (S122)
For $|\omega|\ll\Lambda$ the logarithm dominates and the universal infrared tail is
$A(\omega)\xrightarrow{|\omega|\to 0}\frac{2\Gamma}{\rho_{0}^{2}}\,\frac{1}{\ln^{2}(\Lambda/|\omega|)}.$ (S123)
Thus the spectral weight is suppressed only logarithmically toward zero frequency, reflecting the accumulation of memory across many time scales.
(IV) Peak / hump scale away from criticality: For finite $\Delta_{\mathbf{q}}$, Eq. (S121) is maximized when
$\Delta_{\mathbf{q}}+\rho_{0}\ln\!\left(\frac{\Lambda}{|\omega_{*}|}\right)=0,$ (S124)
which yields
$|\omega_{*}|=\Lambda\exp\!\left(\frac{\Delta_{\mathbf{q}}}{\rho_{0}}\right).$ (S125)
For $\Delta_{\mathbf{q}}>0$ this scale lies above the infrared window ($|\omega_{*}|\gtrsim\Lambda$) and no low–energy peak develops, whereas for $\Delta_{\mathbf{q}}<0$ a pronounced low–frequency hump appears at $|\omega_{*}|\ll\Lambda$.
(V) Universal scaling collapse: Define the scaling variable
$u(\omega,\mathbf{q})\equiv\frac{\Delta_{\mathbf{q}}}{\rho_{0}}+\ln\!\left(\frac{\Lambda}{|\omega|}\right).$ (S126)
Using $\Gamma=(\pi/2)\rho_{0}$, Eq. (S121) reduces to the universal form
$A(\omega,\mathbf{q})=\frac{\pi}{\rho_{0}}\,\frac{1}{u(\omega,\mathbf{q})^{2}+(\pi/2)^{2}},$ (S127)
so that all momentum and tuning dependence enters only through $u$, implying a collapse of line shapes onto a single curve.
(VI) Finite–temperature infrared cutoff: At finite temperature the logarithmic growth is regulated by the thermal scale. A minimal phenomenological implementation consistent with Appendix D is
$|\omega|\;\rightarrow\;\max\{|\omega|,cT\},\qquad c=O(1),$ (S128)
yielding
$A_{T}(\omega,\mathbf{q})=\frac{2\Gamma}{\left[\Delta_{\mathbf{q}}+\rho_{0}\ln(\Lambda/\max\{|\omega|,cT\})\right]^{2}+\Gamma^{2}}.$ (S129)
For $|\omega|\ll T$ the logarithm saturates, producing a finite infrared spectral weight rather than the zero–temperature logarithmic suppression.
Physical interpretation. Unlike conventional overdamped criticality, where $\mathrm{Im}\,\Sigma^{R}(\omega)$ vanishes linearly with $|\omega|$, the memory–dominated regime exhibits frequency–independent damping ($\propto\mathrm{sgn}(\omega)$) accompanied by a logarithmically singular reactive part. The resulting spectral function is therefore a broad continuum with slowly varying infrared tails rather than a sharp Lorentzian quasiparticle peak.
(VII) Experimental diagnostics and scaling tests: Figure S1 summarizes a set of practical diagnostics that distinguish the memory–dominated TDOS regime from conventional overdamped critical dynamics such as the Hertz–Millis (HM) theory.
Panel (a) compares the absolute spectral line shapes at $q=0$ predicted by the two frameworks. The HM theory produces a relatively narrow Lorentzian-like peak, reflecting dynamics controlled by a small number of overdamped bosonic modes. In contrast, Eq. (S121) yields a much broader continuum with slowly varying logarithmic tails, arising from the collective response of an extensive continuum of slow relaxation modes.
Panel (b) illustrates the predicted momentum dependence of the TDOS spectral function. Because the tuning parameter enters through $\Delta_{\mathbf{q}}=r+cq^{2}$, varying the momentum effectively shifts the spectral hump according to the scale $|\omega_{*}|=\Lambda\exp(\Delta_{\mathbf{q}}/\rho_{0})$ derived in Eq. (S125). This produces a characteristic logarithmic displacement of the hump position as $q$ changes.
Panels (c) and (d) demonstrate the most direct experimental test of the TDOS framework. Introducing the scaling variable
$u(\omega,\mathbf{q})=\frac{\Delta_{\mathbf{q}}}{\rho_{0}}+\ln\!\left(\frac{\Lambda}{|\omega|}\right),$ (S130)
all TDOS spectra collapse onto the universal curve (S127). This collapse reflects the fact that the full line shape depends only on the combined variable $u$. By contrast, spectra generated by the conventional Hertz–Millis form do not collapse under the same transformation, providing a clear diagnostic distinction.
Panel (e) illustrates the finite–temperature regulation discussed in Eq. (S129). Replacing $|\omega|$ by $\max\{|\omega|,cT\}$ cuts off the logarithmic infrared growth and produces saturation of the low–frequency spectral weight. The resulting spectra evolve smoothly from the zero–temperature form to a temperature–dominated plateau.
Finally, panel (f) shows a convenient experimental diagnostic at $q\simeq 0$. In the regime $|\omega|\ll T$, Eq. (S129) predicts that the infrared spectral weight varies systematically with $\ln(\Lambda/T)$. Measuring $A_{T}(\omega\!\rightarrow 0)$ as a function of temperature therefore provides a direct way to extract the TDOS parameter $\rho_{0}$ and test the predicted logarithmic scaling.
Taken together, these diagnostics provide a practical experimental protocol for identifying the memory–dominated TDOS universality class. In particular, the logarithmic peak shift and the universal collapse of line shapes constitute strong signatures of collective dynamics governed by an extensive slow–mode reservoir rather than by a single overdamped critical mode.
Appendix F Appendix F: Possible connections between flat TDOS and scale–free dynamical phenomena
In this Appendix we briefly discuss a broader dynamical perspective suggested by the relaxation–spectrum framework developed in the main text. The central quantity of the theory is the density of relaxation rates associated with the Liouvillian generator of the coarse–grained dynamics.
Near dynamical critical regimes, the spectrum of relaxation rates naturally develops a dense continuum extending toward vanishing decay rates. In particular, an important limiting case is a finite infrared density of slow modes,
$\rho(\lambda\rightarrow 0)=\rho_{0}\neq 0.$ (S131)
This condition corresponds to a situation in which relaxation processes exist across arbitrarily long time scales. As a result, the system cannot be characterized by a single relaxation time, and long–time memory effects emerge naturally.
(I) Long–time memory kernel: The dynamical kernel associated with the relaxation spectrum can be written as
$K(t)=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\,e^{-\lambda t},$ (S132)
where $\Lambda$ denotes the ultraviolet cutoff of the relaxation spectrum.
If the TDOS approaches a constant for $\lambda\ll\Lambda$, $\rho(\lambda)\approx\rho_{0}$, the long–time behavior becomes
$K(t)\approx\rho_{0}\int_{0}^{\Lambda}d\lambda\,e^{-\lambda t}=\frac{\rho_{0}}{t}\left(1-e^{-\Lambda t}\right).$ (S133)
For $t\gg\Lambda^{-1}$ one therefore obtains the algebraic decay
$K(t)\sim\frac{\rho_{0}}{t},$ (S134)
indicating the emergence of long–time memory.
(II) Relation to scale–free relaxation statistics: Distributions of relaxation times with power–law statistics are widely encountered in complex systems, including models of self–organized criticality (SOC). In many SOC models the lifetime distribution of events takes the approximate form
$P(\tau)\sim\tau^{-2}.$ (S135)
Using the relation between lifetime and relaxation rate $\lambda=1/\tau$, one finds
$\rho(\lambda)=P(\tau)\left|\frac{d\tau}{d\lambda}\right|=P(1/\lambda)\frac{1}{\lambda^{2}}.$ (S136)
Substituting $P(\tau)\sim\tau^{-2}$ yields
$\rho(\lambda)\sim{\rm const}.$ (S137)
Thus a flat infrared TDOS corresponds to a relaxation spectrum that is consistent with scale–free lifetime statistics. This observation suggests a possible dynamical connection between the TDOS framework and the emergence of SOC–type temporal scaling.
(III) Implications for low–frequency noise: A broad distribution of relaxation times is also known to generate 1/f1/f–type noise spectra. Within the present framework the noise spectrum can be written as
$S(\omega)\sim\int d\lambda\,\frac{\rho(\lambda)\lambda}{\lambda^{2}+\omega^{2}}.$ (S138)
When the infrared TDOS approaches a constant, the dominant contribution arises from $\lambda\lesssim\omega$, which yields approximately
$S(\omega)\sim\frac{1}{|\omega|}.$ (S139)
This provides a natural dynamical interpretation of the ubiquitous $1/f$ noise observed in many correlated systems.
(IV) Relation to the infrared spectral function: As shown in Appendix E, a flat TDOS produces a characteristic logarithmic structure in the retarded susceptibility,
$\chi_{R}^{-1}(\omega,q)=\Delta_{q}+\rho_{0}\ln\!\left(\frac{\Lambda}{|\omega|}\right)+i\Gamma\,{\rm sgn}(\omega).$ (S140)
The resulting spectral function
$A(\omega,q)=\frac{2\Gamma}{\left[\Delta_{q}+\rho_{0}\ln(\Lambda/|\omega|)\right]^{2}+\Gamma^{2}}$ (S141)
describes a broad continuum rather than a sharp quasiparticle peak. Such continua have been reported in several strongly correlated systems, including strange metals, although the precise microscopic origin remains an open question.
(V) Consequences for superconducting pairing: In the main text we have shown that the same infrared spectral organization modifies the Cooper channel. The pairing susceptibility takes the form
$\chi_{\Delta}^{-1}(\omega)=\frac{1}{g_{\rm pair}}-\Pi(\omega),$ (S142)
with
$\Pi(\omega)\sim\int d\lambda\frac{\rho(\lambda)}{\lambda^{2}+\omega^{2}}.$ (S143)
For $\rho(\lambda\rightarrow 0)=\rho_{0}$ one finds
$\Pi(\omega)\sim\frac{1}{|\omega|},$ (S144)
which leads to the algebraic transition scale
$T_{c}\sim\rho_{0}E_{\rm pair}.$ (S145)
The slow–mode reservoir therefore acts as a dynamical amplifier of intrinsic electronic pairing correlations.
Summary. These observations suggest that the infrared organization of the relaxation spectrum may provide a common dynamical language for several scale–free phenomena observed in correlated systems. Within this perspective the central object is not a single quasiparticle mode but the entire spectrum of collective relaxation processes. Further investigation of this spectral viewpoint may help clarify the role of dynamical criticality in a wide range of nonequilibrium and strongly correlated systems.